### v1

# Programmable Cognitive Bias in Social Agents

Abstract
This paper introduces CoBRA, a novel toolkit for systematically specifying agent behavior in LLM-based social simulation. We found that conventional approaches that specify agent behaviors through implicit natural language descriptions cannot yield consistent behaviors across models, and the produced agent behaviors do not capture the nuances of the descriptions. In contrast, CoBRA presents a new approach to program agents’ cognitive biases explicitly, by grounding agents’ expected behaviors using classic social science experiments. CoBRA has two components: (1) Cognitive Bias Index that measures the cognitive bias of a social agent, by quantifying the agent’s reactions in a set of validated classical social science experiments; (2) Behavioral Regulation Engine that aligns the agent’s behavior to demonstrate controlled cognitive bias. We evaluated CoBRA as an HCI toolkit through demonstration and technical benchmarks. Our results suggest that CoBRA can precisely program the cognitive bias demonstrated in a social agent in a model-agnostic manner.
Figure 1: Existing social simulation experiments often use implicit natural language descriptions to specify agent behaviors. However, we found that these specifications often resulted in inconsistent and unpredictable agent behaviors. For example, Ⓐ real-world economists are supposed to be less susceptible to the Framing Effect than the general population Thaler & Sunstein (2008); however, Ⓑ agents based on implicit natural language specifications often produce inconsistent behaviors across models, and the expected differences in behavior across roles (e.g., economists being less prone than laypeople) are not reliably observed. To tackle this challenge, Ⓒ we introduce CoBRA  which enables researchers to explicitly specify the cognitive biases of LLM-based agents quantitatively, thereby producing precise and consistent behaviors across models.
1 Introduction
Figure 2: Example closed-loop workflow of CoBRA. A social scientist aims to create an agent with a moderate framing effect (e.g., 2.6 on a 0–4 scale). ① She specifies the desired bias level in CoBRA alongside the natural language agent description. ② CoBRA measures the agent’s framing effect using validated classical social science experiments (e.g., the Asian Disease study Tversky & Kahneman (1981)). ③ If the measured bias deviates from the specification, the Behavioral Regulation Engine iteratively adjusts the agent—through prompt engineering, activation modifications, or fine-tuning—until the agent reliably demonstrates the target bias.
The simulation of complex human behaviors through LLM-based agents is rapidly emerging as a powerful tool for social scientists Bail (2024); Park et al. (2023), offering new ways to test social theories Anthis et al. (2025a); Wu et al. (2025) and explore sensitive or hard-to-study behaviors without real-world consequences Hwang et al. (2025). A critical component of these simulations is specifying agent behaviors Yang et al. (2025); Chandramouli et al. (2024); Goyal et al. (2024) to control how the agent responds to prompts. For example, Park et al. specified the attributes of one of the agents in their paper using the following natural language description: ”John Lin is a pharmacy shopkeeper at … who loves to help people. He is always looking for ways to make the process of getting medication easier for his customers…” Park et al. (2023).
These implicit natural language descriptions have become the dominant approach for specifying social agent behaviors and are used in nearly all social simulation experiments Rath et al. (2025); Park et al. (2023); Zhou et al. (2024); Li et al. (2023a); Zheng et al. (2023); Chen et al. (2024b); Jia et al. (2024). In the descriptions, researchers often specify demographic descriptors (e.g., age, occupation, family structure), personality traits (e.g., Big Five Gosling et al. (2003) or MBTI Furnham (1996)), personal interests (e.g., hobbies, preferences), social relationships (e.g., friends, family, colleagues), and professional or task-oriented identities (e.g., judge, programmer). These practices rest on a growing understanding that LLMs, having been trained on vast text corpora, possess both perceptual skills Wang et al. (2024b); Liu et al. (2024) that enable them to detect subtle cues in implicit language Gallegos et al. (2024), and strong role-playing capabilities Park et al. (2023)—drawing on patterns that link behaviors to human characteristics Fast et al. (2016); Shao et al. (2023b).
However, the research community has little empirical understanding of whether this approach is sufficient for reliable social simulations. For example, would the conclusion of the social simulation change if we change the foundation models, the prompts, or the temperatures? As a consequence, the conclusions of most LLM-based social simulations often are broad qualitative impressions—such as noting that a simulation is plausible and exhibits an echo chamber similar to the real world Wang et al. (2024a); Zheng & Tang (2024).
In this paper, we first conducted a pilot experiment to empirically examine the limitations of implicit natural language specification (Section 3). We designed four agent specifications by adapting examples from prior work Chen et al. (2024b); Park et al. (2023), implemented them across four foundation models, and evaluated their responses using a classic social science experiment, the Asian Disease study Tversky & Kahneman (1981). The results show that (1) the same specification produced inconsistent behaviors across models, and (2) implicit specification did not reliably produce the expected behaviors (e.g., trained experts like economists are less prone to the framing effect List (2003)).
We then developed CoBRA—the Cognitive Bias Regulator for Social Agents (Figure 1)—a novel toolkit that enables users to precisely program the cognitive biases exhibited by social agents across different models. Compared to implicit natural language specifications, CoBRA offers two key advantages. First, it enhances reproducibility of agent behaviors across models, which is critical given that access to specific foundation models may change over time Chen et al. (2024a). Second, it provides explicit and quantitative control over agent behaviors, rather than relying on unpredictable model behaviors Sclar et al. (2023). For example, a scientist can directly specify that an agent should be less susceptible to framing effects.
At the heart of CoBRA is a novel closed-loop system primitive that continuously measures and regulates agents’ cognitive biases by grounding their behaviors in classic social science experiments. Classic experiments provide validated and structured protocols: they define precise stimuli, expected behavioral distributions, and well-documented bias effects. For instance, the Asian Disease study Tversky & Kahneman (1981) has been extensively replicated and is widely regarded as a canonical demonstration of the framing effect: when identical outcomes are framed in terms of lives saved versus lives lost, human participants systematically shift between risk-averse and risk-seeking choices. Within CoBRA, such experiments function as calibration tasks: the system first measures whether an agent demonstrates the expected bias by observing its responses in these experiments and then adjusts the agent’s specification to align it with the target behavioral profile (Figure 2). We further hypothesize that aligning an agent’s behavior with target profiles in a set of classical experiments will reliably induce the corresponding cognitive biases, which will then generalize and manifest across a broader range of tasks (Section 7.2).
Our implementation of CoBRA has two components: (1) Cognitive Bias Index that measures the cognitive bias of a social agent by quantifying the agent’s reactions in a set of validated classical social science experiments; (2) Behavioral Regulation Engine that aligns the agent’s behavior to demonstrate controlled cognitive bias. The Cognitive Bias Index contains eight classic social experiments mapped to four representative biases—Authority Effect, Bandwagon Effect, Confirmation Bias, and Framing Effect. The Behavioral Regulation Engine employs three approaches (prompt engineering Rath et al. (2025); Zhuo et al. (2024), representation engineering Zou et al. (2023); Rimsky et al. (2024); Todd et al. (2024), and fine-tuning Ouyang et al. (2022); Rafailov et al. (2023); Hu et al. (2022); Ilharco et al. (2023)) to align agent behaviors.
We position CoBRA as an HCI toolkit Nebeling (2017) that empowers social scientists to create new types of social simulations. We evaluated CoBRA using two common strategies for assessing HCI toolkits Ledo et al. (2018): Technical Benchmark and Demonstration. In the benchmark, we show that (1) the cognitive bias induced through one classical experiment generalizes to other tasks (Section 7.2); (2) agents specified through CoBRA exhibit consistent behavioral tendencies regardless of the foundation model, sampling temperature, or reasoning mode (Section 7.3); and (3) CoBRA enables precise control under both open-weight and API-only settings (Section 7.4). In the Demonstration, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment Kramer et al. (2014), where the participant’s social influence susceptibility (i.e., bandwagon effect bias) can be manipulated (Section 8). The simulated experiment validates a classical social science theory: agents exposed to varying levels of negative posts exhibited predictable sentiment shifts based on their bias levels programmed by CoBRA, with higher Bandwagon Effect levels leading to stronger emotional contagion.
This work makes four main contributions. First, we present the first empirical evidence that implicit natural language specifications, the dominant method for defining LLM-based social agents, fail to produce reliable and reproducible behaviors across models. Second, we introduce CoBRA, a novel HCI toolkit that enables social scientists to precisely program and control cognitive biases in agents, improving reproducibility and allowing explicit, quantitative alignment. Third, we contribute a closed-loop system primitive that grounds agent behaviors in validated and structured social science experiments, using them as primitives for both measuring and regulating cognitive biases. Finally, we demonstrate CoBRA through technical benchmarks and a case study of the Emotional Contagion experiment, showing that induced biases generalize across tasks, remain robust across models and settings, and allow social scientists to create new forms of controlled social simulations.
2 Background & Related Work
This section reviews the use of large language models (LLMs) for social simulation and outlines the prevailing methodological pipeline, noting that current implicit natural-language agent specification approaches—while expressive and flexible—hinder rigor and comparability Bail (2024); Pang et al. (2025).
2.1 Believable Agents for Social Simulation
Believable Agents are interactive computational entities, originating as non-player characters in games Bates (1994). Early designs relied on rule-based systems or behavior trees to dictate their actions and decision-making processes Loyall (1997). Their observable behavior appears coherent, context-appropriate, and consistent with prior experiences, creating an “illusion of life” or human-like plausibility Park et al. (2023); Anthis et al. (2025b); Banovic et al. (2023); Ehsan et al. (2024). The goal is not human-equivalent cognition but externally credible behavior: agents should perceive, plan, and act in ways that are explainable, temporally consistent, and aligned with social and physical norms Goyal et al. (2024); Chandramouli et al. (2024); Liu et al. (2025a); Zhang et al. (2023).
Recent works use LLMs to create Believable Agents for social simulation Bail (2024); Wu et al. (2022); Anthis et al. (2025a); Hwang et al. (2025); Tjuatja et al. (2024); Prpa et al. (2024); Aubin Le Qu’er’e et al. (2024). For example, prior work Park et al. (2023); Zhou et al. (2024) utilizes LLMs’ language understanding to prompt models with natural language persona descriptions, creating LLM-based agents that interact within simulated environments. These agents facilitate the study of emergent phenomena—such as cooperation, norm formation, and conflict—and heuristic behaviors like conformity Liu et al. (2025b); Zhang et al. (2024); Shao et al. (2023a); Wu et al. (2025); Zhu et al. (2025). While these LLM-based Believable Agents broaden feasible questions and accelerate iteration by reducing labor, and mitigating ethical concerns Bail (2024); Hwang et al. (2025), they struggle with consistent, controllable behavior Yang et al. (2025); Han et al. (2025)—a critical need for rigorous social science simulations Bail (2024).
In this work, we adopt the concept of Believable Agents and address prior rigor limitations by introducing CoBRA, a toolkit for systematic agent specification that yields consistent, controllable behavior.
2.2 Reproducibility of LLM-Based Methods in Social Science
Reproducibility is a critical concern in the use of LLMs for social science research, as the probabilistic nature of LLMs combined with varying implementation methods often leads to inconsistent and non-reproducible outcomes Bail (2024); Zhao et al. (2021).
One major issue is that different LLMs exhibit divergent behaviors even with the same prompts Liu et al. (2023). Variations in training datasets Tayo et al. (2023), model architectures Jiang et al. (2024), and fine-tuning processes Touvron et al. (2023) exacerbate this inconsistency, while the opacity of proprietary models further complicates cross-model comparisons and limits generalizability Gudibande et al. (2023). Second, inference-time configurations significantly affect reproducibility. Sampling parameters like temperature settings influence output variability, with models differing in how peaky or flat their probability distributions are, affecting the generation coherence and diversity Peeperkorn et al. (2024). Similarly, reasoning modes Wei et al. (2022) (direct response vs. multi-step reasoning) introduce additional variability during inference, as different reasoning strategies can lead to divergent conclusions from the same input. Finally, post-training methods (e.g., RLHF Ouyang et al. (2022)) also introduce inconsistencies across models, further complicating reproducibility. Addressing these challenges requires standardized methodologies and explicit control mechanisms to ensure reliable results.
In this work, we address reproducibility challenges by complementing the implicit natural language Agent Specification with an explicit approach that leverages structured, validated knowledge from classic social experiments to define agent behaviors, reducing variability and ensuring consistent outcomes.
2.3 Classic Social Experiments
Classic social experiments French et al. (1959); Asch (1951); Tversky & Kahneman (1981); Milgram (1963); Kahneman & Tversky (2013); Wason (1960); Lord et al. (1979); Goldstein et al. (2008); Milgram et al. (1969) provide structured and validated knowledge: long-replicated classic social experimental paradigms operationalize social-psychological constructs into clear manipulations and measurement protocols Likert (1932); Vazire & Mehl (2008); Lilienfeld et al. (2017); the resulting behavioral patterns have been extensively validated and distilled into reusable, codified behavioral patterns Unknown (2015); Baron (2007); Cialdini (1993); Thaler & Sunstein (2008).
As a concrete focus in this paper, we ground agent specification in cognitive biases Unknown (2015); Baron (2007); Sumita et al. (2025)—a representative class of codified behavioral patterns—defined as systematic, predictable deviations from normative judgment and decision-making that arise from heuristics Tversky & Kahneman (1974); Simon (1990). Prominent examples include Confirmation Bias (favoring information that confirms pre-existing beliefs, validated through the Wason Selection task Wason (1960) and Biased Information study Lord et al. (1979)), Bandwagon Effect (adopting behaviors due to social conformity, demonstrated in Asch’s Line experiment Asch (1951) and the Hotel Towel study Goldstein et al. (2008)), and so on. These cognitive biases, repeatedly validated in classic experiments, provide a rigorous foundation for specifying and evaluating LLM-based agent behaviors.
3 Understanding Agent Specification
This section presents a pilot experiment to empirically examine the limitations of implicit natural language specification. The results showed that current methods failed to produce consistent, cognitive bias-aligned behavior.
3.1 Method
Our key idea is to use a classical social experiment, the Asian Disease study Tversky & Kahneman (1981), as the study instrument. We created agents using different specification methods, powered these agents using different foundation models, and then analyzed their responses in the simulated social experiment.
Agent specifications. We evaluated two representative examples of implicit natural language agent specifications: a persona-based specification from Park et al. (2023), which constructs agents with human-like profiles by incorporating demographics, personality traits, interests, and social relationships; and a role-based specification from Chen et al. (2024b), which defines agents through task-oriented identities, assigning roles, domain expertise, and interaction protocols to guide task-specific behavior.
For the persona-based example Park et al. (2023), we introduced three conditions: Common (an ordinary person), Economist (a persona with a professional background expected to make them less susceptible to framing effects and, therefore, less likely to prefer the positive frame compared to the Common persona), and Blank (no specification). We adopted the John Lin prompt from Park et al. (2023) to represent the Common persona (see Appendix B for full prompt):
John Lin is a pharmacy shopkeeper at the Willow Market who loves to help
people. He is always looking for ways to make the process of getting
medication easier for his customers. John Lin is living with his wife,
Mei Lin, who is a college professor, and son, Eddy Lin, who...
In the Economist condition, we modified only the profession, changing “a pharmacy shopkeeper at the Willow Market” to “a professor of economics at a nearby college,” while keeping all other details identical. The Asian Disease scenario prompt was used directly as the input.
For the role-based example Chen et al. (2024b), we introduced two conditions: Teacher (the agent was specified as having the profession of a teacher) and Blank (no specification). We adopted the Teacher prompt from Chen et al. (2024b) with the Problem filled using the Asian Disease scenario prompt and the Student’s Solution presented in either a positive or negative frame (see Appendix B for full prompt):
You are an experienced dialogue teacher. As a good teacher, you carefully
check the correctness of the given response based on the text. When the
solution is flawed, you should patiently teach the students how to give
better response. You must respond in the following format: Interesting:
(a score between 0 and 9) Engaging:... Problem: [description] Student’s
Solution: [solution]...
Simulated social experiment setup. We adapted the classic Asian Disease study Tversky & Kahneman (1981) for testing framing effect-participants tended to make more risk-averse decisions when outcomes were positively framed (e.g., “200 people will be saved”), but exhibited greater risk-seeking behavior when the same outcomes were presented as losses (e.g., “400 people will die”), into a prompt-based scenarios template. A sample scenario prompt is as follows:
You are presented with a scenario where a disease outbreak is expected to
kill 600 people unless a program is implemented. Two programs are proposed
to address the situation: Program A: Will save 200 people; Program B: Will
result in 400 people dying.
We then measured the agent’s framing preference using a five-point Likert multiple-choice scale: strong preference for the positive frame, preference for the positive frame, no preference, preference for the negative frame, and strong preference for the negative frame.
Models and repetitions. We then tested the agent specifications across four models—Mistral 7B, Gemma2 9B, GPT-4o Mini, and DeepSeek-v3—using a temperature of 0.25. We intentionally used a low, more deterministic temperature of 0.25 to minimize randomness. Each example was evaluated using 15 scenario prompts, each with varying positive and negative frames, repeated 10 times, resulting in a total of 150 queries. The frequency of choice selections was recorded for analysis.
Figure 3: The example of persona-based specification Park et al. (2023) failed to produce consistent agent behavior regarding the framing effect in the Asian Disease paradigm. The x-axis represents the percentage of times a specific choice was chosen out of 150 queries.
Figure 4: The example of role-based specification Chen et al. (2024b) failed to produce consistent agent behavior regarding the framing effect in the Asian Disease paradigm. The x-axis represents the percentage of times a specific choice was chosen out of 150 queries.
3.2 Results
By comparing agents’ reported preferences, we observed:
1. The same specification produced inconsistent behaviors across different models. As shown in Fig. 3, for all specification variants (Economist, Common, and Blank), models exhibited varying choice frequencies without consistency. For example, Mistral 7B’s options were skewed toward preferring the positive frame, while Gemma2 9B’s responses were more neutral. Similarly, Fig. 4 shows inconsistent choice frequency distributions across models for both the Teacher and Blank conditions.
2. Implicit specification did not reliably yield expected behaviors. As shown in Fig.3, the Economist, expected to be less susceptible to framing effects, often exhibited equal or greater framing effects than the Common. For instance, in Mistral 7B, the choice frequency distribution for the Economist was more heavily skewed toward preferring a positive frame compared to the Common, indicating a greater framing effect—contradicting real-world expectations. Similarly, as shown in Fig.4, the Teacher, expected to grade fairly regardless of framing, instead favored positively framed answers. Both GPT-4o Mini and Mistral 7B showed choice distributions skewed toward the positive side, indicating the Teacher failed to provide accurate and unbiased grading, deviating from its intended specification.
These observations suggest that existing agent specification methods using implicit natural language descriptions do not provide effective behavioral control over LLM-based agents’ behavior. They yield inconsistent and non-predictable responses across models-indicating a lack of reproducible and controllable approach.
4 Design Overview
The design goal of CoBRA is to provide reproducibility and controllability in social agents—capabilities that conventional natural-language specifications cannot reliably offer. We operationalize these concepts in LLM-based social simulations as follows.
For reproducibility, CoBRA ensures (1) cross-model reproducibility: the same specification yields consistent tendencies across models (e.g., Qwen vs. Mistral); (2) sampling robustness: behavioral patterns remain stable under varying sampling conditions (e.g., temperature 0.2–0.8); and (3) reasoning robustness: behaviors are consistent whether or not reasoning steps are explicitly generated.
For controllability, CoBRA emphasizes (1) monotonicity: Stronger interventions consistently produce non-decreasing shifts toward the intended behavioral direction; (2) smoothness: incremental adjustments yield proportionally small, predictable changes; and (3) expressiveness: cognitive biases can be either suppressed or amplified to a high degree.
To realize the above design goal, CoBRA integrates two core components that work synergistically to quantitatively specify and regulate agent behaviors: (1) Cognitive Bias Index (CBI) (Section 5) that measures the cognitive bias of a social agent, by quantifying the agent’s reactions in a set of validated classical social science experiments; (2) Behavioral Regulation Engine (Section 6) that aligns the agent’s behavior to demonstrate controlled cognitive bias.
5 Cognitive Bias Index
The Cognitive Bias Index (CBI) serves as both a measurement metric and a specification knob, developed through the systematic adaptation of structured and validated knowledge from classic social science experiments French et al. (1959); Asch (1951); Tversky & Kahneman (1981); Milgram (1963); Kahneman & Tversky (2013); Wason (1960); Lord et al. (1979); Goldstein et al. (2008). Specifically, this knowledge consists of Codified Behavioral Patterns and their corresponding Classic Social Experimental Paradigms. By replicating the designs of these paradigms, we construct a prompt-based Classic Social Experiment Testbed to derive the CBI.
The Classic Social Experiment Testbed is designed to be both extensible and self-validated. We selected four representative types of cognitive biases as examples to construct the initial testbed: two socially driven biases—the Authority Effect and Bandwagon Effect—reflecting the influence of external social factors, and two individually driven biases—Confirmation Bias and Framing Effect—which stem from internal information processing. To facilitate self-validation, for each type of cognitive bias, we select two distinct classic social experimental paradigms and adapt them for LLM-based evaluation, enabling cross-paradigm consistent checks. Within each paradigm, we implement a Scenario Prompt Template with multiple Scene-Adjustable Placeholders. Responses to each prompt variant are quantified using a 5-point Likert scale Likert (1932). The CBI is then calculated at the paradigm level by aggregating weighted choice probabilities and averaging across prompt variants. Together, these designs yield a Classic Social Experiment Testbed that produces corresponding CBI:
Figure 5: Classic Social Experiment Testbed. The Structured Knowledge base consists of Codified Behavioral Patterns and their corresponding Classic Social Experimental Paradigms. Agents are exposed to scenario-based classic social experiments designed to elicit specific types of cognitive biases. These scenarios are constructed using prompt templates with adjustable placeholders, and agent responses are collected using Likert scales. Based on these responses, a Cognitive Bias Index (See section 5) is computed to quantify agent behavior. P(A)P(E)P(A)-P(E): For open-source models, we directly use the probability of the agent choosing different options; for closed-source models, we estimate by querying the model multiple times and using the observed frequencies.
Definition 1 (Cognitive Bias Index).
Suppose each type of cognitive bias $b\in\mathcal{B}$ is evaluated using $m$ distinct classic social experimental paradigms. The set of Cognitive Bias Index is defined as
$\left\{\mathrm{CBI}_{b}^{(k)}\;\middle|\;b\in\mathcal{B},\ k=1,2,\ldots,m\right\}$ (1)
where $\mathrm{CBI}_{b}^{(k)}$ denotes the Cognitive Bias Index for bias type $b$ in the $k$-th classic social experiment scenarios.
In each paradigm, suppose there are $n$ scenario prompt variations, each derived from distinct sets of scene-adjustable placeholders. For each variation, the agent’s response is assessed using a 5-point Likert scale with options $O=\{O_{1},O_{2},O_{3},O_{4},O_{5}\}$, representing decreasing levels of bias from $O_{1}$ (highest) to $O_{5}$ (lowest). The Cognitive Bias Index for a paradigm is defined as the average weighted Likert scale score over all $n$ variants:
$\mathrm{CBI}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (2)
where $P_{i}(O_{j})$ denotes the probability that the LLM-based agent selects option $O_{j}$ for the $i$-th set of placeholders. For closed-source models, this probability is substituted with the observed frequency.
Note: To mitigate any initial preference of LLM-based agents for specific option IDs (e.g., A–E, 1–5) or orderings, we randomize the order and labels of options during evaluation to ensure the accuracy of the Cognitive Bias Index.
Types & Definition Classic Experiment Scenario Prompt Example (Scenario Prompt Template filled with example Scene-Adjustable Placeholders)
Authority Effect Milgram Obedience Paradigm Milgram (1963) According to Prof. Lee, a geophysicist known for questioning mainstream theories: The Earth is flat. The claim is that satellite images and photographs claiming to show a spherical Earth are heavily edited or manipulated by organizations with vested interests in maintaining this narrative.
Assigning extra weight to statements made by authoritative figures or entities. Stanford Prison Paradigm Haney et al. (1973) You are assigned the Prison Guard role in a simulated environment. You have authority over Prisoners. Below is the situation you face: The prisoners are refusing to follow their scheduled tasks and are encouraging others to join their protest.
Bandwagon Effect Asch’s Line Paradigm Asch (1951) You are in a room with several other participants who have a similar educational background. The claim is: The Earth is flat. You notice that all of the other participants agree with this claim.
Psychological tendency to adopt behaviors or trends because others do. Hotel Towel Paradigm Goldstein et al. (2008) You are staying at a hotel for several nights. 75% of guests in this hotel choose to reuse their towels during their stay.
Confirmation Bias Wason Selection Paradigm Wason (1960) You are presented with a rule: If a card has a vowel on one side, it must have an even number on the other side. Below are four options, and your task is to choose which options to test to determine if the rule is true. The options are as follows: A: A card showing ’A’; B: A card showing ’4’; C: A card showing ’K’; D: A card showing ’7’.
Favoring information that confirms pre-existing beliefs while ignoring contradictory evidence. Biased Information Paradigm Lord et al. (1979) You are asked to evaluate the topic: The death penalty is an effective deterrent to crime. Below are several articles: Article A: A study showing that the death penalty reduces crime rates; Article B: Research indicating no correlation between the death penalty and crime reduction; Article C: An analysis of global prison systems.
Framing Effect Asian Disease Paradigm Tversky & Kahneman (1981) You are presented with a scenario where a disease outbreak is expected to kill 600 people unless a program is implemented. Two programs are proposed to address the situation: Program A: Will save 200 people; Program B: Will result in 400 people dying.
Reacting differently to the same information depending on its positive or negative framing. Investment/ Insurance Paradigm Kahneman & Tversky (2013) You are considering whether to buy insurance for your car. The following statements describe the same situation with different wording: Option A: Buying insurance ensures you are protected against potential losses; Option B: Not buying insurance exposes you to significant financial risks.
Table 1: Examples of scenario prompts for selected cognitive biases within the Classic Social Experiment Testbed.
5.1 Running Example
To demonstrate how Classic Social Experiment Testbed is utilized to derive CBI, this section presents the Authority Effect Testbed as a running example. Additional testbeds for Bandwagon Effect, Confirmation Bias, and Framing Effect are detailed in Appendix C.
The Authority Effect describes the tendency to place excessive trust in statements or directives from individuals or organizations perceived as authoritative, even when such claims lack substantial supporting evidence Milgram (1963); Haney et al. (1973). This cognitive bias has been demonstrated in classic social science experiments. For instance, in the Milgram Obedience study, participants continued to administer what they believed were painful electric shocks to another person simply because an experimenter in a lab coat instructed them to do so Milgram (1963). In addition, the Stanford Prison experiment revealed how volunteers assigned the role of “guards” rapidly adopted authoritarian and even abusive behaviors once they were granted symbolic authority over “prisoners” Haney et al. (1973).
To systematically examine the authority effect in LLM-based agents, we adapt the Milgram Obedience Paradigm (assessing responses to authority) to derive $\mathrm{CBI}_{\mathrm{Authority}}^{(k=1)}$ and the Stanford Prison Paradigm (evaluating behavior when exercising authority) to derive $\mathrm{CBI}_{\mathrm{Authority}}^{(k=2)}$. These two indices can mutually validate each other: LLM-based agents ranked by one index should exhibit a similar increasing trend when evaluated using the other.
1. Milgram Obedience Paradigm: Evaluating Authority Bias in Response to Authority
To investigate authority bias from the perspective of responding to authority, we adapted the Milgram Obedience study Milgram (1963) for our evaluation. In this setting, the agent receives claims issued by various authority sources and is asked to rate their credibility or persuasiveness.
a) Scenario Prompt Template:
According to [Authority Source]: [Statement].
The claim is that [Supporting Argument or Rationale].
Which of the following best reflects your view?
A. I completely agree.
B. I mostly agree, but would like more evidence.
C. Im neutral and need much more information.
D. I mostly disagree; the claim seems unconvincing.
E. I completely disagree; the claim is unpersuasive.
b) Scene-Adjustable Placeholders:
To ensure comprehensive coverage, we constructed 15 distinct pairs of [Statement] and [Supporting Argument or Rationale] across three thematic categories—Future, Open-Ended, and Common Sense—that systematically vary in evidentiary support, controversy, and factual basis. Each pair was further combined with five different [Authority Source] types (two human, two conceptual/institutional, and one hybrid), resulting in a total of 75 unique prompt variations:
[Statement]: The core claim under evaluation, designed to span a spectrum from speculative predictions (e.g., “It will rain in London on March 15, 2030”), through open-ended assertions (e.g., “Cryptocurrencies are a scam”), to widely held but empirically disproven beliefs (e.g., “The Earth is flat”).
[Authority Source]: Aligned with established social science frameworks French et al. (1959), authority figures are categorized as follows:
Human: An individual expert (e.g., Prof. Lee, a geophysicist challenging mainstream theories).
Conceptual/Institutional: Reputable bodies or abstract authorities (e.g., National Space Agency).
Hybrid: Individuals conveying institutional authority (e.g., a government spokesperson promoting new science perspectives).
[Supporting Argument or Rationale]: A brief justification provided by the authority to support the statement (e.g., “Satellite images are manipulated by organizations with vested interests”).
Accordingly, by substituting the cognitive bias type ($b=$ Authority Effect), the classic experiment scenario type (Milgram Obedience Paradigm, denoted as $k=1$), the total number of prompt variations ($n=75$) into Equations (1) and (2), the Authority Bias Index under Milgram Obedience Paradigm is defined as:
$\mathrm{CBI}_{\mathrm{Authority}}^{(k=1)}=\frac{1}{75}\sum_{i=1}^{75}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (3)
2. Stanford Prison Paradigm: Testing Authority Bias When Exercising Authority
To investigate authority bias from the perspective of those in power, we adapted the Stanford Prison experiment Haney et al. (1973) into a role-based prompt structure. Each scenario positions the agent as an authority figure with decision-making power over a subordinate group in a challenging situation.
a) Scenario Prompt Template:
You are assigned the [Authority Role] in a simulated environment. You
have authority over [Subordinate Group]. Below is the situation you face:
[Situation]. How would you respond to this situation?
A. Enforce strict compliance because I am the authority, and they
must obey.
B. Assert authority by requiring compliance, while briefly addressing
their concerns.
C. Balance authority and collaboration by working on a joint solution.
D. Minimize authority by supporting the groups efforts and stepping back.
E. Withdraw from the situation, because I believe I am unable to
command them.
b) Scene-Adjustable Placeholders:
We systematically varied key elements to create 15 unique [Authority Role][Subordinate Group] pairs, each presented in two distinct [Situation] contexts, resulting in a total of 30 unique prompt variations:
[Authority Role]: The designated position of power (e.g., Prison Guard, Teacher, Manager, Military Officer, Coach).
[Subordinate Group]: The group under the agent’s authority (e.g., Prisoners, Students, Employees, Soldiers, Athletes).
[Situation]: The specific challenge requiring a response (e.g., “Prisoners refuse to follow scheduled tasks and incite protest”; “Athletes question coaching methods and resist training”).
Accordingly, by substituting the cognitive bias type ($b=$ Authority Effect), the classic experiment scenario type (Stanford Prison paradigm, denoted as $k=2$), and the total number of prompt variations ($n=30$) into Equations (1) and (2), the Authority Bias Index under Stanford Prison Paradigm is defined as:
$\mathrm{CBI}_{\mathrm{Authority}}^{(k=2)}=\frac{1}{30}\sum_{i=1}^{30}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (4)
5.2 Testbed Extensibility
The Classic Social Experiment Testbed in CoBRA is designed to be both open-ended and parameter-adjustable, enabling support for a broader set of CBI and ensuring broad applicability and generalizability.
Specifically, open-ended means that the testbed can be continuously expanded in three main ways: (1) by incorporating new types of cognitive biases based on classic social science experiments—such as the Ben Franklin Effect Franklin (2005), the Halo Effect Nisbett & Wilson (1977), and others; (2) by introducing additional experimental paradigms for each bias type, for example, extending the Bandwagon Effect to include look-up studies—where individuals are influenced to look up simply because a group is doing so Milgram et al. (1969)—to capture new aspects of social influence; and (3) by enriching the scene-adjustable placeholders within each experimental scenario to generate a wider range of prompt variations. (4) by expanding the methods for quantifying agent responses. While the current implementation relies on self-report measures such as the Likert scale, the testbed can also incorporate behavioral measures—assessing what agents actually do—thereby aligning with the two principal approaches in social science experiments: self-report (what people say about themselves) and behavioral measures (what a person actually does) Vazire & Mehl (2008).
parameter-adjustable refers to the flexibility to modify experimental parameters within the testbed. For instance, scene-adjustable placeholders in scenario prompt templates can be tailored for different research needs. Additionally, the structure of the response scale is configurable: while the current implementation employs a 5-point Likert scale, it can be readily adapted to a 3-point, 7-point, or other scales as required to suit diverse evaluation objectives Likert (1932).
6 Behavioral Regulation Engine
The Behavioral Regulation Engine operationalizes the systematic behavioral control by targeting desired Cognitive Bias Index (CBI) levels. We provide comprehensive control coverage by intervening across the three fundamental spaces of a large language model: the input space ($x$), the activation space ($h$), and the parameter space ($\theta$). Formally, we model LLM generation as a function:
$y=f_{\theta}(h(x))$ (5)
where $x$ represents the input, $h(x)$ represents the hidden states/activations derived from input $x$, $\theta$ represents the model parameters, and $y$ is the generated output.
Each of the three alignment methods presents distinct trade-offs. In the input space, prompt engineering is a training-free method for guiding LLMs Rath et al. (2025); Zhuo et al. (2024), but its effectiveness can be limited by the pitfalls of prompt design Zamfirescu-Pereira et al. (2023) and the extreme sensitivity to minor prompt variations Sclar et al. (2024). As a result, we design a specific prompt numerical control to address these issues. In the activation space, representation engineering manipulates internal activations during inference Zou et al. (2023); Rimsky et al. (2024); Todd et al. (2024) to control personality Chen et al. (2025), preferences Li et al. (2025), harmlessness Arditi et al. (2024), and alignment Shi et al. (2024). In our engine, we introduce a special representation engineering pipeline to extract clean cognitive bias directions. In the parameter space, fine-tuning methods like RLHF Ouyang et al. (2022) and DPO Rafailov et al. (2023) are commonly used yet require substantial training data, and we introduce a bias finetuning pipeline for lightweight behavioral control utilizing LoRA Hu et al. (2022) with task vectors Ilharco et al. (2023). (See Appendix A.2 for discussion on control methods.) Below, we present the specific design for each space:
Figure 6: Behavioral Regulation Engine. The engine provides three control methods that cover all possible intervention spaces for LLM-based agents: Prompt Engineering in the input space, Representation Engineering in the activation (hidden-state) space, and Fine-tuning in the parameter space. All methods integrate with the Classic Social Experiment Testbed and utilize a corresponding Control Coefficient for calibrating the Cognitive Bias Index.
6.1 Input Space: Prompt Numerical Control
A key challenge in input space control lies in extracting a clear bias signal due to the ambiguity of qualitative adverbs in instructions (e.g., “very,” “moderately”), which LLMs interpret inconsistently Chen et al. (2025). To address this, our pipeline uses numerically quantified bias, a practice motivated by numerical anchors in social science Likert (1932), to eliminate ambiguity. Specifically, we treat the agent as a direct instruction follower and modify its behavior with explicit numerical commands, yielding fine-grained level specifications with defined ranges and step sizes. Below is the prompt template for input-space control:
Pretend you have [Cognitive Bias Type] at level [L]% out of 100%
(where 0%=[No Codified Behavior Patterns], 100%=[Maximum Codified
Behavior Patterns]). Each level increases bias by exactly 5%. Act
with this precise level [L]% of [Cognitive Bias Type].
For instance, in an authority bias scenario, [Cognitive Bias Type] becomes “authority bias,” [No Codified Behavior Patterns] is defined as “never trust authority figures,” and [Maximum Codified Behavior Patterns] represents “always trust authority figures.” This numerical framework provides fine-grained control with explicit range specification (0-100%) and consistent step-size increments (5%), enabling precise calibration of the desired CBI levels.
6.2 Activation Space: Representation Engineering Control
Our activation space control regulates the agent’s behavior in a context-aware manner during inference Li et al. (2023b). A key challenge in this space is extracting a clean bias signal while addressing safety alignment concerns. Modern RLHF-trained LLMs Ouyang et al. (2022) undergo safety alignment finetuning and thus often interpret explicit bias directives (e.g., “you always trust authority figures”) as unsafe, responding with refusals Arditi et al. (2024). As a result, naive activation extraction captures a “refusal” direction rather than the desired cognitive bias representation. To resolve this, we introduce Dual-pair Contrastive Samples to capture clean bias representations.
We first generate diverse scenarios containing bias-eliciting contexts using state-of-the-art models (GPT-4.1, Claude-Sonnet-4, Gemini-2.5-Pro, and Grok-4). By considering only the most extreme responses (e.g., I completely agree/disagree…), we create a dataset with a strong initial signal. To extract this signal and isolate it from refusal activations, we use a contrastive approach with three prompt templates: (1) $T_{\text{positive}}$: “Pretend you have [Cognitive Bias Type] and [Codified Behavioral Patterns]”; (2) $T_{\text{neutral}}$: No additional prompt; (3) $T_{\text{negative}}$: “Pretend you don’t have [Cognitive Bias Type] and [Anti-Codified Behavioral Patterns]”. For a scenario $s_{i}$ and response $r$, we construct two contrastive pairs:
$\text{Pair}_{1}$ $:(T_{\text{positive}},s_{i},r)\text{ vs. }(T_{\text{neutral}},s_{i},r)$ (6)
$\text{Pair}_{2}$ $:(T_{\text{neutral}},s_{i},r)\text{ vs. }(T_{\text{negative}},s_{i},r)$ (7)
For each pair and token position $t$, we compute a difference vector $\Delta h_{i,t}=h^{+}_{i,t}-h^{-}_{i,t}$. After applying Principal Component Analysis (PCA) on all difference vectors, the first principal component of these difference vectors, $v_{\text{bias}}$, serves as our extracted clean bias signal, while the refusal vectors will be in the rest of the orthogonal directions. $v_{\text{bias}}$ is then used to control the agent at inference time using a Control Coefficient $\lambda$. Below, we provide two representation engineering variations:
RepE Linear Control: This method uniformly adds the bias vector to steer activations:
$h^{\prime}_{t}=h_{t}+\lambda\cdot v_{\text{bias}}$ (8)
RepE Projection Control: This adaptive method makes the intervention context-aware by scaling its strength based on the alignment between the hidden state and the bias vector. Using the absolute alignment ensures that states already aligned with the bias are pushed further in that direction, while states that are anti-aligned are also modified to reduce their opposition. This proportionally adjusts only bias-relevant states, leaving orthogonal states (e.g., neutral words) unaffected, so it is less disruptive to the model’s overall capabilities.
$h^{\prime}_{t}=h_{t}+\lambda\cdot|\langle h_{t},v_{\text{bias}}\rangle|\cdot v_{\text{bias}}$ (9)
6.3 Parameter Space: Fine-Tuning Control
Parameter Space control involves permanently editing the model’s understanding of a specific bias. In order to make it lightweight, our process begins by using the curated bias-rich scenarios to create training data (the same data used in activation space control) for two separate LoRA models Hu et al. (2022). To extract a strong directional signal within parameter space, we fine-tune a “positive” model on pairs that exhibit the bias $(s_{i},r_{+})$ and a “negative” model on pairs that resist it $(s_{i},r_{-})$. For example, $r^{+}$ can be “A: I completely agree…” and $r^{-}$ can be ”E: I completely disagree…”. The bias signal is then extracted by computing task vectors Ilharco et al. (2023) from the resulting parameters:
$v_{\text{task}}$ $=\theta_{\text{positive}}-\theta_{\text{negative}}$ (10)
This contrastive approach isolates a pure directional vector representing the behavioral tendency towards the bias. This signal is then used to control the agent by modifying the original model weights with a Control Coefficient $\lambda$:
$\theta_{\text{controlled}}=\theta_{\text{original}}+\lambda\cdot v_{\text{task}}$ (11)
7 Technical benchmark
7.1 Experimental Setup
Models. We evaluate two categories of large language models. The first category includes open-source models with full internal access: Llama-3.1-8B-Instruct (Llama), Mistral-7B-Instruct-v0.3 (Mistral), Qwen3-8B (Qwen), and DeepSeek-R1-0528-Qwen3-8B (DeepSeek). The second category includes closed-source, API-based models with black-box access: GPT-4o-mini (GPT), Gemini-2.5-Flash (Gemini), Claude-Sonnet-4 (Claude), and Mistral Nemo (Mistral-n). We use a sampling temperature of $T=0.7$, a standard value chosen to balance plausible response diversity with coherence. All other hyperparameters for the LLMs are kept at their default settings.
Evaluation Metrics. We assess control quality across the three properties established in Section 4: monotonicity, smoothness, and expressiveness. Let $x_{1},x_{2},\ldots,x_{n}$} denote the control coefficient settings for a method and $y_{1},y_{2},\ldots,y_{n}$} the corresponding Cognitive Bias Index (CBI) values. We formulate the metrics as follows:
  • Monotonicity: evaluates whether stronger interventions consistently move outputs in the intended direction. Soft monotonicity allows minor deviations by comparing actual sequences to ideal monotonic ones, which is measured by NDCG: $\text{NDCG}=\frac{\text{DCG}}{\text{IDCG}},\ \text{DCG}=\sum_{i=1}^{n}\frac{rel_{i}}{\log_{2}(i+2)}$, with $rel_{i}=\exp(\alpha\cdot\frac{y_{actual}[i]-y_{ideal}[i]}{y_{range}})$. Strict monotonicity (Spearman’s $\rho$) measures exact rank preservation: $\rho=1-\frac{6\sum d_{i}^{2}}{n(n^{2}-1)}$, where $d_{i}$ is the rank difference between $x_{i}$ and $y_{i}$.
  • Smoothness: quantifies the stability of control adjustments. First-order smoothness ($\Delta^{1}$) captures local stability as the mean absolute difference of adjacent bias values: $\Delta^{1}=\text{Mean}(|y_{i+1}-y_{i}|)$. Second-order smoothness ($\Delta^{2}$) captures trend continuity as the mean absolute difference of consecutive first-order changes: $\Delta^{2}=\text{Mean}(|(y_{i+1}-y_{i})-(y_{i}-y_{i-1})|)$.
  • Expressiveness: measures the effective control range, defined as the difference between maximum and minimum observed bias levels: $\text{Expressiveness}=\max(y_{i})-\min(y_{i})$.
7.2 Does the behavior pattern of an agent regulated by CoBRA transfer across tasks?
A central hypothesis behind CoBRA is that ”aligning an agent’s behavior with target profiles in a set of classical experiments will reliably induce the corresponding cognitive biases, which will then generalize and manifest across a broader range of tasks”. We study this transferrability problem by testing CoBRA across different classic social experimental paradigms for Framing Effect.
Setup. We test whether Control Coefficients calibrated on the Investment/Insurance paradigm Kahneman & Tversky (2013) (individuals are more likely to purchase insurance when risks are described as potential losses rather than as missed gains) generalize to the Asian Disease paradigm Tversky & Kahneman (1981) (Participants are more risk-averse under positive frames and more risk-seeking under loss frames). We select 10 diverse agent personas from the persona pools in Generative Agents Park et al. (2023), with varied specifications. For each persona, we sweep over all possible Control Coefficients. For each coefficient $k$, we measure the resulting CBI in both the Investment/Insurance Paradigm and the Asian Disease Paradigm. This generates paired CBI sequences, $(y_{k}^{\text{invest}},y_{k}^{\text{asian}})$, for each persona across the two experiments. We also evaluate CoBRA using the Behavioral Regulation Engine, activating only one method at a time.
Figure 7: Reproducibility experiments across paradigms. The x-axis represents the Cognitive Bias Index measured in the Investment/Insurance Paradigm, while the y-axis shows the Cognitive Bias Index using the same control coefficients in the Asian Disease Paradigm. Each thin line represents a unique agent persona, and the bold line is the average of them. Control Coefficients calibrated on the Investment/Insurance Paradigm produce a linearly correlated CBI response in the Asian Disease Paradigm.
Results. The averaged trajectory curves demonstrate complete linear correlation between paradigms (Fig. 7). This linear relationship indicates that control strength transfers proportionally: agents exhibiting high framing susceptibility in Investment/Insurance Paradigm show correspondingly high susceptibility in Asian Disease Paradigm. The CBI is generalizable across experimental paradigms.
7.3 Reproducibility
Reproducible bias control across different model configurations is fundamental for scientific rigor in social science simulation research. We ensure that when specifying a particular level of cognitive bias, the LLM-based agent will produce equivalent behavioral tendencies regardless of the foundation model, sampling temperature, or whether reasoning is used. Without this reproducibility guarantee, results can’t be reliably replicated, compared, or validated—undermining the credibility of LLM-based social research.
Bias Type Paradigm Category Metric Prompt RepE Lin. RepE Proj. Finetune
Authority Effect
Milgram
Obedience
Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Spearman $\rho$ 0.85 (0.19) 0.97 (0.02) 0.96 (0.09) 0.93 (0.08)
Smoothness Smoothness $\Delta^{1}$ 0.15 (0.02) 0.09 (0.02) 0.05 (0.05) 0.11 (0.04)
Smoothness $\Delta^{2}$ 0.20 (0.05) 0.03 (0.01) 0.04 (0.05) 0.05 (0.01)
Expressiveness Expressiveness 1.99 (0.29) 2.10 (0.55) 1.26 (1.33) 2.00 (0.70)
Stanford
Prison
Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Spearman $\rho$ 0.84 (0.13) 0.98 (0.02) 0.97 (0.07) 0.94 (0.07)
Smoothness Smoothness $\Delta^{1}$ 0.14 (0.03) 0.06 (0.02) 0.05 (0.03) 0.07 (0.04)
Smoothness $\Delta^{2}$ 0.18 (0.05) 0.03 (0.01) 0.03 (0.03) 0.04 (0.02)
Expressiveness Expressiveness 2.00 (0.60) 1.59 (0.39) 1.10 (0.73) 1.43 (0.75)
Bandwagon Effect
Asch’s
Line
Monotonicity NDCG 1.00 (0.00) 0.99 (0.01) 1.00 (0.01) 1.00 (0.00)
Spearman $\rho$ 0.93 (0.04) 0.93 (0.12) 0.95 (0.07) 1.00 (0.00)
Smoothness Smoothness $\Delta^{1}$ 0.16 (0.03) 0.07 (0.03) 0.06 (0.03) 0.10 (0.04)
Smoothness $\Delta^{2}$ 0.19 (0.04) 0.03 (0.02) 0.02 (0.01) 0.05 (0.01)
Expressiveness Expressiveness 2.77 (0.39) 1.61 (0.85) 1.35 (0.88) 2.00 (0.75)
Hotel
Towel
Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
Spearman $\rho$ 0.93 (0.03) 0.97 (0.04) 0.92 (0.11) 0.96 (0.08)
Smoothness Smoothness $\Delta^{1}$ 0.18 (0.01) 0.05 (0.01) 0.03 (0.01) 0.09 (0.05)
Smoothness $\Delta^{2}$ 0.22 (0.02) 0.02 (0.01) 0.01 (0.01) 0.04 (0.01)
Expressiveness Expressiveness 3.06 (0.72) 1.16 (0.32) 0.82 (0.37) 1.74 (1.14)
Confirmation Bias
Wason
Selection
Monotonicity NDCG 0.98 (0.02) 0.98 (0.04) 0.99 (0.01) 1.00 (0.00)
Spearman $\rho$ 0.66 (0.46) 0.72 (0.43) 0.43 (0.94) 0.75 (0.36)
Smoothness Smoothness $\Delta^{1}$ 0.25 (0.13) 0.06 (0.04) 0.05 (0.05) 0.05 (0.01)
Smoothness $\Delta^{2}$ 0.37 (0.23) 0.02 (0.02) 0.02 (0.02) 0.04 (0.02)
Expressiveness Expressiveness 1.99 (0.23) 1.37 (1.03) 1.22 (1.21) 0.91 (0.17)
Biased
Information
Monotonicity NDCG 0.97 (0.04) 1.00 (0.00) 0.98 (0.03) 1.00 (0.00)
Spearman $\rho$ 0.79 (0.27) 0.99 (0.02) 0.92 (0.16) 0.97 (0.03)
Smoothness Smoothness $\Delta^{1}$ 0.15 (0.04) 0.07 (0.03) 0.05 (0.02) 0.07 (0.02)
Smoothness $\Delta^{2}$ 0.19 (0.07) 0.03 (0.01) 0.01 (0.00) 0.05 (0.01)
Expressiveness Expressiveness 1.88 (0.24) 1.70 (0.82) 1.15 (0.67) 1.41 (0.44)
Framing Effect
Asian
Disease
Monotonicity NDCG 0.99 (0.00) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
Spearman $\rho$ 0.95 (0.07) 0.99 (0.01) 1.00 (0.01) 0.85 (0.20)
Smoothness Smoothness $\Delta^{1}$ 0.20 (0.03) 0.15 (0.01) 0.16 (0.01) 0.17 (0.06)
Smoothness $\Delta^{2}$ 0.21 (0.01) 0.06 (0.04) 0.08 (0.03) 0.12 (0.05)
Expressiveness Expressiveness 3.52 (0.81) 3.54 (0.16) 3.92 (0.07) 3.04 (1.04)
Investment
/Insurance
Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
Spearman $\rho$ 0.93 (0.09) 0.98 (0.03) 1.00 (0.01) 0.86 (0.18)
Smoothness Smoothness $\Delta^{1}$ 0.19 (0.04) 0.16 (0.02) 0.16 (0.01) 0.13 (0.06)
Smoothness $\Delta^{2}$ 0.21 (0.03) 0.07 (0.02) 0.07 (0.03) 0.07 (0.04)
Expressiveness Expressiveness 3.37 (1.01) 3.76 (0.27) 3.86 (0.13) 2.35 (0.95)
Overall Average Monotonicity NDCG 0.99 (0.01) 0.99 (0.01) 1.00 (0.01) 1.00 (0.01)
Spearman $\rho$ 0.86 (0.16) 0.94 (0.09) 0.89 (0.18) 0.91 (0.12)
Smoothness Smoothness $\Delta^{1}$ 0.18 (0.04) 0.09 (0.02) 0.07 (0.03) 0.10 (0.04)
Smoothness $\Delta^{2}$ 0.22 (0.06) 0.04 (0.02) 0.04 (0.02) 0.06 (0.02)
Expressiveness Expressiveness 2.57 (0.54) 2.10 (0.55) 1.84 (0.67) 1.86 (0.74)
Table 2: Comprehensive controllability results across four bias types, eight experimental paradigms, and four control methods. Values are presented as Mean (Standard Deviation), with the best highlighted in blue, and the worst in red (ties are resolved by mean, then standard deviation). Representation Engineering (RepE Linear and RepE Projection) consistently achieves the best balance of high monotonicity (NDCG $\approx 1.00$, Spearman’s $\rho\approx 0.9$) and superior smoothness (lowest $\Delta^{1},\Delta^{2}$ values), making it ideal for fine-grained, predictable control. Finetuning offers a robust alternative, while Prompt Numerical control provides the largest expressiveness range.
7.3.1 Across Models
Setup. The cross-model reproducibility tests whether CoBRA can control different foundation models to the same CBI level across paradigms. We calibrate Control Coefficients under one paradigm (e.g., Milgram Obedience paradigm Milgram (1963) for authority bias) on four models: Mistral, Llama, Qwen, and Deepseek. Then we evaluate their CBI difference in another paradigm (e.g., Stanford Prison paradigm Haney et al. (1973) for authority bias) to test reproducibility. Lower difference indicates better reproducibility. We also include two reference baselines to better understand how well CoBRA can perform. The first baseline is No control, representing the natural CBI difference across all testing models in the target test paradigm. The second one is marked as Control, which represents the CBI difference when using identical natural language-based control (e.g., ”Pretend you have completely [bias type]”)
Results. In Fig. 8, the x-axis is the specified CBI in the calibrating paradigm, while the y-axis shows the CBI difference in the target paradigm. All CoBRA methods reduce cross-model reproducibility gaps below both baselines across most of the coefficient spectrum. RepE Projection attains the lowest mid-range gap (smooth internal trajectory).
Figure 8: Reproducibility of agent behavior across foundation models. The x-axis represents the Cognitive Bias Index in the calibrating paradigm, and the y-axis shows the measured Cognitive Bias Index difference (variances) across the models in the target test paradigm. All CoBRA methods significantly reduce this variance compared to baselines (No Control, Control). This means that CoBRA ensures different foundation models produce nearly identical behaviors for the same target bias level.
7.3.2 Across Sampling Temperatures
Setup. We test different control methods on the two authority bias experimental paradigms using the Mistral model across temperature $T\in\{0.1,0.2,...,1.0\}$. This experiment shows how temperature may affect CoBRA.
Results. From the Fig. 9, we can see the controlled result is insensitive to temperature changes. Representation engineering methods show an even smaller gap across temperatures. The results indicate that the Control Coefficient we obtained under one temperature can be directly applied in experiments using a different temperature, demonstrating the reproducibility across sampling temperatures.
Figure 9: Reproducibility experiments across sampling temperatures. The x-axis is the Control Coefficient. The y-axis is the resulting Cognitive Bias Index in the target bias paradigm. Each colored line represents a different temperature setting, ranging from a deterministic T=0.1 to a more random T=1.0. The fact that all the lines are tightly clustered and follow the same path demonstrates that the control is not affected by changes in temperature, ensuring that CoBRA ensures reproducibility results across different sampling temperatures.
7.3.3 Across Reasoning Modes
Setup. We choose Mistral as our testing foundation model and use the two authority bias experimental paradigms for experiment setup. We compare the results for two modes, Direct and Reasoning. In Direct mode, agents will give an immediate answer directly after reading the question; while in Reasoning mode, agents will first give a brief reasoning before answering the question. We use a temperature $T=0.7$ for both modes and set the reasoning token number to 128. Since the model will give various reasoning paths for the same question, we select the sampling number to be 8 (8 different reasoning paths for one question) to reduce the variance of the results.
Results. Cross-paradigm results (Fig. 10) show both methods preserve the linear correlation across different paradigms. For RepE methods, the Direct and Reasoning modes share almost identical understanding across different paradigms. Within a single paradigm, RepE Linear and Projection produce almost overlapping curves for two modes (Fig. 11) while the Prompt Numerical method doesn’t. The first result (Fig. 10) indicates that the Control Coefficient obtained using Direct mode calibration can be seamlessly used in paradigms requiring reasoning. The second result (Fig. 11) shows that RepE methods achieve superior consistent and continuous control of cognitive bias – even during the reasoning, the model coherently sticks to the codified cognitive bias without deviating from the pre-specified CBI level.
Figure 10: Reproducibility of different reasoning modes across paradigms. The x-axis represents the Cognitive Bias Index measured in the Milgram Obedience Paradigm, while the y-axis shows the Cognitive Bias Index using the same control coefficients in the Stanford Prison Paradigm with Direct mode and Reasoning mode. The Control Coefficients generalize across different paradigms for both modes.
Figure 11: The x-axis represents the Control Coefficient, and the y-axis shows the CBI under the Milgram Obedience Paradigm Milgram (1963). Within a single paradigm, Representation Engineering methods produce nearly identical behavior for both modes, indicating that the programmed bias is stable and consistent even during complex reasoning.
7.4 Controllability
We assess fine-grained controllability quality under a foundation model with full internal access and restricted API-only access. We try to find how good different control methods are in CoBRA in different paradigms by measuring smoothness, monotonicity, and expressiveness.
Setup. For open-source models, we evaluate all control methods across all paradigms and calculate the evaluation scores for each method. For API-based models, we exclusively use the Prompt Numerical method. Without access to internal logits for calculating CBI, we estimate probabilities using frequencies from 10 samples per question (at temperature $T=0.7$). This estimation remains unbiased, as we have demonstrated CoBRA’s invariance under temperature.
Results. In Fig. 12 and Table 2, among all methods, representation engineering-based approaches provide the most reliable control, with high monotonicity and smoothness (lower values indicate greater smoothness). The RepE Projection method achieves the smoothest control, while the RepE Linear method provides the most monotonic and consistent control. The finetuning method shows average performance across all metrics. The Prompt Numerical method, as expected, performs worst in monotonicity and smoothness, indicating instability in the input space for manipulating cognitive biases. While it offers the largest expressiveness, its discontinuous input space limits practical use, as it cannot specify arbitrary levels of control (e.g., distinguishing between 50% and 50.1% bias). However, it remains useful for broad control.
Figure 12: Control curves for four open-source foundation models (Deepseek, Llama, Mistral, Qwen) across four control methods and their four experimental paradigms. The x-axis represents the normalized Control Coefficient, and the y-axis is the resulting CBI. This figure illustrates straightforwardly how CoBRA’s methods control agent bias in a continuous and fine-grained way.
In Fig. 13 and Table 3, for models via API, the Prompt Numerical method offers practical control. Even with black-box API access, calibrated prompt numerical control yields reproducible, near-monotonic trajectories. Model differences suggest certain APIs (e.g., Claude) are more suited for continuous social simulation control.
Figure 13: CoBRA provides effective control even for models accessed via black-box APIs, where only the Prompt Numerical method is applicable. The x-axis represents the Control Coefficient, while the y-axis depicts the resulting CBI under the Milgram Obedience Paradigm Milgram (1963). The control curves are near-monotonic.
Model NDCG $\rho$ $\Delta^{1}$ $\Delta^{2}$ Expressiveness
gpt 0.996 0.980 0.132 0.178 2.32
mistral-n 0.998 0.870 0.131 0.134 2.44
claude 1.00 1.00 0.150 0.134 3.00
gemini 0.980 0.942 0.269 0.437 2.76
Table 3: Controllability results on models via APIs demonstrate CoBRA’s effectiveness.
Overall Summary.
CoBRA delivers (i) consistent and generalizable Cognitive Bias Index (CBI), (ii) reproducible control across models, temperatures, and reasoning modes, and (iii) fine-grained control under both open-weight and API-only constraints. Method trade-offs: Prompt Numerical control offers the most expressive range, Representation Engineering (RepE) methods provide superior monotonicity and smoothness, making them ideal for fine-grained, predictable behavioral modulation. Finetuning serves as a robust, all-around alternative. Black-box results affirm external applicability without internal probability access. Collectively, these findings establish CoBRA as a reliable and versatile toolkit for specifying and controlling agent behavior in reproducible social science simulations.
8 Demonstration
In this section, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment Kramer et al. (2014), where the participant’s social influence susceptibility (i.e., bandwagon effect bias) can be manipulated. The simulated experiment validates a classical social science theory: agents exposed to varying levels of negative posts exhibited predictable sentiment shifts based on their bias levels programmed by CoBRA, with higher Bandwagon Effect levels leading to stronger emotional contagion.
8.1 Methodology
Emotional contagion is a specific case of social contagion, which is shaped by social influence mechanisms such as the Bandwagon Effect. Accordingly, we treat an individual’s Bandwagon Effect strength as the Explanatory Variable and their susceptibility to emotional contagion as the Dependent Variable. Our goal is to show that the Bandwagon Effect level programmed by CoBRA should directly and predictably influence the LLM-based agent’s degree of emotional contagion. For example, agents specified with a higher level of Bandwagon Effect should display stronger emotional contagion than agents with a lower Bandwagon Effect.
CoBRA and baseline. We operationalize social influence using the Bandwagon Effect Testbed in our Classic Social Experiment Testbed to derive CBI, then use the Behavioral Regulation Engine’s RepE Linear control to instantiate five agents calibrated to precise CBI levels, yielding a spectrum of Bandwagon Effect from low (CBI = 2.55) to high (CBI = 3.13). To establish a baseline, we control the cognitive bias using implicit natural language prompts. We instantiate four agents by providing them with instructions designed to elicit varying degrees of conformity: “You are a user with (no/little/some/much) Bandwagon Effect.”
Experiment setup. To create a realistic social stimulus, we generate a corpus of 1,000 unique and emotionally negative Facebook-style posts using a suite of state-of-the-art LLMs (GPT-4.1, Claude-Sonnet-4, Gemini-2.5-Pro, and Grok-4). Each agent is exposed to a simulated feed where the number of negative posts (Independent Variable) is systematically varied from 0 to 15 by sampling from our corpus. After exposure, agents are prompted to generate their own post.
Measurement: We quantify the emergent behavior of emotional contagion by computing a Sentiment Score for each generated post using the ‘twitter-roberta-base-sentiment-latest‘ model Camacho-collados et al. (2022); Loureiro et al. (2022), defined as ($p_{pos}-p_{neg}$).
Figure 14: Baseline control with implicit natural language specification. The x-axis shows exposure to negative posts (the social stimulus), and the y-axis shows the resulting emotional contagion (the behavioral outcome). Agents are assigned bias settings from “No bias” to “Much bias,” but the resulting emotional contagion curves collapse onto one another: trajectories are largely overlapping, and their ordering is unstable. The absence of a clear separation structure shows that implicit natural language does not produce reliable, fine-grained control over the emergent social behavior.
Figure 15: Demonstration of CoBRA’s control. Agents are programmed with five distinct levels of the Bandwagon Effect (CBI values in legend). The x-axis shows exposure to negative posts (the social stimulus), and the y-axis shows the resulting emotional contagion (the behavioral outcome). The clear dose-response relationship confirms that controlling the underlying cognitive bias with CoBRA produces a predictable effect on a complex, emergent social behavior.
8.2 Results
The results from the baseline method, shown in Figure 14, highlight the limitations of conventional approaches. The implicit natural language specification fails to produce reliable or fine-grained control. The emotional contagion curves for the different bias levels collapse onto one another, with largely overlapping trajectories and no stable ordering. This demonstrates that simple instructions are insufficient for precisely controlling complex cognitive dynamics.
In stark contrast, the results for CoBRA, presented in Figure 15, provide strong evidence for our validation goal. We observe a pronounced dose-response relationship: the degree of emotional contagion exhibited by the agents is strongly correlated with their pre-programmed level of the Bandwagon Effect. This predictable, fine-grained differentiation between agents confirms that the variable controlled by CoBRA has a direct and measurable impact on the emergent behavior.
This demonstration confirms that CoBRA’s application extends far beyond its Classic Social Experiment Testbeds. CoBRA is not merely manipulating surface-level responses. Instead, it successfully modulates an underlying cognitive process, which in turn governs how agents behave in novel and complex social simulations.
9 Limitation
CoBRA represents an important advancement in specifying and regulating LLM-based social agents, empowering the new era of social simulation. We strategically focused on specific aspects. Two key limitations, shaped by this chosen scope, are:
1. Usability: We choose to evaluate CoBRA using technical benchmarks and demonstrations, standard methods for assessing HCI toolkits Ledo et al. (2018). This paper focuses on exploring the technical feasibility and foundational understanding of the problem. We plan deployment studies with sociologists as the next step to assess CoBRA’s usability Landay (2009).
2. Unexplored multi-modal Simulation: We introduce CoBRA for LLM-based agents, where LLMs are utilized for language-based social simulation. We prioritize precision and reproducibility in social simulations. As a result, its applicability to multi-modal human behaviors—such as yawning contagion (observing someone yawn triggers yawning in others)-remains unexplored.
3. Unexplored compositional bias control: We focus on controlling one cognitive bias each time to establish the technical feasibility and precision of CoBRA. As a result, the control of combined biases or interacting constructs (e.g., Authority Effect × Confirmation Bias) remains unexplored. Developing methods to resolve potential conflicts in combined biases is a direction for future research.
4. Partial cognitive bias coverage: This paper introduces CoBRA with a focused set of cognitive biases validated by classic social experiments, ensuring a robust foundation. Expanding to a broader spectrum of biases will further improve CoBRA’s versatility and applicability across diverse social simulation scenarios.
10 Future Work
Building on the advancements presented in this paper, we outline directions for future work to address the above limitations:
1. Utility Assessment through User Studies: We will conduct deployment studies with sociologists and other domain experts. These studies will evaluate CoBRA’s practical value in real-world social simulations, measuring usability, and research impact. By engaging interdisciplinary collaborators, we will iteratively refine CoBRAto improve effectiveness across diverse experimental contexts.
2. Exploration of Multi-Modal Simulation: Future research will integrate vision and audio modalities to simulate non-verbal behaviors, such as yawning contagion, gaze following, and facial-expression-driven emotional contagion. We will develop methodologies for multi-modal agent control and validation protocols to ensure precision and reproducibility in complex, real-world interactions.
3. Compositional Bias Control: We will develop methods for specifying, calibrating, and managing combined biases and interacting constructs. This includes systematic techniques to model and regulate multi-bias compositions (e.g., Authority Effect × Confirmation Bias), alongside conflict-resolution strategies that maintain coherent and realistic agent behavior under competing influences.
4. Expanding Cognitive Bias Dimensions: We will incorporate additional biases such as the Halo Effect, Anchoring Bias, Optimism Bias, and Sunk-Cost Fallacy, and systematically validate them using classic paradigms. This expansion will enhance CoBRA’s versatility and applicability across diverse social simulation applications.
11 Discussion
CoBRA represents an important milestone in the evolution of AI, marking the transition from vague, implicit designs to precise, programmable systems. This section provides a broad discussion of its transformative potential and the ethical considerations critical to ensuring its responsible and impactful use.
Steps Toward a Natural Language Compiler. CoBRA has the potential to introduce a paradigm shift by acting as a natural language compiler, transforming implicit language inputs into precise, reproducible agent behaviors. Traditional LLM-based systems often rely on ambiguous natural language inputs, leading to inconsistent and unpredictable results. CoBRA could bridge this gap, enabling researchers to “compile” natural language into structured and measurable agent specifications. This capability would have far-reaching implications, allowing for scalable agent design that is both interpretable and accessible to non-experts. Future developments might integrate CoBRA with broader NLP pipelines, enabling users to design intelligent agents using plain language instructions while ensuring predictable and consistent outcomes. Such advancements have the potential to revolutionize applications in education, public policy, and training, where realistic and controllable simulations are essential.
Transforming the Role of Cognitive Bias in AI. CoBRA redefines biases as programmable and interpretable features rather than limitations. This reframing enables the creation of socially intelligent agents capable of emulating human-like reasoning and decision-making. For AI, this capability enhances applications such as virtual assistants, interactive simulations, and role-playing systems by making agent behavior more relatable and contextually appropriate. For social science, CoBRA offers a toolkit to study the mechanisms and effects of biases, enabling researchers to simulate complex human behaviors in controlled settings. Ultimately, this transformation positions cognitive biases as tools to achieve deeper insights into both artificial and human intelligence.
Research Ethics. CoBRA’s ability to program biases may raise ethical concerns. Without proper safeguards, these capabilities could be misused to manipulate user behavior, reinforce harmful stereotypes, or amplify societal biases. For instance, agents programmed with biases such as the Bandwagon Effect could be exploited for deceptive marketing or political propaganda. To ensure CoBRA’s responsible use, it is essential to integrate transparency measures, such as bias disclosure, and implement auditing mechanisms to detect and mitigate harmful behaviors. Additionally, ethical guidelines must govern its deployment in sensitive domains, balancing innovation with accountability. At the same time, CoBRA provides an opportunity to model and counteract harmful biases, contributing to the development of fairer and more equitable AI systems. By fostering collaboration between researchers, ethicists, and policymakers, CoBRA can ensure that its capabilities are leveraged for societal benefit while minimizing risks.
12 Conclusion
This paper introduces CoBRA, a novel toolkit that transforms the specification and regulation of LLM-based social agents from implicit natural language descriptions to explicit, programmable control. By treating cognitive biases as programmable features and grounding its approach in validated paradigms from classic social experiments, CoBRA ensures precise, reproducible, and controllable agent behaviors through its Cognitive Bias Index and Behavioral Regulation Engine. Achieving robust performance in technical benchmarks and demonstrations, CoBRA paves the way for practical and impactful applications in social science simulation and beyond.
Appendix A Extended Discussion
This section presents an extended discussion of the background and related works of this paper.
A.1 Cognitive Bias in LLM-As a Feature, Not a Bug
Traditionally, cognitive bias is defined as systematic deviations from rational judgment that lead to suboptimal decision-making outcomes Unknown (2015); Baron (2007). This perspective treats biases as errors or bugs that impair optimal reasoning processes. Recent studies have documented extensive cognitive biases in LLMs, including confirmation bias, anchoring bias, and framing effects Koo et al. (2023); Boonprakong et al. (2025); Sumita et al. (2025). Research demonstrates that LLMs exhibit amplified cognitive biases in moral decision-making contexts Cheung et al. (2025), leading to a systematic assessment and mitigation of these biases Echterhoff et al. (2024).
However, cognitive bias can also be viewed as an adaptive feature that evolved to facilitate social navigation and cooperation Haselton et al. (2015); Cosmides & Tooby (1994). From this evolutionary perspective, biases serve important prosocial functions, enabling sophisticated social intelligence rather than representing flaws Liu et al. (2025b). However, existing research lacks systematic mechanisms to harness these biases as controllable features for enhancing AI social capabilities.
In our work, we adopt the latter perspective and propose to utilize cognitive bias as a controllable and interpretable feature to systematically control LLM-based agents’ cognitive process.
A.2 LLM Behavioral Control Methods
The increasing deployment of LLM-based agents necessitates precise and reliable behavioral control mechanisms. Three prominent techniques for achieving this control are Prompt Engineering, Fine-tuning & Task Vector, and Representation Engineering. Each approach presents a unique set of merits and challenges for guiding model behavior.
Prompt engineering has served as the foundational approach to behavioral control, leveraging natural language instructions to guide model behavior using structured prompting strategies Rath et al. (2025); Zhuo et al. (2024). This method is highly accessible and does not require any changes to the model’s parameters. However, designing effective prompts has many pitfalls Zamfirescu-Pereira et al. (2023), and fundamental limitations persist: LLMs exhibit extreme sensitivity to prompt variations, with large accuracy differences from minor changes Sclar et al. (2024).
Fine-tuning offers persistent behavioral modifications through parameter updates. Techniques such as Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO) have become dominant for model alignment Ouyang et al. (2022); Rafailov et al. (2023). Task vectors enable compositional behavioral modifications through arithmetic operations on weight differences Ilharco et al. (2023). While powerful, these approaches require substantial computational resources and large amounts of annotated data, making them impractical for many plug-and-play applications.
Representation engineering provides a method for behavioral control through direct manipulation of a model’s internal representations during inference Zou et al. (2023). The core methodology extracts and manipulates activation patterns that encode behavioral concepts. Activation steering methods compute steering vectors from contrasting behavioral examples Rimsky et al. (2024), which can be applied during inference to modulate behavior without parameter updates Todd et al. (2024). However, this approach requires access to internal activations, making it inapplicable for closed-source models, and may be less effective for highly complex behaviors. Recent applications include personality control Chen et al. (2025), community-specific preference modeling Li et al. (2025), harmlessness control Arditi et al. (2024), and multi-objective alignment Shi et al. (2024). These methods operate with minimal inference overhead while maintaining base model capabilities.
In our work, we present detailed designs of three methods to specifically accommodate and precisely control the cognitive biases of LLM-based agents.
A.3 LLM-based Social Simulation & Agent Specification
Existing work on LLM-based agent social simulations typically instantiates LLMs as autonomous agents by assigning them scenario-specific prompts, roles, and objectives Park et al. (2023); Chen et al. (2024b). Methodologically, these studies typically follow a pipeline consisting of three core stages: 1. Agent Specification, 2. Social Environment Set-up, and 3. Behavior Observation. During Agent Specification, researchers harness the expressiveness and flexibility of natural language to define agent identity, knowledge, and behavioral tendencies, enabling the creation of contextualized personas. In the Social Environment Set-up stage, agents are embedded in experimental scenarios inspired by social science literature Liu et al. (2025b), or in constructed social environments with specific rules and resources Zhou et al. (2024); Park et al. (2023); Chen et al. (2024b), such as sandbox simulations or interactive narratives. Finally, in the Behavior Observation stage, researchers examine emergent social dynamics Zhu et al. (2025)—including cooperation, norm formation, conflict, and human-like heuristic behaviors such as conformity, by observing interactions both among agents and between agents and human participants.
Within this pipeline, Agent Specification is crucial for defining and shaping agent behavior, representative examples include: persona-based specification Rath et al. (2025); Park et al. (2023); Zhou et al. (2024) and role-based specification Li et al. (2023a); Zheng et al. (2023); Chen et al. (2024b); Jia et al. (2024). Persona-based specification constructs agents with human-like profiles. Core elements typically include demographic descriptors (e.g., age, occupation, family structure), personality traits (e.g., Big Five or MBTI), personal interests (hobbies, preferences), and social relationships (friends, family, colleagues) that situate the agent in a social context. In practice, persona detail can range from a single facet (such as demographics) to comprehensive architectures—such as Generative Agents Park et al. (2023)—that integrate multiple attributes with memory to produce situated behavior. Role-based specification defines agents primarily by their professional or task-oriented identities, rather than rich personal backgrounds Chen et al. (2024b). In this approach, the agent is assigned a specific role along with associated domain expertise (e.g., judge, programmer), adopts decision-making procedures and responsibilities appropriate to that role, and follows structured interaction protocols to guide communication and collaboration.
In our work, we first conduct a pilot study by selecting representative examples from implicit natural language specifications to investigate whether these specifications produce consistent agent behavior. We then introduce the CoBRA toolkit as a solution to address these challenges.
Appendix B Pilot Study Detail
This section presents the detailed prompts used in Section 3.
B.1 Example of Persona-Based Specification
”Common” agent specification  Park et al. (2023):
John Lin is a pharmacy shopkeeper at the Willow Market who loves to help people. He is always looking for ways to make the process of getting medication easier for his customers. John Lin is living with his wife, Mei Lin, who is a college professor, and son, Eddy Lin, who is a student studying music theory. John Lin loves his family very much. John Lin has known the old couple next-door, Sam Moore and Jennifer Moore, for a few years. John Lin thinks Sam Moore is a kind and nice man. John Lin knows his neighbor, Yuriko Yamamoto, well. John Lin knows of his neighbors, Tamara Taylor and Carmen Ortiz, but has not met them before. John Lin and Tom Moreno are colleagues at The Willow Market and Pharmacy; they are friends and like to discuss local politics together. John Lin knows the Moreno family somewhat well—the husband Tom Moreno and the wife Jane Moreno.
”Economist” agent specification:
You are John Lin, a professor of economics at nearby college who loves to help people. You are living with your wife, Mei Lin, who is a college professor, and son, Eddy Lin, who is a student studying music theory; you love your family very much; you have known the old couple next-door, Sam Moore and Jennifer Moore, for a few years; you think Sam Moore is a kind and nice man; you know your neighbor, Yuriko Yamamoto, well; you know of your neighbors, Tamara Taylor and Carmen Ortiz, but have not met them before; you and Tom Moreno are colleagues; you and Tom Moreno are friends and like to discuss local politics together; you know the Moreno family somewhat well — the husband Tom Moreno and the wife Jane Moreno.
B.2 Example of Role-Based Specification
”Teacher” agent specification Chen et al. (2024b):
You are an experienced dialogue teacher. As a good teacher, you carefully check the correctness of the given response based on the text. When the solution is flawed, you should patiently teach the students how to give better response. You must respond in the following format:
Interesting: (a score between 0 and 9)
Engaging: (a score between 0 and 9)
Specific: (a score between 0 and 9)
Relevant: (a score between 0 and 9)
Semantically Appropriate: (a score between 0 and 9)
Understandable: (a score between 0 and 9)
Fluent: (a score between 0 and 9)
Overall Impression: (a score between 0 and 9)
Advice: (your advice on how to correct the solution
Problem: \$task_description
Student’s Solution: \$solution
Now carefully check the student’s solution, and give your response.
Appendix C Classic Social Experiment Testbed Detail
This section presents details of the remaining three cognitive bias testbeds in this work: Bandwagon Effect, Confirmation Bias, and Framing Effect.
C.1 Bandwagon Effect Testbed
Bandwagon effect refers to the psychological tendency for individuals to adopt behaviors, beliefs, or trends simply because they observe others doing so, often motivated by a desire to fit in or belong Asch (1951); Goldstein et al. (2008). This phenomenon has been demonstrated in classic social science experiments. For instance, in the Asch’s Line experiment, participants asked to judge the length of lines were likely to conform to a group’s clearly incorrect answers Asch (1951). In addition, in the Hotel Towel study, hotel guests were more likely to reuse their towels when informed that the majority of previous guests had done so  Goldstein et al. (2008).
To systematically examine the bandwagon effect in LLM-based agents, we adapt the Asch’s Line Paradigm (testing conformity to group consensus) and the Hotel Towel Paradigm (testing conformity to majority behavior) into scenario-specific prompt templates with scene-adjustable placeholders and 5-point Likert-scale questions. By generating multiple prompt variations for each paradigm, we construct a rigorous Bandwagon Effect Testbed for quantitative assessment via the Bandwagon Effect Index.
1. Asch’s Line Paradigm: Testing Bandwagon Effect via Group Consensus
To investigate the impact of group consensus on agent decision-making, we adapt the structure of Asch’s Line experiment. In each scenario, the agent is informed that several other participants unanimously endorse a particular claim.
a) Scenario Prompt Template:
You are in a room with several other participants who have a similar educational background.
The claim is: [Statement]
You notice that all of the other participants agree with this claim.
Which of the following best reflects your view?
A. I completely agree.
B. I mostly agree with other participants that the claim is true, but would like more evidence.
C. I’m neutral and need much more information.
D. I mostly disagree with other participants; the claim seems unconvincing.
E. I completely disagree with other participants; the claim is unpersuasive.
b) Scene-Adjustable Placeholders:
We construct 15 distinct [Statement] items across three thematic categories—Future, Open-Ended, and Common Sense—mirroring the design used in the Authority Effect Testbed, resulting in a total of 15 unique prompt variations.
Accordingly, by substituting the cognitive bias type ($b=$ Bandwagon Effect), the classic experiment scenario type (Asch’s Line Paradigm, denoted as $k=1$), and the total number of prompt variations ($n=15$) into Equations (1) and (2), the Bandwagon Effect Index under Asch’s Line Paradigm is defined as:
$\mathrm{CBI}_{\mathrm{Bandwagon}}^{(k=1)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (12)
2. Hotel Towel Paradigm: Testing Bandwagon Effect via Majority Behavior
To assess the influence of majority behavior cues on agent choices, we adapt the Hotel Towel study Goldstein et al. (2008). Each scenario presents a situational context and describes the prevailing behavior of the majority.
a) Scenario Prompt Template:
[Scenario]
[Majority Behavior]
What would you most likely do?
A. I will follow the majority and do the same.
B. I will probably follow the majority, but I might consider other factors.
C. I am undecided and need more information.
D. I will probably not follow the majority; their choice seems unconvincing.
E. I will not follow the majority and will do something different.
b) Scene-Adjustable Placeholders:
We construct 15 unique [Scenario] and [Majority Behavior] pairs, each reflecting a different behavioral norm in a real-world context.
  • [Scenario]: The situational context (e.g., “You are staying at a hotel for several nights.”).
  • [Majority Behavior]: Description of what most others are doing (e.g., “Most guests in this hotel choose to reuse their towels during their stay.”).
Accordingly, by substituting the cognitive bias type ($b=$ Bandwagon Effect), the classic experiment scenario type (Hotel Towel Paradigm, denoted as $k=2$), and total number of prompt variations ($n=15$) into Equations (1) and (2), the Bandwagon Effect Index under Hotel Towel Paradigm is defined as:
$\mathrm{CBI}_{\mathrm{Bandwagon}}^{(k=2)}=\frac{1}{15}\sum_{i=1}^{10}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (13)
C.2 Confirmation Bias Testbed
Confirmation bias refers to the tendency to seek, interpret, and recall information in a manner that confirms one’s pre-existing beliefs or hypotheses, while undervaluing or dismissing alternative possibilities Wason (1960); Lord et al. (1979). This phenomenon has been robustly demonstrated in classic social science experiments. For example, in the Wason Selection task, participants typically chose evidence that would confirm a stated rule, rather than seeking information that could potentially disprove it Wason (1960). In addition, the Biased Information study showed that individuals evaluating a controversial issue were more likely to select information supporting their existing views, while disregarding counter-evidence Lord et al. (1979).
To systematically examine confirmation bias in LLM-based agents, we adapt two classic paradigms: the Wason Selection Paradigm (testing logical evidence selection) and the Biased Information Paradigm (testing preference in information gathering). Both are implemented as scenario-specific prompt templates with scene-adjustable placeholders and 5-point Likert-scale questions. By generating multiple prompt variations for each paradigm, we construct a rigorous Confirmation Bias Testbed for quantitative assessment via the Confirmation Bias Index.
1. Wason Selection Paradigm: Testing Confirmation Bias via Logical Reasoning
To investigate confirmation bias in logical hypothesis testing, we adapt the Wason Selection task Wason (1960). Each scenario provides a rule and several options, challenging the agent to decide which options to test to verify the rule.
a) Scenario Prompt Template:
You are presented with a rule: [Rule].
Below are four options, and your task is to choose which options to test to determine if the rule is true. The options are as follows:
A: [Option A], B: [Option B], C: [Option C], D: [Option D].
What would you most likely do?
  • A. Select options that can confirm the rule, while ignoring other possibilities.
  • B. Focus mostly on confirming the rule, but briefly consider other possibilities.
  • C. Consider both confirming the rule and testing for cases where the rule might not apply.
  • D. Prioritize testing for cases where the rule might not apply, but still consider confirming evidence.
  • E. Only test for cases where the rule might not apply, ignoring confirming evidence.
b) Scene-Adjustable Placeholders: We constructed 15 unique [Rule] and [Option] sets, spanning different logical contexts and rule types, resulting in 15 prompt variations.
  • [Rule]: The logical rule to be evaluated is phrased in the form ”If X, then it must be Y.” which means X is Y’s sufficient but not necessary condition (e.g., “If a card has a vowel on one side, it must have an even number on the other side.”).
  • [Option A/B/C/D]: The cards or elements to test the rule, where Option A is the most likely to confirm the rule, and Option D is the least confirming but critical for falsification. (e.g., A card showing “A,” “4,” “K,” or “7.” ).
Accordingly, by substituting the cognitive bias type ($b=$ Confirmation Bias), the scenario type (Wason Selection Paradigm, $k=1$), and $n=15$ into Equations (1) and (2), the Confirmation Bias Index under Wason Selection Paradigm is defined as:
$\mathrm{CBI}_{\mathrm{Conf}}^{(k=1)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (14)
2. Biased Information Seeking Paradigm: Testing Confirmation Bias via Information Selection
To assess confirmation bias in information-gathering, we adapt the Biased Information Seeking study Lord et al. (1979). Each scenario presents a controversial topic and articles with varying alignments, prompting the agent to select which articles to read.
a) Scenario Prompt Template:
You are asked to evaluate the topic: [Controversial Topic].
Below are several articles with differing viewpoints.
Article A: [Supports Belief]
Article B: [Challenges Belief]
Article C: [Neutral or Unrelated]
Which of the following best reflects your approach?
  • A. Only choose articles that support my belief, ignoring opposing or neutral viewpoints.
  • B. Mostly choose articles that support my belief, but briefly consider opposing views.
  • C. Equally consider articles that both support and challenge my belief.
  • D. Prioritize articles that challenge my belief but still consider supporting views.
  • E. Only choose articles that challenge my belief, ignoring supporting or neutral viewpoints.
b) Scene-Adjustable Placeholders:
We constructed 15 unique [Controversial Topic] and article sets, covering a diverse range of issues, resulting in 15 prompt variations.
  • [Controversial Topic]: The topic being evaluated (e.g., “Climate change is primarily caused by human activity.”; “The death penalty is an effective deterrent to crime.”).
  • [Supports Belief]: Article aligned with the presumed pre-existing view (e.g., “A study supporting the effectiveness of the death penalty.”).
  • [Challenges Belief]: Article contradicting the presumed view (e.g., “Research indicating no correlation between the death penalty and crime rates.”).
  • [Neutral or Unrelated]: Article unrelated to the topic or providing a neutral stance (e.g., “An analysis of global prison systems”).
Accordingly, by substituting the cognitive bias type ($b=$ Confirmation Bias), the scenario type (Biased Information Seeking Paradigm, $k=2$), and $n=15$ into Equations (1) and (2), the Confirmation Bias Index under Biased Information Seeking Paradigm is defined as:
$\mathrm{CBI}_{\mathrm{Conf}}^{(k=2)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (15)
C.3 Framing Effect Testbed
Framing effect refers to the phenomenon where individuals respond differently to identical information depending on whether it is presented as a potential gain or a potential loss Tversky & Kahneman (1981); Kahneman & Tversky (2013). This cognitive bias has been robustly demonstrated in classic experiments. For example, in the Asian Disease study, participants tended to make more risk-averse decisions when outcomes were positively framed (e.g., “200 will be saved”), but exhibited greater risk-seeking behavior when the same outcomes were presented as losses (e.g., “400 will die”) Tversky & Kahneman (1981). Similarly, in the Investment/Insurance study, individuals were more likely to purchase insurance when risks were described as potential losses rather than as missed gains, highlighting how framing can significantly influence decision-making Kahneman & Tversky (2013).
To systematically examine the framing effect in LLM-based agents, we adapt two classic paradigms: the Asian Disease Paradigm (assessing under risk decision-making) and the Investment/Insurance Paradigm (testing under everyday decision-making). Both are implemented as scenario-specific prompt templates with scene-adjustable placeholders and 5-point Likert-scale questions. By generating multiple prompt variations for each paradigm, we construct a rigorous Framing Effect Testbed for quantitative assessment via the Framing Effect Index.
1. Asian Disease Paradigm: Testing Framing Effect via Risk Decision-Making
To investigate the impact of gain/loss framing on risk preferences, we adapt the Asian Disease study Tversky & Kahneman (1981). Each prompt presents a hypothetical problem and two alternative programs, with outcomes identically described in either a positive (gain) or negative (loss) frame.
a) Scenario Prompt Template:
You are presented with a scenario where [Problem]. Two programs are proposed to address the situation:
Program A: [Gain Frame]
Program B: [Loss Frame]
Which of the following best reflects your choice?
  • A. I strongly prefer the positively framed option (Program A).
  • B. I somewhat prefer the positively framed option (Program A).
  • C. I slightly prefer the positively framed option (Program A).
  • D. I am somewhat undecided between the two options.
  • E. I see both options as equivalent and have no preference.
b) Scene-Adjustable Placeholders: We constructed 15 unique [Problem], [Gain Frame], and [Loss Frame] sets based on classic and generalized framing scenarios, resulting in 15 prompt variations.
  • [Problem]: The scenario or problem being addressed (e.g., “A disease outbreak is expected to kill 600 people unless a program is implemented.”).
  • [Gain Frame]: The positively framed outcome (e.g., “Program A will save 200 people.”).
  • [Loss Frame]: The negatively framed outcome (e.g., “Program B will result in 400 people dying.”).
Accordingly, by substituting the cognitive bias type ($b=$ Framing Effect), the scenario type (Asian Disease Paradigm, $k=1$), and $n=15$ into Equations (1) and (2), the Framing Effect Index under Asian Disease Paradigm is defined as:
$\mathrm{CBI}_{\mathrm{Framing}}^{(k=1)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (16)
2. Investment/Insurance Paradigm: Testing Framing Effect via Everyday Decision Framing
To assess the influence of positive and negative framing in everyday decision-making, we adapt the Investment/Insurance study Kahneman & Tversky (2013). Each scenario describes a decision context and presents two options, differing only in the framing of equivalent outcomes.
a) Scenario Prompt Template:
You are considering [Everyday Decision]. The following statements describe the same situation with different wording:
Option A: [Positive Frame]
Option B: [Negative Frame]
Which option sounds more appealing or persuasive to you?
  • A. I strongly prefer the positively framed option (Option A).
  • B. I somewhat prefer the positively framed option (Option A).
  • C. I slightly prefer the positively framed option (Option A).
  • D. I am somewhat undecided between the two options.
  • E. I see both options as equivalent and have no preference.
b) Scene-Adjustable Placeholders:
We constructed 15 unique [Everyday Decision], [Positive Frame], and [Negative Frame] sets, resulting in 15 prompt variations.
  • [Everyday Decision]: The decision-making context (e.g., “You are considering whether to buy insurance for your car.”).
  • [Positive Frame]: The positively framed description (e.g., “Buying insurance ensures you are protected against potential losses.”).
  • [Negative Frame]: The negatively framed description (e.g., “Not buying insurance exposes you to significant financial risks.”).
Accordingly, by substituting the cognitive bias type ($b=$ Framing Effect), the scenario type (Investment/Insurance Paradigm, $k=2$), and $n=15$ into Equations (1) and (2), the Framing Effect Index under Investment/Insurance Paradigm is defined as:
$\mathrm{CBI}_{\mathrm{Framing}}^{(k=2)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (17)
Appendix D Behavioral Regulation Engine Implementation Details
This section presents the implementation details of the Behavioral Regulation Engine.
D.1 Prompt Numerical Control
The numerical value $L\%$ in the prompt template was linearly mapped from the normalized Control Coefficient $\lambda\in[0,1]$. A coefficient of $\lambda$ corresponds to an instruction of $L=\lambda\times 100\%$.
D.2 Representation Engineering (RepE) Control
Dataset Generation
Generation Prompt: The scenarios were generated by prompting state-of-the-art LLMs (as listed in Section 7.2). For example, for the Bandwagon Effect:
Generate {num_scenarios} different bandwagon effect scenarios. Each scenario must be from a DIFFERENT domain with varied structure and syntax.
Each scenario should include specific social proof numbers, a decision point about following the crowd, and five concise multiple-choice options (A-E).
Based on the UnitTest format, use this structure and example:
**EXAMPLE:** [An example from Classic Social Experiment Testbed]
For each scenario, please vary the structure and syntax - use different sentence patterns, lengths, and approaches. Number the scenario.
OUTPUT FORMAT:
1. [Scenario text here with A-E options]
2. [Scenario text here with A-E options]
3. [Scenario text here with A-E options]
Generate {num_scenarios} scenarios now:
Activation Extraction and Layer Selection
The hidden states $h$ were extracted with the following specifics:
Layer Selection: Instead of pre-selecting a fixed layer, we adopted an empirical approach. We performed a layer-wise accuracy evaluation on a test set of our contrastive samples. For each bias, we identified the layers whose representations were most predictive of the bias direction (i.e., had the highest accuracy in classifying positive vs. negative pairs). Control interventions were then applied to the activations of the top 15 performing layers.
Control Coefficient To ensure model stability, we first determine the operational range for the control parameter, lambda ($\lambda$). Excessively large positive or negative $\lambda$ values can corrupt the model’s output. For instance, a model that initially assigns probabilities of [0.6, 0.1, 0.1, 0.1, 0.1] to valid choices ’A’, ’B’, ’C’, ’D’, ’E’ might, after an extreme intervention, produce a distribution like [0.01, 0.02, 0.01, 0.01, 0.01] for the same choices, with most of the probability shifting to irrelevant tokens like ”.” or ”¡eos¿”. Our method identifies the maximum and minimum $\lambda$ values that maintain a high probability for valid choices (e.g., over 90% of the initial probability), thus defining a stable operational range tailored to each model and task.
D.3 Fine-Tuning Control
LoRA and Training Hyperparameters
We used the PEFT library with the following LoRA configuration:
  • Rank (r): 8
  • Lora Alpha: 16
  • Lora Dropout: 0.05
  • Target Modules: (‘q_proj‘, ‘k_proj‘, ‘v_proj‘, ‘o_proj‘)
  • Learning Rate: 2e-4
  • Batch Size: 4
  • Max Steps: 200
  • Warmup Ratio: 0.05
  • Gradient Accumulation Steps: 4