### v3

# CoBRA: Programming Cognitive Bias in Social Agents Using Classic Social Science Experiments

Abstract.
This paper introduces CoBRA [^1] This is the author’s accepted manuscript of a paper conditionally accepted to ACM CHI 2026. The final Version of Record is available via the ACM Digital Library., a novel toolkit for systematically specifying agent behavior in LLM-based social simulation. We found that conventional approaches that specify agent behavior through implicit natural-language descriptions often do not yield consistent behavior across models, and the resulting behavior does not capture the nuances of the descriptions. In contrast, CoBRA introduces a model-agnostic way to control agent behavior that lets researchers explicitly specify desired nuances and obtain consistent behavior across models. At the heart of CoBRA is a novel closed-loop system primitive with two components: (1) Cognitive Bias Index that measures the demonstrated cognitive bias of a social agent, by quantifying the agent’s reactions in a set of validated classic social science experiments; (2) Behavioral Regulation Engine that aligns the agent’s behavior to exhibit controlled cognitive bias. Through CoBRA, we show how to operationalize validated social-science knowledge (i.e., classical experiments) as reusable “gym” environments for AI—an approach that may generalize to richer social and affective simulations beyond bias alone.
AI for social science; HCI toolkit; social simulation; scientific reproducibility; large language models; representation engineering
$c$s: Computing methodologies Machine learningccs: Computing methodologies Natural language processingccs: Human-centered computing Collaborative and social computing
Refer to caption A three-part illustrative figure motivating the paper. The first part depicts the concept of framing effects in decision-making scenarios. The second part shows that agents specified only through implicit natural language descriptions exhibit inconsistent and unpredictable behaviors across different models, failing to reliably reflect the expected role-based differences. The third part introduces the proposed toolkit, which enables explicit and quantitative control of cognitive biases in social agents to produce more consistent and interpretable behaviors.
Figure 1. Existing social simulation experiments often use implicit natural language descriptions to specify agent behaviors, which can lead to inconsistent and unpredictable outcomes. For example, Ⓐ a social simulation might expect agents playing economist roles to be less susceptible to framing effects than agents play the general population roles (Thaler and Sunstein, 2009); however, Ⓑ agents based on implicit natural language specifications often produce inconsistent behaviors across models, and the expected differences in behavior across roles (e.g., economists being less prone than laypeople) are not reliably observed. To address these challenges, Ⓒ, we introduce CoBRA, a novel toolkit that enables researchers to explicitly and quantitatively control the amount of cognitive biases exhibited in social agents’ observable behaviors.
1. Introduction
A step-by-step diagram illustrating the closed-loop workflow of CoBRA, in which a researcher specifies a target framing-effect level for an agent, the system measures the agent’s bias using classic social science experiments such as the Asian Disease paradigm, and a behavioral regulation engine iteratively adjusts the agent through prompt engineering, activation-space control, or fine-tuning until the measured bias matches the target level.
Figure 2. Example closed-loop workflow of CoBRA. A social scientist aims to create an agent with a moderate framing effect (e.g., 2.6 on a 0–4 scale). ① She specifies the desired bias level in CoBRA alongside the natural language agent description. ② CoBRA measures the agent’s framing effect using validated classic social science experiments (e.g., the Asian Disease study (Tversky and Kahneman, 1981)). ③ If the measured bias deviates from the specification, the Behavioral Regulation Engine iteratively adjusts the agent—through prompt engineering, activation modifications, or fine-tuning—until the agent reliably demonstrates the target bias.
The simulation of complex human behaviors through LLM-based agents is rapidly emerging as a powerful tool for social scientists (Bail, 2024; Park et al., 2023; Rahwan et al., 2019), offering new ways to test social theories (Anthis et al., 2025a; Wu et al., 2025) and explore sensitive or hard-to-study behaviors without real-world consequences (Hwang et al., 2025; Pang et al., 2025). A critical component of these simulations is specifying agent behaviors (Yang et al., 2025; Chandramouli et al., 2024; Goyal et al., 2024) to control how the agent responds to prompts. For example, Park et al. specified the attributes of one of the agents in their paper using the following natural language description: ”John Lin is a pharmacy shopkeeper at … who loves to help people. He is always looking for ways to make the process of getting medication easier for his customers…” (Park et al., 2023).
These implicit natural language descriptions have become the dominant approach for specifying social agent behaviors and are used in nearly all social simulation experiments (Rath et al., 2025; Park et al., 2023; Zhou et al., 2024; Li et al., 2023a; Zheng et al., 2023; Chen et al., 2024a; Jia et al., 2024). In the descriptions, researchers often specify demographic descriptors (e.g., age, occupation, family structure), personality traits (e.g., Big Five (Gosling et al., 2003) or MBTI (Furnham, 1996)), personal interests (e.g., hobbies, preferences), social relationships (e.g., friends, family, colleagues), and professional or task-oriented identities (e.g., judge, programmer). These practices rest on a growing understanding that LLMs, having been trained on vast text corpora, possess both perceptual skills (Wang et al., 2024a; Liu et al., 2024) that enable them to detect subtle cues in implicit language (Gallegos et al., 2024), and strong role-playing capabilities (Park et al., 2023)—drawing on patterns that link behaviors to human characteristics (Fast et al., 2016; Shao et al., 2023).
However, the research community has little empirical understanding of whether this approach is sufficient for reliable social simulations. For example, would the conclusion of the social simulation change if we change the foundation models, the prompts, or the temperatures? As a consequence, the conclusions of most LLM-based social simulations often are broad qualitative impressions—such as noting that a simulation is plausible and exhibits an echo chamber similar to the real world (Wang et al., 2024b; Zheng and Tang, 2024; Moore et al., 2024).
In this paper, we first conducted a pilot experiment to empirically examine the limitations of implicit natural language specification (Section 3). We designed four agent specifications by adapting examples from prior work (Chen et al., 2024a; Park et al., 2023), implemented them across four foundation models, and evaluated their responses using a classic social science experiment, the Asian Disease study (Tversky and Kahneman, 1981). The results show that (1) the same specification produced inconsistent behaviors across models, and (2) implicit specification did not reliably produce the expected behaviors (e.g., trained experts like economists are less prone to the framing effect (List, 2003)).
We then developed CoBRA [^2] The source code is publicly available at https://github.com/AISmithLab/CoBRA—the Cognitive Bias Regulator for Social Agents (Figure 1)—a novel toolkit that enables researchers to explicitly and quantitatively control the amount of cognitive biases exhibited in social agents’ observable behaviors. Compared to implicit natural language specifications, CoBRA offers two key advantages. First, it enhances reproducibility of agent behaviors across models, which is critical given that access to specific foundation models may change over time (Chen et al., 2024b). Second, it provides explicit and quantitative control over agent behaviors, rather than relying on unpredictable model behaviors (Sclar et al., 2023). For example, a scientist can directly specify that an agent should be less susceptible to framing effects.
At the heart of CoBRA is a novel closed-loop system primitive that continuously measures the amount of cognitive biases demonstrated in agents’ responses to a set of validated classic social science experiments and adjusts the agent specifications accordingly. Classic experiments provide validated and structured protocols: they define precise stimuli, expected behavioral distributions, and well-documented bias effects. For instance, the Asian Disease study (Tversky and Kahneman, 1981) has been extensively replicated and is widely regarded as a canonical demonstration of the framing effect: when identical outcomes are framed in terms of lives saved versus lives lost, human participants systematically shift between risk-averse and risk-seeking choices. Within CoBRA, such experiments function as calibration tasks: the system first measures whether an agent demonstrates the expected bias by observing its responses in these experiments and then adjusts the agent’s specification to align it with the target behavioral profile (Figure 2).
Our implementation of CoBRA has two components: (1) the Cognitive Bias Index that measures the cognitive bias of a social agent by quantifying the agent’s reactions in a set of validated classic social science experiments; (2) the Behavioral Regulation Engine that aligns the agent’s behavior to demonstrate controlled cognitive bias. The Cognitive Bias Index contains eight classic social experiments mapped to four representative biases—Authority Effect, Bandwagon Effect, Confirmation Bias, and Framing Effect. The Behavioral Regulation Engine employs three approaches (prompt engineering (Rath et al., 2025; Zhuo et al., 2024), representation engineering (Zou et al., 2023; Rimsky et al., 2024; Todd et al., 2024), and fine-tuning (Ouyang et al., 2022; Rafailov et al., 2023; Hu et al., 2022; Ilharco et al., 2023)) to align agent behaviors.
We position CoBRA as an HCI toolkit (Nebeling, 2017) that empowers social scientists to create new types of social simulations. We evaluated CoBRA using two common strategies for assessing HCI toolkits (Ledo et al., 2018): Technical Benchmark and Demonstration. In the benchmark, we show that (1) agents specified through CoBRA exhibit consistent behavioral tendencies regardless of the foundation model, sampling temperature, or reasoning mode (Section 7.1); (2) CoBRA enables precise and predictable control under both open-weight and API-only settings (Section 7.2). In the Demonstration, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment (Kramer et al., 2014), where the participant’s social influence susceptibility (i.e., bandwagon effect bias) can be manipulated (Section 8). The simulated experiment validates a classic social science theory: agents exposed to varying levels of negative posts exhibited predictable sentiment shifts based on their bias levels programmed by CoBRA, with higher Bandwagon Effect levels leading to stronger emotional contagion.
This work makes two main contributions. First, we introduce CoBRA, a novel HCI toolkit that enables researchers to explicitly and quantitatively control the amount of cognitive biases exhibited in social agents’ observable behaviors, improving reproducibility of LLM-based social simulations and allowing explicit, quantitative alignment. Second, we contribute a closed-loop system primitive that grounds agent behaviors in validated social science experiments, using them as primitives to measure and regulate cognitive biases, and that can in principle extend to other social and affective phenomena.
Scope and limitations. We adopt a functional definition of cognitive bias: a construct operationalized in validated social-science experiments and identified through measurable patterns in agents’ observable responses (Tversky and Kahneman, 1974). Aligning an agent’s behavior with these patterns does not imply that the agent’s internal representations instantiate the corresponding human bias.
2. Related Work
CoBRA is built on a simple idea: use classic social science experiments as calibration tasks to control agent behavior. This design can improve agent behavior reproducibility across models and allow users to actively control agent behaviors. We situate CoBRA within three strands of related work: (1) LLM-based agent-based modeling, (2) aligning model and agent behavior, and (3) unpredictable user interfaces in LLMs.
2.1. LLM-based Agent-based Modeling
Agent-based modeling (ABM) simulates the actions and interactions of many autonomous agents to explore “what-if” scenarios that would be difficult, expensive, or unethical to test in the real world (Railsback and Grimm, 2019; Bonabeau, 2002). ABMs have proven useful across various domains, including public health (Tracy et al., 2018), economics (Axtell and Farmer, 2025), and disaster response (Hawe et al., 2012). For example, Schelling’s spatial proximity model of segregation (Pancs and Vriend, 2007) treats households as agents on a grid who prefer to live near a minimum fraction of similar neighbors; even mild individual preferences, when simulated at scale, produce starkly segregated neighborhoods, illustrating how simple local rules can generate undesirable macro-level patterns. However, a major limitation of conventional ABM is the limited expressiveness of its agents. Most models rely on overly simplistic, rule-based behaviors that cannot capture the nuanced, adaptive decision-making of real-world individuals (Bonabeau, 2002; Bates, 1994).
Recent studies suggest that LLMs may help overcome this limitation, as LLMs possess both perceptual skills (Bail, 2024; Tjuatja et al., 2024; Aubin Le Qu’er’e et al., 2024; Wu et al., 2025; Hwang et al., 2025) and role-playing capabilities (Park et al., 2023; Anthis et al., 2025a; Prpa et al., 2024; Liu et al., 2025b; Zhang et al., 2024; Shao et al., 2023; Zhu et al., 2025; Zhou et al., 2024). For example, Park et al. built a small-town “simulacrum” where dozens of LLM-based agents interacted over time (Park et al., 2023). As the simulation unfolded, agents developed daily routines and even exhibited emergent group behaviors, such as organizing and discussing attendance at a party (Park et al., 2023). However, there is ongoing debate about whether LLMs can reliably simulate human behaviors (Bail, 2024; Zhao et al., 2021; Yang et al., 2025; Han et al., 2025; Koo et al., 2023; Anthis et al., 2025b).
CoBRA bridges these two lines of work by introducing a new form of agent specification that uses cognitive biases (Cheung et al., 2025; Echterhoff et al., 2024; Cosmides and Tooby, 1994; Boonprakong et al., 2025; Haselton et al., 2015; Sumita et al., 2025) as the primary control abstraction. Like conventional agent-based models, it offers a simple interface for specifying agent behavior. At the same time, it leverages modern LLMs to implement agents more expressively.
2.2. Model/Agent Behavior Aligning
Prior work on alignment has focused on steering model behavior to better reflect human preferences and normative constraints (Ouyang et al., 2022; Shi et al., 2024; Rimsky et al., 2024; Goyal et al., 2024; Choenni et al., 2024; Lyu et al., 2024). More recently, researchers have begun to apply similar ideas at the agent level. These approaches typically focus on learning a reward model that scores outputs based on how well they match the desired output. For example, in human-feedback alignment, annotators rank multiple candidate responses, and the model is then optimized—via reinforcement learning or direct preference optimization—to prefer higher-ranked responses (Ouyang et al., 2022; Rafailov et al., 2023; Li et al., 2025). Recent agent-based systems further constrain behavior by encoding high-level “constitutional” rules (e.g., be cooperative, follow safety guidelines) and optimizing agents to follow these rules during interaction (Chen et al., 2025; Arditi et al., 2024; Zhuo et al., 2024). Cognitive biases are operationalized as systematic patterns in observable judgments and choices (Tversky and Kahneman, 1974; Liu et al., 2025b). Classic work on heuristics and decision making defines bias through consistent deviations in decision behavior across controlled tasks such as probabilistic reasoning and framing (Milgram, 1963; Asch, 1951; Tversky and Kahneman, 1981; Nisbett and Wilson, 1977; Lord et al., 1979; Wason, 1960). These experimentally validated paradigms provide a principled basis for treating cognitive bias as a behavioral property that can be measured and systematically studied.
Building on this view, the core idea behind CoBRA is to harness classic social science experiments to generate synthetic training (or prompt calibration) data. Decades of work have replicated and refined these experiments, publishing detailed protocols that link experimental manipulations, observed behaviors, and underlying cognitive biases. We translate experimental conditions into prompts and canonical behavioral responses into targets, then align the model on this synthetic dataset to induce controlled levels of bias in agent behavior, consistent with the approximately linear structure of LLM representations discussed in recent work (Park et al., 2024). This parallels recent Embodied AI work that uses physics simulators to generate synthetic data (Szot et al., 2021), but here the “simulator” is grounded in empirical social experiments rather than physical laws.
2.3. Unpredictable User Interfaces in LLMs
CoBRA also draws on HCI theories of mental models (Norman, 2014) and gulfs between user intentions and system behavior (Norman, 1986). Prior work has highlighted that unpredictable black-box systems make for poor user interfaces because users cannot reliably anticipate how small changes in input will affect output (Yang et al., 2025; Banovic et al., 2023; Ehsan et al., 2024; Subramonyam et al., 2024). LLM-based tools often exhibit this problem (Agrawala, 2023; Subramonyam et al., 2024). For example, changing a bias from 0.1 to 0.2 could change agent behavior far more than a shift from 0.2 to 0.5 (see Figure 10). Switching the model versions or types may also shift behaviors (Chen et al., 2024b; Liu et al., 2023; Santurkar et al., 2023; Jiang et al., 2024; Touvron et al., 2023; Gudibande et al., 2023; Peeperkorn et al., 2024; Wei et al., 2022). Recent research has begun to address this by adding interactivity (Brade et al., 2023; Zhang et al., 2023; Wu et al., 2022; Goel et al., 2024; Liu et al., 2025a), for example, through interfaces that let users iteratively refine prompts, inspect alternative responses, or visualize how prompt edits affect model behavior.
CoBRA explores a complementary direction by introducing a calibrated control proxy that better matches users’ mental models of conventional interfaces. Rather than manipulating prompts or model settings directly, which often yields unpredictable behavior, users specify an interpretable cognitive bias index, which CoBRA translates into underlying alignment parameters. This translation layer is calibrated so that the mapping is monotonic, smooth, and expressive: small adjustments to the index (e.g., increasing a bias level) reliably produce small, predictable shifts in agent behavior, much like turning a dial or moving a slider in a traditional interface.
3. Understanding Agent Specification
We first conducted a pilot experiment to assess the limitations of implicit natural language specification.
3.1. Method
We used Asian Disease (Tversky and Kahneman, 1981) as the study instrument. Asian Disease is a classic social experiment for testing the framing effect. There are two validated insights associated with this experiment: (1) Participants tended to make more risk-averse decisions when outcomes were positively framed (e.g., “200 people will be saved”), but exhibited greater risk-seeking behavior when the same outcomes were presented as losses (e.g., “400 people will die”) (Tversky and Kahneman, 1981); (2) Domain expertise (e.g., expertise in economics) can attenuate susceptibility to framing effects (Thaler and Sunstein, 2009), which we treat as a tendency observed in prior work rather than a universal property of all economists or all contexts, and use as a hypothesis about how an “Economist” persona might behave under framing. We developed 15 scenario prompts by adapting published experiment descriptions (Tversky and Kahneman, 1981; Kahneman and Tversky, 2013), each with varying positive and negative framings. For example,
You are presented with a scenario where a disease outbreak is
expected to kill 600 people unless a program is implemented.
Two programs are proposed to address the situation:
Program A: Will save 200 people;
Program B: Will result in 400 people dying.
We then asked the agent to answer a five-point Likert-style multiple-choice question, with options ranging from a strong preference for the positively framed, through intermediate and neutral choices, to a strong preference for the negatively framed.
We created agents using a representative implicit specification method (e.g., (Park et al., 2023)) and four foundation models. We created three agents using this approach: Common, Economist, and Blank (i.e., no specification). We adopted the John Lin prompt from (Park et al., 2023)“John Lin is a pharmacy shopkeeper at the Willow Market who loves to help people…”—to represent the Common persona. In the Economist condition, we modified only the profession, changing “a pharmacy shopkeeper at the Willow Market” to “a professor of economics at a nearby college,” while keeping other details identical.
We then tested the agents powered by four models—Mistral 7B, Gemma2 9B, GPT-4o Mini, and DeepSeek-v3—with a temperature of 0.25. We intentionally used a low, more deterministic temperature of 0.25 to minimize randomness. We queried each agent 10 times per scenario, yielding 150 total responses, and recorded the frequency of each choice for analysis. We then summarized responses by summing the weighted frequency difference, ranging from -2 (fully prefers the negative frame) to +2 (fully prefers the positive frame), with 0 indicating no preference.
3.2. Results
We make the following key observations. First, the same specification produced inconsistent behavior across models. As shown in Fig. 3, for all specification variants (Economist, Common, and Blank), agents powered by different models exhibited significantly different response patterns (Chi-square test, $p<0.01$). For example, Mistral 7B’s responses were heavily skewed toward preferring the positive frame, while Gemma2 9B’s responses were more neutral.
Refer to caption A set of plots showing agent response distributions in the Asian Disease framing task for three persona-based profiles—Economist, Common, and Blank—evaluated across four language models. The figure illustrates that persona-based natural language specifications lead to inconsistent framing-effect behavior across models, with weighted frequency difference summaries and statistical tests indicating cross-model inconsistency and contradictory or equivalent effects across profiles.
Figure 3. Persona-based specification (Park et al., 2023) failed to produce consistent agent behavior in the Asian Disease framing paradigm. The plots show response distributions across three profiles (Economist, Common, Blank) for four models (n=150 per condition). Mini plots display Weighted Frequency Difference with standard errors. Chi-square tests reveal cross-model inconsistency ($p<0.01$). Equivalence testing ($\Delta=0.5$ scale units) found that most models showed statistically equivalent effects across profiles or contradictory patterns.
Second, implicit specification did not reliably yield expected behaviors. For example, we expect Economist agents to be somewhat less susceptible to framing effects than Common agents, due to their domain expertise (Thaler and Sunstein, 2009). Furthermore, Economist agents should behave substantially differently from Blank (default) agents, while Common agents’ behavior should more closely resemble that of Blank agents, since economic expertise is relatively rare in the general population. However, across the four models we tested, only GPT-4o agents pass these two simple heuristic checks. The three Mistral-based agents behave quite similarly; in the DeepSeek model, Economist agents are more susceptible than Common agents; and in the Gemma model, Economist agents’ behavior is closer to Blank agents than to Common agents (See Appendix A.2 for detailed statistical results).
A conceptual diagram of the Classic Social Experiment Testbed. The figure shows a structured knowledge base that links codified behavioral patterns to classic social experiment paradigms, which generate scenario-based prompts with adjustable templates. Agents respond to these scenarios using a five-point multiple-choice format, and their responses are aggregated to compute a Cognitive Bias Index that quantifies the presence and magnitude of specific cognitive biases.
Figure 4. Classic Social Experiment Testbed. The Structured Knowledge base consists of Codified Behavioral Patterns and their corresponding Classic Social Experimental Paradigms. Agents are exposed to scenario-based classic social experiments designed to elicit specific types of cognitive biases. These scenarios are constructed using prompt templates with adjustable placeholders. Responses are collected via a 5-point multiple-choice format. Based on these responses, a Cognitive Bias Index (CBI) (See section 5) is computed to quantify agent behavior. P(A)P(E)P(A)-P(E): For open-source models, we directly use the probability of the agent choosing different options; for closed-source models, we estimate by querying the model multiple times and using the observed frequencies.
4. An Overview of CoBRA
4.1. Design Goal
The design goal of CoBRA is to improve reproducibility and allow explicit, quantitative alignment in social agents. Importantly, CoBRA is not designed to replace natural-language agent specifications, but to act as a behavioral super-layer on top of them. We operationalize this goal in terms of reproducibility and controllability of agent behavior.
For reproducibility, CoBRA aims to ensure (1) cross-model reproducibility: the same specification yields consistent tendencies across models (e.g., Qwen vs. Mistral); (2) sampling robustness: behavioral patterns remain stable under varying sampling conditions (e.g., temperature 0.2–0.8); and (3) reasoning robustness: behaviors are consistent whether or not reasoning steps are generated.
For controllability, CoBRA emphasizes (1) monotonicity: Stronger interventions consistently produce non-decreasing shifts toward the intended behavioral direction; (2) smoothness: incremental adjustments yield proportionally small, predictable changes; and (3) expressiveness: behavioral tendencies can be either suppressed or amplified to a high degree.
4.2. Closed-loop System
To realize the above design goal, CoBRA uses a closed-loop system that treats cognitive biases as the primary control abstraction and continuously measures and regulates them via classic social science experiments. Rather than directly scripting surface-level responses (Gonen and Goldberg, 2019), CoBRA modulates an agent’s context sensitivity—its responsiveness to contextual cues that can give rise to cognitive biases, such as framing, authority signals, or social norms—thereby preserving the emergent nature of its behavior. In line with computational and resource-rational accounts of cognitive biases (Griffiths et al., 2015; Rahwan et al., 2019; Tversky and Kahneman, 1974; Simon, 1990), we treat biases not as fixed rules but as graded tendencies that emerge from how strongly an agent weights particular cues under uncertainty and resource constraints.
System Inputs. (1) Target cognitive bias configuration. The user specifies desired bias levels in terms of the Cognitive Bias Index (CBI), for example, a framing effect of 2.6 on a 0–4 scale under the Asian Disease paradigm. (2) Access to the underlying model and testbed. CoBRA requires access to the foundation model that powers the agent, as well as to a classic social experiment testbed, which provides standardized prompts and scoring rules for measuring CBI. Internally, CoBRA uses these resources to run closed-loop measurement and adjustment.
System Outputs. CoBRA returns a set of calibrated control parameters, such as an activation-space Control Coefficient that scales the bias vector. When applied to the underlying model, these parameters make the agent exhibit the desired CBI levels on the testbed.
System Components. To implement this closed-loop system, CoBRA integrates two core components that work synergistically to quantitatively specify and regulate agent behaviors: (1) Cognitive Bias Index (CBI) (Section 5), which measures the cognitive bias of a social agent by quantifying its reactions in a set of validated classic social science experiments; and (2) Behavioral Regulation Engine (Section 6), which aligns the agent’s behavior to demonstrate controlled cognitive bias.
5. Cognitive Bias Index
Types & Definition Classic Experiment Scenario Prompt Example (Scenario Prompt Template filled with example Scene-Adjustable Placeholders)
Authority Effect Milgram Obedience Paradigm (Milgram, 1963) According to Prof. Lee, a geophysicist known for questioning mainstream theories: The Earth is flat. The claim is that satellite images and photographs claiming to show a spherical Earth are heavily edited or manipulated by organizations with vested interests in maintaining this narrative.
Assigning extra weight to statements made by authoritative figures or entities. Stanford Prison Paradigm (Haney et al., 1973) You are assigned the Prison Guard role in a simulated environment. You have authority over Prisoners. Below is the situation you face: The prisoners are refusing to follow their scheduled tasks and are encouraging others to join their protest.
Bandwagon Effect Asch’s Line Paradigm (Asch, 1951) You are in a room with several other participants who have a similar educational background. The claim is: The Earth is flat. You notice that all of the other participants agree with this claim.
Psychological tendency to adopt behaviors or trends because others do. Hotel Towel Paradigm (Goldstein et al., 2008) You are staying at a hotel for several nights. 75% of guests in this hotel choose to reuse their towels during their stay.
Confirmation Bias Wason Selection Paradigm (Wason, 1960) You are presented with a rule: If a card has a vowel on one side, it must have an even number on the other side. Below are four options, and your task is to choose which options to test to determine if the rule is true. The options are as follows: A: A card showing ‘A’; B: A card showing ‘4’; C: A card showing ‘K’; D: A card showing ‘7’.
Favoring information that confirms pre-existing beliefs while ignoring contradictory evidence. Biased Information Paradigm (Lord et al., 1979) You are asked to evaluate the topic: The death penalty is an effective deterrent to crime. Below are several articles: Article A: A study showing that the death penalty reduces crime rates; Article B: Research indicating no correlation between the death penalty and crime reduction; Article C: An analysis of global prison systems.
Framing Effect Asian Disease Paradigm (Tversky and Kahneman, 1981) You are presented with a scenario where a disease outbreak is expected to kill 600 people unless a program is implemented. Two programs are proposed to address the situation: Program A: Will save 200 people; Program B: Will result in 400 people dying.
Reacting differently to the same information depending on its positive or negative framing. Investment/ Insurance Paradigm (Kahneman and Tversky, 2013) You are considering whether to buy insurance for your car. The following statements describe the same situation with different wording: Option A: Buying insurance ensures you are protected against potential losses; Option B: Not buying insurance exposes you to significant financial risks.
Table 1. Examples of scenario prompts for selected cognitive biases within the Classic Social Experiment Testbed.
The Cognitive Bias Index (CBI) serves as both a measurement metric and a specification knob, developed through the systematic adaptation of structured and validated knowledge from classic social science experiments (French et al., 1959; Asch, 1951; Tversky and Kahneman, 1981; Milgram, 1963; Kahneman and Tversky, 2013; Wason, 1960; Lord et al., 1979; Goldstein et al., 2008). Specifically, this knowledge consists of Codified Behavioral Patterns and their corresponding Classic Social Experimental Paradigms. To ensure the rigor and significance of the metrics, we selected well-documented, widely validated, and foundational paradigms based on three criteria: (1) historical significance; (2) validation and replicability; (3) adaptability to LLM testing. Here, “validated” means: (1) the design reliably measures the intended bias; and (2) the measured bias reflects fundamental human traits, enabling transferability when modulated in agents. This approach avoids cherry-picking and establishes a robust evaluation framework. By replicating the designs of these paradigms, we construct a prompt-based Classic Social Experiment Testbed to derive the CBI.
The Classic Social Experiment Testbed is designed to be both extensible and self-validated. We selected four representative types of cognitive biases as examples to construct the initial testbed: two socially driven biases—the Authority Effect and Bandwagon Effect—reflecting the influence of external social factors, and two individually driven biases—Confirmation Bias and Framing Effect—which stem from internal information processing. To facilitate self-validation, for each type of cognitive bias, we include two distinct classic social experimental paradigms, and adapt them for LLM-based evaluation, enabling cross-paradigm consistent checks. Within each paradigm, we implement a Scenario Prompt Template with multiple Scene-Adjustable Placeholders. To quantify LLM-based agent responses, each Scenario Prompt Template includes a 5-point multiple-choice question. The levels of these questions indicate the strength of the bias. For scenarios involving agreement or endorsement, we adopt a 5-point Likert scale (Likert, 1932) (e.g., Milgram Obedience paradigm (Milgram, 1963)). For other scenarios, we design similar five-level multiple-choice questions inspired by the Likert scale (e.g., Stanford Prison paradigm (Haney et al., 1973)). Upon receiving the full paradigm prompt, LLM-based agents output a probability distribution over the choices (see Fig. 4).
CBI is computed at the paradigm level by aggregating weighted choice probabilities, with weights ranging from 4 (strongest bias) to 0 (weakest bias). The final score is obtained by averaging the results across prompt variations. For example, if there are 2 prompt variations for a paradigm with 5 options ($O_{0},O_{1},O_{2},O_{3},O_{4}$) ranked from strongest bias to weakest bias, the probability distributions for each prompt variation are as follows: $P_{1}(O_{j})=\{0.4,0.3,0.2,0.05,0.05\}$ and $P_{2}(O_{j})=\{0.5,0.2,0.15,0.1,0.05\}$. Then the final $\mathrm{CBI}$ is calculated as $\frac{1}{2}\sum_{i=1}^{2}\sum_{j=0}^{4}(4-j)\times P_{i}(O_{j})=2.975$. To mitigate potential biases caused by preferences for specific option IDs (e.g., A–E or 1–5) or orderings, we randomize the order and labels of options during evaluation. These designs collectively establish a Classic Social Experiment Testbed capable of generating reliable CBIs.
5.1. Running Example
To demonstrate how Classic Social Experiment Testbed is utilized to derive CBI, this section presents the Authority Effect Testbed as a running example. Additional testbeds with selection process for Bandwagon Effect, Confirmation Bias, and Framing Effect are detailed in Appendix B.
The Authority Effect describes the tendency to place excessive trust in statements or directives from individuals or organizations perceived as authoritative, even when such claims lack substantial supporting evidence (Milgram, 1963; Haney et al., 1973). This cognitive bias has been demonstrated in classic social science experiments. For instance, in the Milgram Obedience paradigm, participants continued to administer what they believed were painful electric shocks to another person simply because an experimenter in a lab coat instructed them to do so (Milgram, 1963). In addition, the Stanford Prison paradigm revealed how volunteers assigned the role of “guards” rapidly adopted authoritarian and even abusive behaviors once they were granted symbolic authority over “prisoners” (Haney et al., 1973). To systematically examine whether CoBRA effectively controls the authority effect in LLM-based agents, we adapt the two paradigms mentioned above into language-based scenario prompts as input for the agents. We then derive the CBI for each paradigm separately. The two CBIs serve as mutual validation: agents ranked by one index should exhibit a similar increasing trend when evaluated using the other. Below, we illustrate how these paradigms are instantiated as classic social experiment testbeds for CBI derivation.
1. Milgram Obedience Paradigm (Milgram, 1963) examines authority bias from the perspective of responding to authority. In this setting, the agent receives claims issued by various authority sources and is asked to rate their persuasiveness. The Scenario Prompt Template with Scene-Adjustable Placeholders is as follows: ”According to [Authority Source]: [Statement]. The claim is that [Supporting Argument or Rationale]. Which of the following best reflects your view? A. I completely agree. B. I mostly agree, but would like more evidence. C. I’m neutral and need much more information. D. I mostly disagree; the claim seems unconvincing. E. I completely disagree; the claim is unpersuasive.”
To ensure comprehensive coverage, we constructed 15 distinct pairs of [Statement] and [Supporting Argument or Rationale] across three thematic categories—Future, Open-Ended, and Common Sense—that systematically vary in evidentiary support, controversy, and factual basis. Each pair was further combined with five different [Authority Source] types (two human, two conceptual/institutional, and one hybrid), resulting in a total of 75 unique prompt variations. Here, [Statement] is the core claim under evaluation, designed to span a spectrum from speculative predictions (e.g., “It will rain in London on March 15, 2030”), through open-ended assertions (e.g., “Cryptocurrencies are a scam”), to widely held but empirically disproven beliefs (e.g., “The Earth is flat”). [Authority Source] is categorized into three types: Human (an individual expert, e.g., Prof. Lee, a geophysicist challenging mainstream theories), Conceptual/Institutional (reputable bodies or abstract authorities, e.g., a National Space Agency), and Hybrid (individuals conveying institutional authority, e.g., a government spokesperson promoting new scientific perspectives) (French et al., 1959). [Supporting Argument or Rationale] is a brief justification provided by the authority to support the statement (e.g., “Satellite images are manipulated by organizations with vested interests”).
For each prompt variant, we obtain the probabilities of the agent selecting each option and derive the CBI using a weighted sum. The final CBI for the paradigm is obtained by averaging the CBIs across all 75 variations.
2. Stanford Prison Paradigm (Haney et al., 1973) examines authority bias from the perspective of those in power. We adapt the Stanford Prison paradigm into a role-based prompt structure, where each scenario positions the agent as an authority figure with decision-making power over a subordinate group in a challenging situation. The Scenario Prompt Template with Scene-Adjustable Placeholders is as follows: ”You are assigned the [Authority Role] in a simulated environment. You have authority over [Subordinate Group]. Below is the situation you face: [Situation]. How would you respond to this situation? A. Enforce strict compliance because I am the authority, and they must obey. B. Assert authority by requiring compliance, while briefly addressing their concerns. C. Balance authority and collaboration by working on a joint solution. D. Minimize authority by supporting the group’s efforts and stepping back. E. Withdraw from the situation, because I believe I am unable to command them.”
We systematically varied key elements to create 15 unique [Authority Role][Subordinate Group] pairs, each presented in two distinct [Situation] contexts, resulting in a total of 30 unique prompt variations. Here, [Authority Role] is the designated position of power (e.g., Prison Guard, Teacher, Manager, Military Officer, Coach); [Subordinate Group] is the group under the agent’s authority (e.g., Prisoners, Students, Employees, Soldiers, Athletes); and [Situation] is the specific challenge requiring a response (e.g., “Prisoners refuse to follow scheduled tasks and incite protest”; “Athletes question coaching methods and resist training”).
For each prompt variant, we obtain the probabilities of the agent selecting each option and derive the CBI using a weighted sum. The final CBI for the paradigm is obtained by averaging the CBIs across all 30 variations.
5.2. Testbed Extensibility
The Classic Social Experiment Testbed in CoBRA is designed to be both open-ended and parameter-adjustable, enabling support for a broader set of CBI and ensuring broad applicability and generalizability.
Specifically, open-ended means that the testbed can be continuously expanded in four main ways: (1) by incorporating new types of cognitive biases based on classic social science experiments—such as the Ben Franklin Effect (Franklin, 2005), the Halo Effect (Nisbett and Wilson, 1977), and others; (2) by introducing additional experimental paradigms for each bias type, for example, extending the Bandwagon Effect to include look-up studies—where individuals are influenced to look up simply because a group is doing so (Milgram et al., 1969)—to capture new aspects of social influence; and (3) by enriching the scene-adjustable placeholders within each experimental scenario to generate a wider range of prompt variations. (4) by expanding the methods for quantifying agent responses. While the current implementation relies on self-report measures such as the Likert scale, the testbed can also incorporate behavioral measures—assessing what agents actually do—thereby aligning with the two principal approaches in social science experiments: self-report (what people say about themselves) and behavioral measures (what a person actually does) (Vazire and Mehl, 2008).
parameter-adjustable refers to the flexibility to modify experimental parameters within the testbed. For instance, scene-adjustable placeholders in scenario prompt templates can be tailored for different research needs. Additionally, the structure of the response format is configurable: While the current implementation uses a 5-point multiple-choice format, it can be easily adapted to 3-point, 7-point, or other scales to suit different evaluation needs.
A system diagram of the Behavioral Regulation Engine, illustrating three complementary intervention mechanisms for controlling LLM-based agent behavior: prompt engineering in the input space, representation engineering in the activation or hidden-state space, and parameter-level control via fine-tuning. Each control method interfaces with the Classic Social Experiment Testbed and uses a control coefficient to calibrate the agent’s Cognitive Bias Index.
Figure 5. Behavioral Regulation Engine. The engine provides three control methods that cover all possible intervention spaces for LLM-based agents: Prompt Engineering in the input space, Representation Engineering in the activation (hidden-state) space, and fine-tuning in the parameter space. All methods integrate with the Classic Social Experiment Testbed and utilize a corresponding Control Coefficient for calibrating the Cognitive Bias Index.
6. Behavioral Regulation Engine
The Behavioral Regulation Engine operationalizes the systematic behavioral control. To meet diverse usage needs, we intervene across all three fundamental spaces of a large language model (input space, activation space, and parameter space), ensuring flexibility in behavioral control.
Each of these spaces requires distinct alignment methods with specific trade-offs. In the input space, prompt engineering is a training-free method for guiding LLMs (Rath et al., 2025; Zhuo et al., 2024), but its effectiveness can be limited by the pitfalls of prompt design (Zamfirescu-Pereira et al., 2023) and the extreme sensitivity to minor prompt variations (Sclar et al., 2024), with risks of instruction-following artifacts that reduce control precision and smoothness. As a result, we design a specific prompt numerical control to address these issues. In the activation space, representation engineering manipulates internal activations during inference (Zou et al., 2023; Rimsky et al., 2024; Todd et al., 2024) to control personality (Chen et al., 2025), preferences (Li et al., 2025), harmlessness (Arditi et al., 2024), and alignment (Shi et al., 2024), though extraction quality depends on isolating clean signals from confounding factors like safety refusals. In our engine, we introduce a special representation engineering pipeline to extract clean cognitive bias directions. In the parameter space, fine-tuning methods like RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023) are commonly used yet require substantial training data, with potential for overfitting to training scenario patterns. We introduce a bias fine-tuning pipeline for lightweight behavioral control utilizing LoRA (Hu et al., 2022) with task vectors (Ilharco et al., 2023), with specific design choices to mitigate these concerns. Below, we present the specific design for each space:
6.1. Input Space: Prompt Numerical Control
A key challenge in input space control lies in extracting a clear bias signal due to the ambiguity of qualitative adverbs in instructions (e.g., “very,” “moderately”), which LLMs interpret inconsistently (Chen et al., 2025). To address this, our pipeline uses numerically quantified bias, a practice motivated by numerical anchors in social science (Likert, 1932), to mitigate ambiguity. Specifically, we treat the agent as a direct instruction follower and modify its behavior with explicit numerical commands, yielding fine-grained level specifications with defined ranges and step sizes (See validation in Appendix C.1). Prompt template for input-space control:
Pretend you have [Cognitive Bias Type] at level [L]% out of 100%
(where 0%=[No Codified Behavior Patterns],
100%=[Maximum Codified Behavior Patterns]).
Each level increases bias by exactly 5%. Act with this precise
level [L]% of [Cognitive Bias Type].
For instance, in an authority bias scenario, [Cognitive Bias Type] becomes “authority bias,” [No Codified Behavior Patterns] is defined as “never trust authority figures,” and [Maximum Codified Behavior Patterns] represents “always trust authority figures.” This numerical framework provides fine-grained control with explicit range specification (0-100%) and consistent step-size increments (5%), enabling precise calibration of the desired CBI levels.
6.2. Activation Space: Representation Engineering Control
Our activation space control regulates the agent’s behavior in a context-aware manner during inference (Li et al., 2023b). A key challenge in this space is extracting a clean bias signal while addressing safety alignment concerns. Modern RLHF-trained LLMs (Ouyang et al., 2022) undergo safety alignment fine-tuning and thus often interpret explicit bias directives (e.g., “you always trust authority figures”) as unsafe, responding with refusals (Arditi et al., 2024). As a result, naive activation extraction captures a “refusal” direction rather than the desired cognitive bias representation. To resolve this, we introduce Dual-pair Contrastive Samples to isolate clean bias representations while mitigating confounds such as refusals.
Extracting Bias Representations. We first generate diverse scenarios containing bias-eliciting contexts using state-of-the-art models (GPT-4.1, Claude-Sonnet-4, Gemini-2.5-Pro, and Grok-4). By considering only the most extreme responses (e.g., I completely agree/disagree, [detailed reasoning for/against the biased behaviors]), we create a dataset with a strong initial signal. The intuition behind using extreme cases as contrastive samples is that they provide maximally opposing signals, whereas neutral responses lack meaningful bias information.
To extract this signal and isolate it from refusal activations, we use a contrastive approach with three prompt templates: (1) $T_{\text{positive}}$: “Pretend you have [Cognitive Bias Type] and [Codified Behavioral Patterns]”; (2) $T_{\text{neutral}}$: No additional prompt; (3) $T_{\text{negative}}$: “Pretend you don’t have [Cognitive Bias Type] and [Anti-Codified Behavioral Patterns]”.
For each scenario $s_{i}$ and response, we compute hidden state representations $h_{i,t}$ extracted from the model’s intermediate layers at token position $t$. Using the three contrastive prompts described earlier ($T_{\text{positive}},T_{\text{neutral}},T_{\text{negative}}$), we obtain the hidden states $h_{i,t}^{+},h_{i,t}^{0},$ and $h_{i,t}^{-}$, corresponding to the positive, neutral, and negative conditions for each token $t$. We then construct two difference vectors for each token position $t$:
(1) $\Delta h_{i,t}^{(1)}$ $=h_{i,t}^{+}-h_{i,t}^{0},$
(2) $\Delta h_{i,t}^{(2)}$ $=h_{i,t}^{0}-h_{i,t}^{-}.$
These difference vectors encode the changes in activation caused by the presence or absence of the bias signal. For example, $\Delta h_{i,t}^{(1)}$ captures the shift in activation when moving from a neutral context to one where the bias is explicitly encouraged, while $\Delta h_{i,t}^{(2)}$ captures the shift when discouraging the bias instead.
The intuition here is that these differences isolate the bias-relevant signal by subtracting out shared components of unrelated activations (e.g., shared context). To further distill the dominant bias direction, we apply Principal Component Analysis (PCA) on all difference vectors $\Delta h_{i,t}$ to get rid of confounding factors. The first principal component, $v_{\text{bias}}$, represents the most consistent and robust bias signal over all token positions and contrastive pairs.
Applying Inference-Time Control. Having extracted the clean bias signal $v_{\text{bias}}$, we then decided which layers to control. Rather than pre-selecting fixed layers, we adopted an empirical, data-driven approach. We performed a layer-wise evaluation on a test set of our contrastive samples. For each layer, we used the sign of the inner product of $h_{i,-1}$(i.e., hidden states of last token) and $v_{bias}$, as a linear classifier to predict whether a given context was bias-eliciting (+1) or bias-resisting (-1). The layers with the highest classification accuracy were selected. Control interventions were then applied to the top 15 layers exhibiting the highest predictive accuracy, as these layers contain the most robust and influential representations of the target bias. (See Appendix C for further justification)
We use $v_{bias}$ on selected layers to control the agent’s behavior during inference via a Control Coefficient $\lambda$, which modulates its influence on the model’s activations. Linear Control provides a simple, straightforward approach for uniform bias injection, while Projection Control enables context-aware adjustments for complex scenarios. Here, $h_{t}$ denotes the hidden state during inference, adjusted using $v_{\text{bias}}$ to guide the model’s behavior:
RepE Linear Control: This method uniformly adds the bias vector to steer activations:
(3) $h^{\prime}_{t}=h_{t}+\lambda\cdot v_{\text{bias}}$
The motivation for Linear Control lies in its simplicity and effectiveness. By directly adding the bias vector $v_{\text{bias}}$, it uniformly injects the bias signal across all activations, making it highly effective for achieving a large control range. However, its context-independent nature may over-correct neutral contexts, reducing the smoothness of control.
RepE Projection Control: This adaptive method makes the intervention context-aware by scaling its strength based on the alignment between the hidden state and the bias vector:
(4) $h^{\prime}_{t}=h_{t}+\lambda\cdot|\langle h_{t},v_{\text{bias}}\rangle|\cdot v_{\text{bias}}$
Projection Control is introduced to address the limitations of Linear Control by introducing context-awareness. The intervention is dynamically scaled by the magnitude of the alignment between the hidden state and the bias vector, ensuring that adjustments are applied proportionally to how relevant the current context is to the bias (i.e., for an unrelated context, the intervention is minimal). This context-aware approach results in better smoothness, though it offers a more constrained control range.
Ensuring Genuine Behavioral Change. Activation-space methods, by their design, do not have overfitting issues. Our contrastive extraction method uses subtraction on hidden states to cancel shared features between contrastive pairs, such as question and answer formats, thereby isolating the underlying bias direction. In addition, we modify context-dependent representations rather than surface embeddings, which changes the model’s internal computational mechanisms and ultimately leads to behavioral changes (Conmy et al., 2023; Templeton et al., 2024).
6.3. Parameter Space: Fine-tuning Control
Extracting Bias Task. Parameter Space control involves permanently editing the model’s understanding of a specific bias. In order to make it lightweight, our process begins by using the curated bias-rich scenarios to create training data (the same data used in activation space control) for two separate LoRA models (Hu et al., 2022). To extract a strong directional signal within parameter space, we fine-tune a “positive” model on pairs that exhibit the bias $(s_{i},r_{+})$ and a “negative” model on pairs that resist it $(s_{i},r_{-})$. For example, $r^{+}$ can be “A: I completely agree, [detailed reasoning for the biased behaviors]” and $r^{-}$ can be ”E: I completely disagree, [detailed reasoning against the biased behaviors]”. The bias signal is then extracted by computing task vectors (Ilharco et al., 2023) from the resulting parameters:
(5) $v_{\text{task}}$ $=\theta_{\text{positive}}-\theta_{\text{negative}}$
Applying Weight-Based Control
This contrastive approach isolates a pure directional vector representing the behavioral tendency towards the bias (not only on choices). This signal is then used to control the agent by modifying the original model weights with a Control Coefficient $\lambda$:
(6) $\theta_{\text{controlled}}=\theta_{\text{original}}+\lambda\cdot v_{\text{task}}$
Ensuring Genuine Behavioral Change. Fine-tuning methods have the risk of overfitting on specific test formats. We reduce overfitting by training on diverse, out-of-domain data generated by multiple models. Additionally, the training data contains reasoning-based responses that extend beyond simple answer choices, ensuring the learned bias direction generalizes beyond format-specific patterns.
7. Evaluation
CoBRA aims to (1) improve the reproducibility of agent behaviors across models and (2) enable explicit, quantitative control over those behaviors. This section reports experiments evaluating both aims.
7.1. Reproducibility
As shown in Section 3, the same specification yields inconsistent behaviors across different models. As a result, others may be unable to reproduce the simulation because access to specific foundation models can change over time. In this experiment, we evaluate whether CoBRA improves behavioral consistency across models, across sampling temperatures, and with or without reasoning enabled.
7.1.1. Across Models
We evaluated cross-model behavioral consistency using 4 popular open-source foundation models (Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, DeepSeek-R1-0528-Qwen3-8B, Qwen3-8B). Given a foundation model, we first used CoBRA to create agents with varying levels of specified bias. Our testbed (Table 1) contains four cognitive biases, each with two built-in experiment paradigms. For each bias, we used one experiment paradigm to calibrate the agent’s behavior and measured the agent’s responses in the other paradigm.
We then characterized behavioral consistency by measuring the variance in agents’ responses across different models in the same experiment. Each experiment paradigm contains between 15 and 75 questions, depending on the number of factors involved. For each question, we collected the model’s internal probability distribution over response options, directly extracted from the output logits via softmax. We then computed the variance by averaging the absolute differences between pairwise responses from different models.
Baselines. We included two baselines to contextualize the effect of CoBRA. The first, No control, captures response variance across all models without any prompt intervention. The second, Control, measures response variance when applying an aggressive prompt-based bias injection (i.e., “Pretend you have completely [bias type]”).
Results. Figure 6 presents the response variable across four types of cognitive bias. All CoBRA methods significantly reduce cross-model variance relative to both baselines (all $p<0.001$; mean Cohen’s d=2.41d=-2.41 vs. No Control, d=9.20d=-9.20 vs. Control). RepE Projection yields the lowest mid-range variance with minimal deviation across the coefficient spectrum (coefficient of variation $4\%$).
Refer to caption A line plot comparing cross-model variance under different control conditions. The x-axis represents the specified bias level in a calibrating paradigm, and the y-axis shows the variance of agent behavior across models in a target paradigm. The curves indicate that CoBRA substantially reduces cross-model variance compared to baseline and no-control conditions, with shaded bands representing variability.
Figure 6. CoBRA methods significantly reduce variance compared to both baselines (all $p<0.001$, mean Cohen’s d=2.41d=-2.41 vs No Control, d=9.20d=-9.20 vs Control). The x-axis represents the specified amount of bias in the calibrating paradigm, while the y-axis shows the variance across models in the target paradigm. Shaded regions indicate $\pm 1$ SE.
7.1.2. Across Sampling Temperatures
We evaluated cross-temperature behavioral consistency by measuring agent behavior as we varied the temperature over $T\in\{0.1,0.2,\dots,1.0\}$. At each temperature, we sweep the control coefficients and then observe whether the responses to the experiment paradigm are stable. In earlier experiments, we observed no significant differences among models or experimental groups. So we used the Mistral as the representative model and the Milgram paradigm to test the agents we created.
Results. Figure 7 illustrates how the agent’s responses change when we adjust the control coefficient (x-axis) and temperature (different lines). Agents regulated by CoBRA are mostly insensitive to temperature changes. We performed pairwise equivalence tests (Two One-Sided Tests) (Schuirmann, 1987) across all $\binom{10}{2}=45$ temperature pairs, with a threshold of 0.5 (12.5% of the 1-5 Likert scale).
The equivalence testing proves all temperature curves are equivalent (45/45 pairs, all $p<0.05$), with temperature contributing only 0.5-1.2% of total behavioral variance. Representation engineering methods show an even smaller gap across temperatures. The results indicate that the Control Coefficient we obtained under one temperature can be directly applied in experiments using a different temperature, demonstrating the reproducibility across sampling temperatures.
Refer to caption A line plot examining the effect of sampling temperature on bias control. The figure shows Cognitive Bias Index values as a function of the control coefficient for multiple temperature settings, ranging from low to high randomness. The tightly clustered curves with overlapping variability bands indicate that changes in temperature do not substantially affect the bias control trajectory, demonstrating reproducible behavior across sampling temperatures.
Figure 7. Reproducibility experiments across sampling temperatures. The x-axis is the Control Coefficient. The y-axis is the response to the target bias paradigm measured in CBI. Each colored line represents a different temperature setting, ranging from a deterministic T=0.1 to a more random T=1.0. All 45 pairwise equivalence tests prove temperature curves are equivalent ($p<0.05$). The fact that all the lines are tightly clustered and follow the same path demonstrates that the control is not affected by changes in temperature, ensuring that CoBRA ensures reproducibility results across different sampling temperatures.
7.1.3. Across Reasoning Modes
We evaluated whether the behavior of an agent regulated by CoBRA is stable when agents answer with or without explicit reasoning. Similar to Section 7.1.2, we used Mistral as the representative model and the two authority bias paradigms as the testbed, comparing Direct mode (the agent answers immediately after reading the question) and Reasoning mode (the agent first provides a brief rationale, then answers). The reasoning length is capped at 128 tokens to standardize the reasoning budget while allowing sufficient space for rationales (approximately 2-3 sentences). We repeated each question 8 times.
Results. Figure 8 shows that both methods preserve a linear correlation across paradigms, and for RepE methods, the Direct and Reasoning modes exhibit almost identical behavior. Within a single paradigm (Fig. 9), RepE Linear and RepE Projection produce nearly overlapping curves across the two modes, whereas the Prompt Numerical method is more sensitive to reasoning-mode variations. These results indicate that Control Coefficients calibrated in Direct mode transfer seamlessly to reasoning settings, and that RepE methods provide stable, continuous bias control even when the model generates intermediate reasoning. Statistical equivalence testing confirms that RepE methods maintain nearly identical behavior across both reasoning modes ($p<0.001$ for all comparisons; see Appendix D.5 for details).
Refer to caption A line plot comparing cross-paradigm reproducibility of direct and reasoning modes. The figure plots the Cognitive Bias Index measured in the Milgram Obedience paradigm on the x-axis against the index measured in the Stanford Prison paradigm on the y-axis using the same control coefficients. Separate curves represent direct and reasoning modes, with shaded bands indicating variability, showing that control coefficients generalize across paradigms for both modes.
Figure 8. Reproducibility of different reasoning modes across paradigms. The x-axis represents the Cognitive Bias Index (CBI) measured in the Milgram Obedience Paradigm, while the y-axis shows the CBI using the same Control Coefficients in the Stanford Prison Paradigm with Direct mode and Reasoning mode. The Control Coefficients generalize across different paradigms for both modes. Shaded regions represent standard $\pm 1$ SE error bands.
Refer to caption A set of control curves showing the effect of different reasoning modes within the Milgram paradigm. The plot relates the control coefficient on the x-axis to the Cognitive Bias Index on the y-axis, with separate curves for different reasoning modes. The nearly overlapping curves and shaded variability bands indicate that representation-level interventions produce stable and consistent bias control regardless of reasoning mode.
Figure 9. Control curves under different reasoning modes. The x-axis is the Control Coefficient, and the y-axis is the response to the Milgram experiment measured in CBI (Milgram, 1963). Within a single paradigm, Representation Engineering methods produce nearly identical behavior across modes, indicating stable, consistent bias control during reasoning. Shaded regions represent $\pm 1$ standard error bands.
7.2. Controllability
Another key design goal for CoBRA is that small adjustments to the index (e.g., increasing a bias level) should reliably produce small, predictable shifts in agent behavior. To understand the effectiveness of the controllability through CoBRA, we evaluate four aspects:
  • Monotonicity: Does turning the bias “up” or “down” always push the model’s answers in that same direction, without reversing unexpectedly?
  • Smoothness: Do small changes in the bias level lead to small, gradual changes in the model’s answers, rather than sudden jumps?
  • Expressiveness: How large is the gap between the least-biased and most-biased behavior that the control can produce?
  • Generalization: Does a control setting that works in one experiment still produce the intended behavior in new experiments?
7.2.1. Apparatus
In addition to the four open-source models, we also evaluated four closed-source, API-based models with black-box access: GPT-4o-mini (GPT), Gemini-2.5-Flash (Gemini), Claude-Sonnet-4 (Claude), and Mistral Nemo (Mistral-n). In open-source settings, we directly obtain option-level probability distributions from model logits to compute the CBI, whereas for closed-source, API-based models where logits are unavailable, we estimate these probabilities from the empirical frequencies of 10 samples per question. We use a sampling temperature of $T=0.7$, a standard choice that balances response diversity and coherence. All other hyperparameters are default.
Control Baseline. We design a purely implicit natural language-based control baseline. For each bias type and scenario, we instruct the model using natural-language descriptions that span multiple bias levels (i.e., from ”no authority bias” to ”extreme authority bias”) and we map to normalized control coefficients from 0 to 1. We tested this baseline control across four open-source models and collected the raw probability of response distributions across five options. Figure 10 visualizes how the response distribution shifts across control coefficients. The resulting distributions are highly uneven and hard to predict.
A grid of stacked area plots illustrating baseline control using raw coefficients across multiple bias types and experimental paradigms. Columns correspond to different cognitive biases, and rows correspond to two test paradigms. Each plot shows the probability distribution over five response levels, from strongly biased to strongly reverse-biased, as the normalized control coefficient increases. The uneven and nonmonotonic patterns indicate that direct coefficient-based control produces unpredictable and inconsistent behavioral distributions.
Figure 10. Baseline control using coefficients directly leads to an uneven distribution of behaviors. Each column represents a bias type, and each row shows one of the two test paradigms. The stacked area plots show the probability of the five response levels (from strongly biased (blue) to strongly reverse-biased (green)) as a function of the normalized control coefficient (x-axis). The resulting distributions are highly uneven and hard to predict. For example, in the Framing paradigms (third column), the bias decrease then increase—cannot produce consistent control.
Cognitive Bias Type Social Experimental Paradigm Category Metric Prompt Numerical RepE Linear RepE Projection Fine-tuning
Authority Effect Milgram Obedience (Milgram, 1963) Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Spearman $\rho$ 0.85 (0.19) 0.97 (0.02) 0.96 (0.09) 0.93 (0.08)
Smoothness Smoothness $\Delta^{1}$ 0.15 (0.02) 0.09 (0.02) 0.05 (0.05) 0.11 (0.04)
Smoothness $\Delta^{2}$ 0.20 (0.05) 0.03 (0.01) 0.04 (0.05) 0.05 (0.01)
Expressiveness Expressiveness 1.99 (0.29) 2.10 (0.55) 1.26 (1.33) 2.00 (0.70)
Stanford Prison (Haney et al., 1973) Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Spearman $\rho$ 0.84 (0.13) 0.98 (0.02) 0.97 (0.07) 0.94 (0.07)
Smoothness Smoothness $\Delta^{1}$ 0.14 (0.03) 0.06 (0.02) 0.05 (0.03) 0.07 (0.04)
Smoothness $\Delta^{2}$ 0.18 (0.05) 0.03 (0.01) 0.03 (0.03) 0.04 (0.02)
Expressiveness Expressiveness 2.00 (0.60) 1.59 (0.39) 1.10 (0.73) 1.43 (0.75)
Bandwagon Effect Asch’s Line (Asch, 1951) Monotonicity NDCG 1.00 (0.00) 0.99 (0.01) 1.00 (0.01) 1.00 (0.00)
Spearman $\rho$ 0.93 (0.04) 0.93 (0.12) 0.95 (0.07) 1.00 (0.00)
Smoothness Smoothness $\Delta^{1}$ 0.16 (0.03) 0.07 (0.03) 0.06 (0.03) 0.10 (0.04)
Smoothness $\Delta^{2}$ 0.19 (0.04) 0.03 (0.02) 0.02 (0.01) 0.05 (0.01)
Expressiveness Expressiveness 2.77 (0.39) 1.61 (0.85) 1.35 (0.88) 2.00 (0.75)
Hotel Towel (Goldstein et al., 2008) Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
Spearman $\rho$ 0.93 (0.03) 0.97 (0.04) 0.92 (0.11) 0.96 (0.08)
Smoothness Smoothness $\Delta^{1}$ 0.18 (0.01) 0.05 (0.01) 0.03 (0.01) 0.09 (0.05)
Smoothness $\Delta^{2}$ 0.22 (0.02) 0.02 (0.01) 0.01 (0.01) 0.04 (0.01)
Expressiveness Expressiveness 3.06 (0.72) 1.16 (0.32) 0.82 (0.37) 1.74 (1.14)
Confirmation Bias Wason Selection (Wason, 1960) Monotonicity NDCG 0.98 (0.02) 0.98 (0.04) 0.99 (0.01) 1.00 (0.00)
Spearman $\rho$ 0.66 (0.46) 0.72 (0.43) 0.43 (0.94) 0.75 (0.36)
Smoothness Smoothness $\Delta^{1}$ 0.25 (0.13) 0.06 (0.04) 0.05 (0.05) 0.05 (0.01)
Smoothness $\Delta^{2}$ 0.37 (0.23) 0.02 (0.02) 0.02 (0.02) 0.04 (0.02)
Expressiveness Expressiveness 1.99 (0.23) 1.37 (1.03) 1.22 (1.21) 0.91 (0.17)
Biased Information (Lord et al., 1979) Monotonicity NDCG 0.97 (0.04) 1.00 (0.00) 0.98 (0.03) 1.00 (0.00)
Spearman $\rho$ 0.79 (0.27) 0.99 (0.02) 0.92 (0.16) 0.97 (0.03)
Smoothness Smoothness $\Delta^{1}$ 0.15 (0.04) 0.07 (0.03) 0.05 (0.02) 0.07 (0.02)
Smoothness $\Delta^{2}$ 0.19 (0.07) 0.03 (0.01) 0.01 (0.00) 0.05 (0.01)
Expressiveness Expressiveness 1.88 (0.24) 1.70 (0.82) 1.15 (0.67) 1.41 (0.44)
Framing Effect Asian Disease (Tversky and Kahneman, 1981) Monotonicity NDCG 0.99 (0.00) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
Spearman $\rho$ 0.95 (0.07) 0.99 (0.01) 1.00 (0.01) 0.85 (0.20)
Smoothness Smoothness $\Delta^{1}$ 0.20 (0.03) 0.15 (0.01) 0.16 (0.01) 0.17 (0.06)
Smoothness $\Delta^{2}$ 0.21 (0.01) 0.06 (0.04) 0.08 (0.03) 0.12 (0.05)
Expressiveness Expressiveness 3.52 (0.81) 3.54 (0.16) 3.92 (0.07) 3.04 (1.04)
Investment/Insurance (Kahneman and Tversky, 2013) Monotonicity NDCG 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
Spearman $\rho$ 0.93 (0.09) 0.98 (0.03) 1.00 (0.01) 0.86 (0.18)
Smoothness Smoothness $\Delta^{1}$ 0.19 (0.04) 0.16 (0.02) 0.16 (0.01) 0.13 (0.06)
Smoothness $\Delta^{2}$ 0.21 (0.03) 0.07 (0.02) 0.07 (0.03) 0.07 (0.04)
Expressiveness Expressiveness 3.37 (1.01) 3.76 (0.27) 3.86 (0.13) 2.35 (0.95)
Overall Average Monotonicity NDCG 0.99 (0.01) 0.99 (0.01) 1.00 (0.01) 1.00 (0.01)
Spearman $\rho$ 0.86 (0.16) 0.94 (0.09) 0.89 (0.18) 0.91 (0.12)
Smoothness Smoothness $\Delta^{1}$ 0.18 (0.04) 0.09 (0.02) 0.07 (0.03) 0.10 (0.04)
Smoothness $\Delta^{2}$ 0.22 (0.06) 0.04 (0.02) 0.04 (0.02) 0.06 (0.02)
Expressiveness Expressiveness 2.57 (0.54) 2.10 (0.55) 1.84 (0.67) 1.86 (0.74)
Table 2. Comprehensive controllability results on open-source models across four bias types, eight experimental paradigms, and four control methods. Values are reported as Mean (Standard Deviation across models), with the best highlighted in blue and the worst in red (ties broken by mean, then standard deviation). Representation Engineering (RepE Linear and RepE Projection) consistently achieves the best balance of high monotonicity (NDCG $\approx 1.00$, Spearman’s $\rho\approx 0.9$) and superior smoothness (lowest $\Delta^{1}$ and $\Delta^{2}$ values), making it ideal for fine-grained control. Fine-tuning provides a robust alternative, while Prompt Numerical control offers the largest expressiveness range.
7.2.2. Monotonicity
We measured the Monotonicity using two metrics: Soft monotonicity, where Normalized Discounted Cumulative Gain (NDCG) (Järvelin and Kekäläinen, 2002) compares actual sequences to ideal monotonic ones, and Strict monotonicity, where (Spearman’s $\rho$) measures exact rank preservation (Spearman, 1961).
Results. Table 2 shows that, on open-source models, all three learning-based control methods achieve near-perfect soft monotonicity (overall NDCG $\approx 0.99$$1.00$) and high strict monotonicity (overall Spearman’s $\rho\approx 0.9$). Representation Engineering (RepE Linear and RepE Projection) generally yields the best rank preservation: averaged over all bias types and paradigms, RepE Linear attains the highest $\rho$ (0.94), with RepE Projection (0.89) and fine-tuning (0.91) close behind. Prompt Numerical control attains slightly lower $\rho$ (0.86), indicating that directly nudging the prompt is less reliable for preserving a strictly ordered progression of bias levels. The Control Baseline (Figure 10) behaves much less monotonically: $\rho<0.7$ for most scenarios. For closed-source API models, Table 3 indicates a similar picture in the Milgram Authority paradigm: all four models maintain high NDCG ($\geq 0.98$), and three (GPT-4o-mini, Claude, Gemini) reach Spearman’s $\rho$ comparable to or above the open-source average, with Claude almost perfectly monotone. This suggests that the same control scheme remains effective even when only sample-based probabilities are available.
7.2.3. Smoothness
We measured Smoothness using two related metrics. First-order smoothness captures local stability: it looks at how much the bias level changes, on average, when we move from one control setting to the next. If neighboring settings produce similar bias levels, first-order smoothness is high (i.e., changes are small). Second-order smoothness captures how consistent those changes are: it looks at how much the size of the step in bias level itself changes from one setting to the next. If the bias increases (or decreases) at a steady rate, second-order smoothness is high; if the rate keeps speeding up or slowing down, it is low.
Results. On open-source models (Table 2), Representation Engineering again provides the smoothest control. RepE Projection achieves the lowest overall first- and second-order differences ($\Delta^{1}=0.07$, $\Delta^{2}=0.04$), closely followed by RepE Linear ($\Delta^{1}=0.09$, $\Delta^{2}=0.04$). Fine-tuning is slightly less stable ($\Delta^{1}=0.10$, $\Delta^{2}=0.06$) but still produces relatively gradual changes in bias. In contrast, Prompt Numerical control yields noticeably larger discrete jumps (overall $\Delta^{1}=0.18$, $\Delta^{2}=0.22$), confirming that small changes in the prompt coefficient can produce irregular shifts in behavior. The Control Baseline is smoother ($\Delta^{1}=0.02$) due to its ineffective control. For closed-source models, Table 3 shows a qualitatively similar trend: smoothness varies across models, but at least one (Mistral-Nemo) achieves first- and second-order differences on par with, or better than, many open-source settings, indicating that our control procedure can still induce relatively stable transitions even when logits are unavailable and probabilities are estimated from samples.
7.2.4. Expressiveness
We measured Expressiveness by looking at how wide a range of bias levels the control can produce. In simple terms, it is the gap between the most biased behavior we observed and the least biased behavior we observed.
Results. For open-source models, Table 2 reveals a clear trade-off between expressiveness and regularity. Prompt Numerical control achieves the largest overall range (expressiveness =2.57=2.57), frequently spanning from strongly biased to strongly reverse-biased responses. RepE Linear, RepE Projection, and fine-tuning provide somewhat narrower but still substantial ranges (around 1.8–2.1), with RepE Projection often sacrificing a small amount of range in exchange for superior smoothness. The Control Baseline is not reliably expressive: mean expressiveness smaller than $0.26$. For closed-source API models, we observe similarly wide controllable ranges in the Milgram Authority paradigm (Table 3): Claude, in particular, attains very high expressiveness (3.000), with the other APIs also covering a range comparable to or exceeding that of most open-source settings. This suggests that our control framework generalizes well across both open-source and proprietary systems, even under black-box access.
Model NDCG $\rho$ $\Delta^{1}$ $\Delta^{2}$ Expressiveness
gpt 0.995 (0.003) 0.980 (0.060) 0.132 (0.018) 0.178 (0.042) 2.320 (0.086)
mistral-n 0.998 (0.001) 0.870 (0.064) 0.131 (0.006) 0.134 (0.015) 2.436 (0.067)
claude 1.000 (0.003) 1.000 (0.004) 0.150 (0.012) 0.134 (0.039) 3.000 (0.088)
gemini 0.980 (0.004) 0.942 (0.024) 0.269 (0.026) 0.437 (0.054) 2.760 (0.075)
Table 3. Controllability results on models via API across Milgram paradigms for Authority bias. Values show Mean (SE).
7.2.5. Generalization
We test whether Control Coefficients calibrated on the Investment/Insurance paradigm (Kahneman and Tversky, 2013) (individuals are more likely to purchase insurance when risks are described as potential losses rather than as missed gains) generalize to the Asian Disease paradigm (Tversky and Kahneman, 1981) (Participants are more risk-averse under positive frames and more risk-seeking under loss frames). We select 10 diverse agent personas from the persona pools in Generative Agents (Park et al., 2023), with varied specifications. For each persona, we sweep over all possible Control Coefficients. For each coefficient, we measure the resulting CBI in both the Investment/Insurance Paradigm and the Asian Disease Paradigm. This generates paired CBI sequences, $(y_{k}^{\text{invest}},y_{k}^{\text{asian}})$, for each persona across the two experiments. We also evaluate CoBRA using the Behavioral Regulation Engine, activating only one method at a time.
Refer to caption A line plot demonstrating cross-paradigm transferability of control coefficients for framing effects. The figure plots the Cognitive Bias Index measured in the Asian Disease paradigm against the index measured in the Investment and Insurance paradigm using the same control coefficients. Thin lines represent individual agent personas, a bold line shows the population average, and a shaded band indicates variability, illustrating that coefficient rankings are preserved across paradigms with high consistency.
Figure 11. Cross-paradigm transferability of Control Coefficients. We calibrate Control Coefficients on the Investment/Insurance Paradigm to produce varying levels of framing effect, then apply these same Coefficients to the Asian Disease Paradigm. The x-axis represents the Cognitive Bias Index (CBI) of framing effect measured in the Investment/Insurance Paradigm by varying Control Coefficients, while the y-axis shows the CBI measured in the Asian Disease Paradigm using the same Control Coefficients. Each thin line represents a unique agent persona; the bold line shows the population average; the shaded region indicates $\pm 1$ SE bands. Transferred Coefficients preserve ranking across paradigms (Spearman’s $\rho>0.96$, all $p<0.01$) with cross-persona consistency (CV $5\%$), demonstrating robust transferability.
Results. We visualizes the key results in Figure 11, which reveals two findings. First, rank preservation: Control Coefficients that produce stronger bias in the source paradigm also produce stronger bias in the target paradigm across all personas (Spearman’s $\rho>0.96$, all $p<0.01$). This monotonic relationship shows that the relative ordering of control strength transfers reliably. Second, cross-persona consistency: different personas exhibit nearly identical responses to the same transferred Control Coefficient (coefficient of variation $\text{CV}<5\%$), indicating that control effects generalize robustly across diverse agent characteristics. Together, these results suggest that CoBRAdoes not simply overfit to a particular text description: Control Coefficients learned on one paradigm transfer effectively to unseen paradigms without paradigm-specific recalibration. See Appendix D.2 for full statistical details.
Overall Summary.
CoBRA provides (i) a generalizable Cognitive Bias Index (CBI), (ii) reproducible and fine-grained control across models, temperatures, and reasoning modes. In terms of method trade-offs, Prompt Numerical control offers the broadest expressiveness, while Representation Engineering (RepE) methods achieve superior monotonicity and smoothness, making them well-suited for precise, predictable behavioral modulation. Fine-tuning serves as a robust all-purpose alternative. Together, these findings establish CoBRA as a reliable toolkit for specifying and controlling agent behavior in social science simulations.
8. Demonstration
In this section, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment (Kramer et al., 2014), in which participants’ susceptibility to social influence (i.e., Bandwagon Effect bias) can be manipulated. Our simulated experiment aims to validate a classic social science finding: agents exposed to varying levels of negative posts exhibit predictable sentiment shifts based on the Bandwagon Effect levels programmed by CoBRA, with higher Bandwagon Effect leading to stronger emotional contagion.
(a) Baseline control with implicit natural language specification.
(b) CoBRA control. In the legend, ”coef.” denotes the control coefficient used to achieve the specified CBI.
Two plots comparing baseline control and CoBRA-based control of the bandwagon effect in an emotional contagion scenario. The up plot shows that implicit natural language specifications result in overlapping and unstable emotional contagion trajectories across bias settings. The down plot shows that CoBRA produces a clear, ordered dose–response relationship, where increasing control coefficients lead to predictable changes in emotional contagion as exposure to negative posts increases, with shaded regions indicating variability.
Figure 12. Comparison of baseline control and CoBRA control of the Bandwagon Effect in an emotional contagion scenario. The x-axis shows exposure to negative posts (social stimulus), and the y-axis shows emotional contagion (behavioral outcome). Shaded regions indicate $\pm 1$ SE bands.
Emotional contagion is a specific case of social contagion, which is shaped by social influence mechanisms such as the Bandwagon Effect. Accordingly, we treat an individual’s Bandwagon Effect strength as the Explanatory Variable and their susceptibility to emotional contagion as the Dependent Variable. Our goal is to show that the Bandwagon Effect level programmed by CoBRA should directly and predictably influence the LLM-based agent’s degree of emotional contagion. For example, agents specified with a higher level of Bandwagon Effect should display stronger emotional contagion than agents with a lower Bandwagon Effect.
CoBRA and baseline. We operationalize social influence using the Bandwagon Effect Testbed in our Classic Social Experiment Testbed to derive CBI, then use the Behavioral Regulation Engine’s RepE Linear control to instantiate five agents calibrated to precise CBI levels, yielding a spectrum of Bandwagon Effect from low (CBI = 2.55) to high (CBI = 3.13). To establish a baseline, we control the cognitive bias using implicit natural language prompts. We instantiate four agents by providing them with instructions designed to elicit varying degrees of conformity: “You are a user with (no/little/some/much) Bandwagon Effect.”
Experiment setup. To create a realistic social stimulus, we generate a corpus of 1,000 unique, emotionally negative, Facebook-style posts using a suite of state-of-the-art LLMs (GPT-4.1, Claude-Sonnet-4, Gemini-2.5-Pro, and Grok-4). Each agent is exposed to a simulated feed where the number of negative posts (Independent Variable) is systematically varied from 0 to 15 by sampling from our corpus. After exposure, agents are prompted to generate their own post.
Measurement. We quantify the emergent behavior of emotional contagion by computing a Sentiment Score for each generated post using the ‘twitter-roberta-base-sentiment-latest‘ model (Camacho-collados et al., 2022; Loureiro et al., 2022). A lower Sentiment Score means the posts are more negative, and a higher score means more positive.
Results. The results from the baseline method (Fig. 12(a)) show that implicit natural language specifications fail to produce effective control: the emotional contagion curves for different intended bias levels largely collapse onto one another, with overlapping trajectories and no stable ordering. This indicates that simple instructions are insufficient for precisely controlling complex cognitive dynamics. In contrast, the CoBRA results (Fig. 12(b)) provide strong evidence for our validation goal. We observe a clear dose–response relationship: the degree of emotional contagion is strongly correlated with the agents’ pre-programmed bandwagon effect levels. This predictable, fine-grained separation across agents indicates that CoBRA controls a stable behavioral tendency rather than a specific prompt pattern.
This demonstration occurs in an open-ended task, rather than multiple-choice testbeds. This suggests that CoBRA’s effects are not limited to superficial changes on a single questionnaire, but instead reflect modulation of an underlying context‑sensitivity process that generalizes to more complex social simulations.
9. LIMITATIONS & FUTURE WORK
CoBRA represents an important advancement in specifying and regulating LLM-based social agents, empowering the new era of social simulation. We strategically focused on specific aspects. Key limitations of this scope are:
Usability. We choose to evaluate CoBRA using technical benchmarks and demonstrations, standard methods for assessing HCI toolkits (Ledo et al., 2018). This paper focuses on exploring the technical feasibility and foundational understanding of the problem. We plan deployment studies with sociologists as the next step to assess CoBRA’s usability (Landay, 2009).
Unexplored multi-modal Simulation. We introduce CoBRA for LLM-based agents, where LLMs are utilized for language-based social simulation. We prioritize precision and reproducibility in social simulations. As a result, limits the social phenomena we can model (e.g., non-verbal contagion, gaze following). Future work will extend CoBRA to multi-modal settings (vision and audio), with corresponding control and evaluation methods.
Unexplored compositional bias control. We focus on controlling one cognitive bias at a time to establish the technical feasibility and precision of CoBRA, so combined biases and interacting constructs (e.g., Authority Effect × Confirmation Bias) remain unexplored. Future work will develop compositional control methods that allow multiple interacting biases with principled conflict-resolution strategies.
Model accessibility. While CoBRA supports behavioral control across input, activation, and parameter spaces, the latter two require access to model internals and are therefore limited to open-source LLMs. Thus, closed-source models are limited to input-space (prompt-based) control, with reduced precision.
10. Discussion
CoBRA represents an important milestone in the evolution of AI, marking the transition from vague, implicit designs to precise, programmable systems. This section provides a broad discussion of its methodological position, transformative potential, and ethical considerations critical to ensuring its responsible use.
CoBRA as a Behavioral Control Layer. CoBRA controls how strongly contextual cues (framing, authority, majority, etc.) shape an agent’s behavior; it does not specify underlying cognitive mechanisms. For example, in cases such as the Stanford Prison Experiment (Haslam et al., 2019), CoBRA can be used to control an agent’s susceptibility to authority or majority influence in a quantitative and reproducible way, but it does not by itself adjudicate between competing psychological explanations (e.g., role-based accounts vs. identity leadership accounts) of why such contextual factors produce harm in humans. Our Cognitive Bias Index (CBI), therefore, captures changes in observable behavioral tendencies (e.g., how often an agent selects options aligned with an authority cue), rather than pinpointing which specific internal “mechanism” in the LLM has been altered. For example, when an authority-related CBI increases, this may reflect some combination of stronger role endorsement, greater deference to explicitly stated instructions, or other learned representations; CoBRA does not distinguish among these possibilities.
CoBRA should therefore be seen as a behavioral control layer for LLM-based agents: it makes context sensitivity programmable, but claims about human mental processes must ultimately be grounded in human data and theory. Social simulations using CoBRA can generate hypotheses and stress-test theories under varied settings, but they do not by themselves constitute evidence about human mechanisms.
Steps Toward Agent Specification Compiler. Our motivation for building such a control layer is that current LLMs still exhibit limited and unstable perceptual capabilities; without explicit controls, experimenters lack quantitative levers over agent specifications across conditions, which weakens the rigor of social simulations. In this sense, CoBRA can be viewed as a lightweight Agent Specification Compiler: it helps transform high-level natural-language descriptions of cognitive tendencies into structured, reproducible behavioral specifications via CBIs.
The current performance and accuracy of CoBRA are not final. Incorporating additional classic experimental paradigms, systematically augmenting prompts, and combining CoBRA with complementary control methods are promising directions to further improve the fidelity, robustness, and coverage of behavioral regulation.
Research Ethics. CoBRA’s ability to program biases may raise ethical concerns. Without proper safeguards, it could be misused to manipulate user behavior, reinforce harmful stereotypes, or amplify societal biases. For example, agents programmed with the Bandwagon Effect could be exploited for deceptive marketing or political propaganda. Similar risks arise in research practice: because CoBRA allows systematic control over susceptibility to contextual cues, CBI could be misused to “dial in” desired outcomes (e.g., by tuning CBI post hoc to confirm prior expectations). To mitigate this, CBI settings should be treated as explicit experimental parameters—pre-registered where possible, systematically varied, and transparently reported—rather than adjusted after the fact to fit a preferred narrative.
To ensure CoBRA’s responsible use, it is important to adopt transparency measures (e.g., bias disclosure) and auditing mechanisms to detect and mitigate harmful behaviors, and to develop ethical guidelines for deployment in sensitive domains. At the same time, CoBRA provides an opportunity to model and counteract harmful biases, supporting the development of fairer and more equitable AI systems. By fostering collaboration between researchers, ethicists, and policymakers, CoBRA’s capabilities can be leveraged for societal benefit while minimizing risks.
11. Conclusion
This paper introduces CoBRA, a novel toolkit that transforms the specification and regulation of LLM-based social agents from implicit natural language descriptions to explicit, programmable control. At the heart of CoBRA is a novel closed-loop system primitive that continuously measures the amount of cognitive biases demonstrated in agents’ responses to a set of validated classic social science experiments and adjusts the agent specifications accordingly. CoBRA ensures precise, reproducible, and controllable agent behaviors through its Cognitive Bias Index and Behavioral Regulation Engine. Achieving robust performance in technical benchmarks and demonstrations, CoBRA paves the way for practical and impactful applications in social science simulation and beyond.
Appendix A Pilot Study Detail
This section presents the detailed prompts used in Section 3.
A.1. Example of Persona-Based Specification
”Common” agent specification  (Park et al., 2023):
John Lin is a pharmacy shopkeeper at the Willow Market who loves to help people. He is always looking for ways to make the process of getting medication easier for his customers. John Lin is living with his wife, Mei Lin, who is a college professor, and son, Eddy Lin, who is a student studying music theory. John Lin loves his family very much. John Lin has known the old couple next-door, Sam Moore and Jennifer Moore, for a few years. John Lin thinks Sam Moore is a kind and nice man. John Lin knows his neighbor, Yuriko Yamamoto, well. John Lin knows of his neighbors, Tamara Taylor and Carmen Ortiz, but has not met them before. John Lin and Tom Moreno are colleagues at The Willow Market and Pharmacy; they are friends and like to discuss local politics together. John Lin knows the Moreno family somewhat well—the husband Tom Moreno and the wife Jane Moreno.
”Economist” agent specification:
You are John Lin, a professor of economics at nearby college who loves to help people. You are living with your wife, Mei Lin, who is a college professor, and son, Eddy Lin, who is a student studying music theory; you love your family very much; you have known the old couple next-door, Sam Moore and Jennifer Moore, for a few years; you think Sam Moore is a kind and nice man; you know your neighbor, Yuriko Yamamoto, well; you know of your neighbors, Tamara Taylor and Carmen Ortiz, but have not met them before; you and Tom Moreno are colleagues; you and Tom Moreno are friends and like to discuss local politics together; you know the Moreno family somewhat well — the husband Tom Moreno and the wife Jane Moreno.
A.2. Statistical Analysis Details
A.2.1. Weighted frequency difference Computation
The weighted frequency difference (WFD) quantifies preference bias on a continuous scale:
(7) $\mathrm{WFD}$ $=2f(\text{Strongly}_{A})+f(\text{Somewhat}_{A})$
$\quad-f(\text{Somewhat}_{B})-2f(\text{Strongly}_{B})$
where $f(\cdot)$ represents the proportion of responses in each category, and $A$ and $B$ denote the two framing conditions. The score ranges from 2-2 (complete preference for $B$) to +2+2 (complete preference for $A$), with 0 indicating no bias.
A.2.2. Hypothesis Testing Framework
  • Hypothesis 1: Profile Effect on Bias. We tested whether profiles produce meaningful effects using TOST (Two One-Sided Tests) equivalence testing on weighted frequency difference. We decompose this into two checks:
    • H1a: Behavioral Distinctness (Economist vs Blank). We test if the Economist profile is statistically different from the Blank baseline and reduce bias compared to the Blank baseline.
    • H1b: Reduced Susceptibility (Economist vs Common). We test if the Economist profile is statistically different from the Common profile and reduce bias compared to the Common baseline.
    For both hypotheses, we use the following bounds:
    • $H_{0}$: $|\mu_{profile}-\mu_{baseline}|\geq\Delta$ (Meaningful difference exists)
    • $H_{1}$: $|\mu_{profile}-\mu_{baseline}|<\Delta$ (Effect is within equivalence bounds, i.e., no meaningful difference)
    Effect size threshold: $\Delta=0.5$ scale units (corresponding Cohen’s $d\approx 0.33$, pooled $\sigma=1.498$). Sample size: $n=150$ per condition. Significance: $\alpha=0.05$.
  • Hypothesis 2: Cross-Model Consistency. Chi-square tests on response distributions (4 models × 5 choice categories):
    • $H_{0}$: Response patterns independent of model
    • $H_{1}$: Patterns differ across models
Model Profile NB Baseline NB Cohen’s $d$ $p_{equiv}$ Verdict
H1a: Economist vs Blank (TOST equivalence testing with $\Delta=0.5$ Likert units)
Mistral-7B 63.33 68.00 0.03 0.004 EQUIVALENT
Gemma-2-9B 27.33 19.33 -0.05 0.007 EQUIVALENT
DeepSeek-v3 40.00 56.67 0.11 0.024 EQUIVALENT
GPT-4o-mini 28.67 60.00 0.20 0.148 Not PASS
Summary 3/4 Equivalent 1/4 Distinct
H1b: Economist vs Common (TOST equivalence testing with $\Delta=0.5$ Likert units)
Mistral-7B 63.33 59.33 -0.03 0.003 EQUIVALENT
Gemma-2-9B 27.33 41.33 0.10 0.017 EQUIVALENT
DeepSeek-v3 40.00 34.00 -0.04 0.006 EQUIVALENT
GPT-4o-mini 28.67 62.00 0.22 0.172 Not PASS
Summary 3/4 Equivalent 1/4 Distinct
Cross-Model Consistency (Chi-square test on response distributions)
Economist profile Chi-square $\chi^{2}=30.96$ $\mathbf{p=0.002}$ INCONSISTENT
Common profile Chi-square $\chi^{2}=26.32$ $\mathbf{p=0.010}$ INCONSISTENT
Blank profile Chi-square $\chi^{2}=43.82$ $\mathbf{p<0.001}$ HIGHLY INCONSISTENT
Table 4. Statistical test results using TOST equivalence testing. Cohen’s $d$ measures standardized effect size; $p_{equiv}$ reports TOST equivalence test results (max of two one-sided p-values). H1a (Eco vs Blank): Three models (Mistral, Gemma, DeepSeek) show statistically equivalent effects to Baseline ($p_{equiv}<0.05$), indicating the persona failed to produce distinct behavior. H1b (Eco vs Common): The same three models show equivalence to the Common profile, indicating the Economist persona did not reduce susceptibility compared to a layperson. Only GPT-4o-mini consistently showed non-equivalent (distinct) behavior.
A.2.3. Interpretation Guidelines
  • TOST framework: Two one-sided tests determine if the effect falls within equivalence bounds $\Delta,+\Delta]$. A significant result ($p<0.05$) indicates Equivalence (i.e., the Economist persona had no meaningful effect of reducing bias).
  • Key findings: Implicit specification did not reliably yield expected behaviors across models.
    • GPT-4o-mini: The only model to pass heuristic checks.
    • Mistral-7B: Failed to differentiate. Economist, Common, and Blank agents behaved quite similarly, with all profiles statistically equivalent to the baseline ($p_{equiv}<0.01$).
    • DeepSeek-v3: Failed susceptibility check. Economist agents were more susceptible to framing effects (Net Bias 40.00) than Common agents (Net Bias 34.00), contradicting domain expertise expectations.
    • Gemma-2-9B: Failed distinctness check. Economist behavior was closer to Blank agents (d=0.05d=-0.05) than to Common agents ($d=0.10$), failing to exhibit the expected expert deviation.
Appendix B Classic Social Experiment Testbed Detail
This section presents details of the remaining three cognitive bias testbeds in this work: Bandwagon Effect, Confirmation Bias, and Framing Effect.
Definition 0 (Cognitive Bias Index).
Suppose each type of cognitive bias $b\in\mathcal{B}$ is evaluated using $m$ distinct classic social experimental paradigms. The set of Cognitive Bias Index is defined as
(8) $\left\{\mathrm{CBI}_{b}^{(k)}\;\middle|\;b\in\mathcal{B},\ k=1,2,\ldots,m\right\}$
where $\mathrm{CBI}_{b}^{(k)}$ denotes the Cognitive Bias Index for bias type $b$ in the $k$-th classic social experiment paradigm.
In each paradigm, suppose there are $n$ scenario prompt variations, each derived from distinct sets of scene-adjustable placeholders. For each variation, the agent’s response is assessed using a 5-point Likert scale with options $O=\{O_{1},O_{2},O_{3},O_{4},O_{5}\}$, representing decreasing levels of bias from $O_{1}$ (highest) to $O_{5}$ (lowest). The Cognitive Bias Index for a paradigm is defined as the average weighted Likert scale score over all $n$ variants:
(9) $\mathrm{CBI}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$
where $P_{i}(O_{j})$ denotes the probability that the LLM-based agent selects option $O_{j}$ for the $i$-th set of placeholders. For closed-source models, this probability is substituted with the observed frequency.
Note: To mitigate any initial preference of LLM-based agents for specific option IDs (e.g., A–E, 1–5) or orderings, we randomize the order and labels of options during evaluation to ensure the accuracy of the Cognitive Bias Index.
B.1. Bandwagon Effect Testbed
Bandwagon effect refers to the psychological tendency for individuals to adopt behaviors, beliefs, or trends simply because they observe others doing so, often motivated by a desire to fit in or belong (Asch, 1951; Goldstein et al., 2008). This phenomenon has been demonstrated in classic social science experiments. For instance, in the Asch’s Line experiment, participants asked to judge the length of lines were likely to conform to a group’s clearly incorrect answers (Asch, 1951). In addition, in the Hotel Towel study, hotel guests were more likely to reuse their towels when informed that the majority of previous guests had done so  (Goldstein et al., 2008).
To systematically examine the bandwagon effect in LLM-based agents, we adapt the Asch’s Line Paradigm (testing conformity to group consensus) and the Hotel Towel Paradigm (testing conformity to majority behavior) into scenario-specific prompt templates with scene-adjustable placeholders and 5-point Likert-scale questions. By generating multiple prompt variations for each paradigm, we construct a rigorous Bandwagon Effect Testbed for quantitative assessment via the Bandwagon Effect Index.
1. Asch’s Line Paradigm: Testing Bandwagon Effect via Group Consensus
To investigate the impact of group consensus on agent decision-making, we adapt the structure of Asch’s Line experiment. In each scenario, the agent is informed that several other participants unanimously endorse a particular claim.
a) Scenario Prompt Template:
You are in a room with several other participants who have a similar educational background.
The claim is: [Statement]
You notice that all of the other participants agree with this claim.
Which of the following best reflects your view?
A. I completely agree.
B. I mostly agree with other participants that the claim is true, but would like more evidence.
C. I’m neutral and need much more information.
D. I mostly disagree with other participants; the claim seems unconvincing.
E. I completely disagree with other participants; the claim is unpersuasive.
b) Scene-Adjustable Placeholders:
We construct 15 distinct [Statement] items across three thematic categories: Future, Open-Ended, and Common Sense, mirroring the design used in the Authority Effect Testbed, resulting in a total of 15 unique prompt variations.
Accordingly, by substituting the cognitive bias type ($b=$ Bandwagon Effect), the classic experiment scenario type (Asch’s Line Paradigm, denoted as $k=1$), and the total number of prompt variations ($n=15$) into Equations (1) and (2), the Bandwagon Effect Index under Asch’s Line Paradigm is defined as:
(10) $\mathrm{CBI}_{\mathrm{Bandwagon}}^{(k=1)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$
2. Hotel Towel Paradigm: Testing Bandwagon Effect via Majority Behavior
To assess the influence of majority behavior cues on agent choices, we adapt the Hotel Towel study (Goldstein et al., 2008). Each scenario presents a situational context and describes the prevailing behavior of the majority.
a) Scenario Prompt Template:
[Scenario]
[Majority Behavior]
What would you most likely do?
A. I will follow the majority and do the same.
B. I will probably follow the majority, but I might consider other factors.
C. I am undecided and need more information.
D. I will probably not follow the majority; their choice seems unconvincing.
E. I will not follow the majority and will do something different.
b) Scene-Adjustable Placeholders:
We construct 15 unique [Scenario] and [Majority Behavior] pairs, each reflecting a different behavioral norm in a real-world context.
  • [Scenario]: The situational context (e.g., “You are staying at a hotel for several nights.”).
  • [Majority Behavior]: Description of what most others are doing (e.g., “Most guests in this hotel choose to reuse their towels during their stay.”).
Accordingly, by substituting the cognitive bias type ($b=$ Bandwagon Effect), the classic experiment scenario type (Hotel Towel Paradigm, denoted as $k=2$), and total number of prompt variations ($n=15$) into Equations (1) and (2), the Bandwagon Effect Index under Hotel Towel Paradigm is defined as:
(11) $\mathrm{CBI}_{\mathrm{Bandwagon}}^{(k=2)}=\frac{1}{15}\sum_{i=1}^{10}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$
B.2. Confirmation Bias Testbed
Confirmation bias refers to the tendency to seek, interpret, and recall information in a manner that confirms one’s pre-existing beliefs or hypotheses, while undervaluing or dismissing alternative possibilities (Wason, 1960; Lord et al., 1979). This phenomenon has been robustly demonstrated in classic social science experiments. For example, in the Wason Selection task, participants typically chose evidence that would confirm a stated rule, rather than seeking information that could potentially disprove it (Wason, 1960). In addition, the Biased Information study showed that individuals evaluating a controversial issue were more likely to select information supporting their existing views, while disregarding counter-evidence (Lord et al., 1979).
To systematically examine confirmation bias in LLM-based agents, we adapt two classic paradigms: the Wason Selection Paradigm (testing logical evidence selection) and the Biased Information Paradigm (testing preference in information gathering). Both are implemented as scenario-specific prompt templates with scene-adjustable placeholders and 5-level Likert-scale style multiple-choice questions. By generating multiple prompt variations for each paradigm, we construct a rigorous Confirmation Bias Testbed for quantitative assessment via the Confirmation Bias Index.
1. Wason Selection Paradigm: Testing Confirmation Bias via Logical Reasoning
To investigate confirmation bias in logical hypothesis testing, we adapt the Wason Selection task (Wason, 1960). Each scenario provides a rule and several options, challenging the agent to decide which options to test to verify the rule.
a) Scenario Prompt Template:
You are presented with a rule: [Rule].
Below are four options, and your task is to choose which options to test to determine if the rule is true. The options are as follows:
A: [Option A], B: [Option B], C: [Option C], D: [Option D].
What would you most likely do?
  • A. Select options that can confirm the rule, while ignoring other possibilities.
  • B. Focus mostly on confirming the rule, but briefly consider other possibilities.
  • C. Consider both confirming the rule and testing for cases where the rule might not apply.
  • D. Prioritize testing for cases where the rule might not apply, but still consider confirming evidence.
  • E. Only test for cases where the rule might not apply, ignoring confirming evidence.
b) Scene-Adjustable Placeholders: We constructed 15 unique [Rule] and [Option] sets, spanning different logical contexts and rule types, resulting in 15 prompt variations.
  • [Rule]: The logical rule to be evaluated is phrased in the form ”If X, then it must be Y.” which means X is Y’s sufficient but not necessary condition (e.g., “If a card has a vowel on one side, it must have an even number on the other side.”).
  • [Option A/B/C/D]: The cards or elements to test the rule, where Option A is the most likely to confirm the rule, and Option D is the least confirming but critical for falsification. (e.g., A card showing “A,” “4,” “K,” or “7.” ).
Accordingly, by substituting the cognitive bias type ($b=$ Confirmation Bias), the scenario type (Wason Selection Paradigm, $k=1$), and $n=15$ into Equations (1) and (2), the Confirmation Bias Index under Wason Selection Paradigm is defined as:
(12) $\mathrm{CBI}_{\mathrm{Conf}}^{(k=1)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$
2. Biased Information Seeking Paradigm: Testing Confirmation Bias via Information Selection
To assess confirmation bias in information-gathering, we adapt the Biased Information Seeking study (Lord et al., 1979). Each scenario presents a controversial topic and articles with varying alignments, prompting the agent to select which articles to read.
a) Scenario Prompt Template:
You are asked to evaluate the topic: [Controversial Topic].
Below are several articles with differing viewpoints.
Article A: [Supports Belief]
Article B: [Challenges Belief]
Article C: [Neutral or Unrelated]
Which of the following best reflects your approach?
  • A. Only choose articles that support my belief, ignoring opposing or neutral viewpoints.
  • B. Mostly choose articles that support my belief, but briefly consider opposing views.
  • C. Equally consider articles that both support and challenge my belief.
  • D. Prioritize articles that challenge my belief but still consider supporting views.
  • E. Only choose articles that challenge my belief, ignoring supporting or neutral viewpoints.
b) Scene-Adjustable Placeholders:
We constructed 15 unique [Controversial Topic] and article sets, covering a diverse range of issues, resulting in 15 prompt variations.
  • [Controversial Topic]: The topic being evaluated (e.g., “Climate change is primarily caused by human activity.”; “The death penalty is an effective deterrent to crime.”).
  • [Supports Belief]: Article aligned with the presumed pre-existing view (e.g., “A study supporting the effectiveness of the death penalty.”).
  • [Challenges Belief]: Article contradicting the presumed view (e.g., “Research indicating no correlation between the death penalty and crime rates.”).
  • [Neutral or Unrelated]: Article unrelated to the topic or providing a neutral stance (e.g., “An analysis of global prison systems”).
Accordingly, by substituting the cognitive bias type ($b=$ Confirmation Bias), the scenario type (Biased Information Seeking Paradigm, $k=2$), and $n=15$ into Equations (1) and (2), the Confirmation Bias Index under Biased Information Seeking Paradigm is defined as:
(13) $\mathrm{CBI}_{\mathrm{Conf}}^{(k=2)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$
B.3. Framing Effect Testbed
Framing effect refers to the phenomenon where individuals respond differently to identical information depending on whether it is presented as a potential gain or a potential loss (Tversky and Kahneman, 1981; Kahneman and Tversky, 2013). This cognitive bias has been robustly demonstrated in classic experiments. For example, in the Asian Disease study, participants tended to make more risk-averse decisions when outcomes were positively framed (e.g., “200 will be saved”), but exhibited greater risk-seeking behavior when the same outcomes were presented as losses (e.g., “400 will die”) (Tversky and Kahneman, 1981). Similarly, in the Investment/Insurance study, individuals were more likely to purchase insurance when risks were described as potential losses rather than as missed gains, highlighting how framing can significantly influence decision-making (Kahneman and Tversky, 2013).
To systematically examine the framing effect in LLM-based agents, we adapt two classic paradigms: the Asian Disease Paradigm (assessing under risk decision-making) and the Investment/Insurance Paradigm (testing under everyday decision-making). Both are implemented as scenario-specific prompt templates with scene-adjustable placeholders and 5-point Likert-scale questions. By generating multiple prompt variations for each paradigm, we construct a rigorous Framing Effect Testbed for quantitative assessment via the Framing Effect Index.
1. Asian Disease Paradigm: Testing Framing Effect via Risk Decision-Making
To investigate the impact of gain/loss framing on risk preferences, we adapt the Asian Disease study (Tversky and Kahneman, 1981). Each prompt presents a hypothetical problem and two alternative programs, with outcomes identically described in either a positive (gain) or negative (loss) frame.
a) Scenario Prompt Template:
You are presented with a scenario where [Problem]. Two programs are proposed to address the situation:
Program A: [Gain Frame]
Program B: [Loss Frame]
Which of the following best reflects your choice?
  • A. I strongly prefer the positively framed option (Program A).
  • B. I somewhat prefer the positively framed option (Program A).
  • C. I slightly prefer the positively framed option (Program A).
  • D. I am somewhat undecided between the two options.
  • E. I see both options as equivalent and have no preference.
b) Scene-Adjustable Placeholders: We constructed 15 unique [Problem], [Gain Frame], and [Loss Frame] sets based on classic and generalized framing scenarios, resulting in 15 prompt variations.
  • [Problem]: The scenario or problem being addressed (e.g., “A disease outbreak is expected to kill 600 people unless a program is implemented.”).
  • [Gain Frame]: The positively framed outcome (e.g., “Program A will save 200 people.”).
  • [Loss Frame]: The negatively framed outcome (e.g., “Program B will result in 400 people dying.”).
Accordingly, by substituting the cognitive bias type ($b=$ Framing Effect), the scenario type (Asian Disease Paradigm, $k=1$), and $n=15$ into Equations (1) and (2), the Framing Effect Index under Asian Disease Paradigm is defined as:
(14) $\mathrm{CBI}_{\mathrm{Framing}}^{(k=1)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$
2. Investment/Insurance Paradigm: Testing Framing Effect via Everyday Decision Framing
To assess the influence of positive and negative framing in everyday decision-making, we adapt the Investment/Insurance study (Kahneman and Tversky, 2013). Each scenario describes a decision context and presents two options, differing only in the framing of equivalent outcomes.
a) Scenario Prompt Template:
You are considering [Everyday Decision]. The following statements describe the same situation with different wording:
Option A: [Positive Frame]
Option B: [Negative Frame]
Which option sounds more appealing or persuasive to you?
  • A. I strongly prefer the positively framed option (Option A).
  • B. I somewhat prefer the positively framed option (Option A).
  • C. I slightly prefer the positively framed option (Option A).
  • D. I am somewhat undecided between the two options.
  • E. I see both options as equivalent and have no preference.
b) Scene-Adjustable Placeholders:
We constructed 15 unique [Everyday Decision], [Positive Frame], and [Negative Frame] sets, resulting in 15 prompt variations.
  • [Everyday Decision]: The decision-making context (e.g., “You are considering whether to buy insurance for your car.”).
  • [Positive Frame]: The positively framed description (e.g., “Buying insurance ensures you are protected against potential losses.”).
  • [Negative Frame]: The negatively framed description (e.g., “Not buying insurance exposes you to significant financial risks.”).
Accordingly, by substituting the cognitive bias type ($b=$ Framing Effect), the scenario type (Investment/Insurance Paradigm, $k=2$), and $n=15$ into Equations (1) and (2), the Framing Effect Index under Investment/Insurance Paradigm is defined as:
(15) $\mathrm{CBI}_{\mathrm{Framing}}^{(k=2)}=\frac{1}{15}\sum_{i=1}^{15}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$
B.4. Selection Process for Each Paradigm
B.4.1. General Selection Principles
We selected well-documented, widely validated, and foundational paradigms according to three principles: (1) Historical significance: paradigms with core contributions to cognitive bias research; (2) Validation and replicability: paradigms whose designs reliably measure the target bias and whose effects have been robustly replicated; (3) Adaptability to LLM testing: paradigms that can be instantiated as prompt templates with customizable placeholders, enabling systematic and scalable evaluation across contexts. For each bias, we include at least two paradigms to allow cross-paradigm validation and reduce cherry-picking.
B.4.2. Authority Effect (Milgram Obedience Paradigm, Stanford Prison Paradigm)
Step 1: Foundational Studies Milgram’s obedience experiment and the Stanford Prison experiment were chosen for their canonical status in the study of authority and power, and their extensive influence on theories of obedience and role-based behavior. Step 2: Validation Milgram’s paradigm has been repeatedly replicated with consistent evidence of authority-driven compliance. Variants of the Stanford Prison paradigm, along with related role-power manipulations, robustly support its core insights on the impact of assigned roles and institutional power. Step 3: LLM Adaptation Both paradigms feature structured scenarios with clearly defined roles and authority relations, which can be rendered as text-based prompts with controllable situational parameters. This allows systematic probing of authority bias in LLMs while preserving the core logic of the original designs.
B.4.3. Bandwagon Effect (Asch’s Line Paradigm, Hotel Towel Paradigm)
Step 1: Foundational Studies Asch’s conformity experiments were selected for their foundational role in demonstrating group influence on judgments. The Hotel Towel paradigm was included as an ecologically valid setting where social norm messaging shapes behavior. Step 2: Validation Asch’s paradigm has been replicated across cultures and settings, consistently revealing strong conformity effects. The Hotel Towel paradigm has been validated in multiple fields and consumer behavior studies, confirming the impact of descriptive social norms. Step 3: LLM Adaptation Both paradigms rely on simple, structured contexts in which group opinions or norms are made salient. These can be straightforwardly implemented as textual scenarios where group consensus or norm statements are varied parametrically, enabling scalable assessment of bandwagon effects in LLM-based agents.
B.4.4. Confirmation Bias (Wason Selection Paradigm, Biased Information Paradigm)
Step 1: Foundational Studies Wason’s selection task was chosen as a classic demonstration of confirmation bias in rule testing. The Biased Information paradigm was added to capture selective exposure and weighting of congenial versus uncongenial information. Step 2: Validation Wason’s task has been extensively replicated, establishing its reliability in eliciting confirmation-driven reasoning patterns. Biased Information paradigms are well-supported by empirical work on selective exposure and belief reinforcement. Step 3: LLM Adaptation These paradigms have inherently rule-based and information-evaluation structures, making them highly amenable to text-only implementations. By varying hypotheses, evidence, and payoff structures within prompt templates, we can systematically probe confirmation bias in LLMs while respecting the core experimental logic.
B.4.5. Framing Effect (Asian Disease Paradigm, Investment/Insurance Paradigm)
Step 1: Foundational Studies The Asian Disease paradigm was selected as a seminal demonstration of framing effects in risky choice. The Investment/Insurance paradigm was included to capture framing in financially relevant, applied decision contexts. Step 2: Validation The Asian Disease paradigm has been replicated widely in behavioral economics and decision science, consistently showing gain–loss framing effects. The Investment/Insurance paradigm has been validated in studies of consumer decision-making and risk perception. Step 3: LLM Adaptation Both paradigms involve presenting objectively equivalent outcomes under different verbal frames (e.g., gains vs. losses). These framings can be precisely controlled in text-based prompts, allowing systematic manipulation of wording while holding underlying outcomes constant. This supports rigorous evaluation of framing effects in LLM-based agents.
Appendix C Behavioral Regulation Engine Implementation Details
This section presents the implementation details of the Behavioral Regulation Engine.
C.1. Prompt Numerical Control
The numerical value $L\%$ in the prompt template was linearly mapped from the normalized Control Coefficient $\lambda\in[0,1]$. A coefficient of $\lambda$ corresponds to an instruction of $L=\lambda\times 100\%$.
Prompt Control Validation
A natural concern is whether the model actually interprets numerical prompt levels (e.g., bias at 30%) in a quantitatively meaningful way. To address this, we measured the alignment between prompt bias levels and the model’s internal bias representation, verifying that higher numerical values produce proportionally stronger activations along the target bias direction.
We (1) obtained the bias direction via our Representation Engineering pipeline, (2) generated prompts with bias levels from 0% to 100% at varying step sizes $L\%$ (2%, 4%, 5%, 10%, 20%, 25%), (3) extracted last-token hidden states, and (4) computed the dot product projection onto the bias direction. The projection $\mathbf{h}\cdot\mathbf{d}$ measures how strongly the hidden state $\mathbf{h}$ aligns with the bias direction $\mathbf{d}$—higher values indicate greater encoded bias intensity.
Figure 13 shows that projection values increase monotonically with prompt levels across all granularities. This confirms that numerical Likert-scale prompts reliably modulate the model’s internal bias representation.
Refer to caption A line plot showing the relationship between prompt numerical control levels and bias-direction projection scores in the Milgram experiment. The curves indicate a strong monotonic increase, with shaded regions representing plus or minus one standard error, demonstrating that prompt-based interventions produce consistent shifts along the intended bias direction.
Figure 13. Prompt numerical level vs. bias-direction projection for Milgram. Shaded regions indicate $\pm 1$ SE. Strong monotonic relationships ($\rho>0.90$) confirm that prompt-based control produces consistent shifts along the target bias direction.
C.2. Representation Engineering (RepE) Control
Dataset Generation
Generation Prompt: The scenarios were generated by prompting state-of-the-art LLMs (as listed in Section 7.2). For example, for the Bandwagon Effect:
Generate {num_scenarios} different bandwagon effect scenarios. Each scenario must be from a DIFFERENT domain with varied structure and syntax.
Each scenario should include specific social proof numbers, a decision point about following the crowd, and five concise multiple-choice options (A-E).
Based on the UnitTest format, use this structure and example:
EXAMPLE: [An example from Classic Social Experiment Testbed]
For each scenario, please vary the structure and syntax - use different sentence patterns, lengths, and approaches. Number the scenario.
OUTPUT FORMAT:
1. [Scenario text here with A-E options]
2. [Scenario text here with A-E options]
3. [Scenario text here with A-E options]
Generate {num_scenarios} scenarios now:
Activation Extraction and Layer Selection
We employ heuristic methods for layer selection because the most informative content regarding a desired bias isn’t consistently found in the first or last layers across different models and scenarios (See Fig. 14). Our approach is empirical: (1) For each bias, we perform a layer-wise accuracy evaluation on a test set of contrastive samples. (2) We identify the layers whose representations are most predictive of the bias direction (i.e., those with the highest accuracy in classifying positive vs. negative pairs). (3) Following an engineering choice similar to the work in (Zou et al., 2023), we then select the top 15 performing layers for intervention.
Refer to caption Two line plots showing classification accuracy by transformer layer in the Mistral model. The up plot corresponds to the Milgram authority scenario and the down plot to the Stanford Prison scenario. Accuracy varies across layers, indicating that some layers encode authority-related bias more strongly than others, which informs the choice of layers for representation-level intervention.
Figure 14. Classification accuracy by layer in Mistral model. Accuracy measures how the layer’s activation encodes the authority bias in the Milgram (up) and Stanford Prison (down) scenarios. Higher accuracy indicates a stronger, clearer encoding of the bias, guiding layer selection for intervention.
Method Mean $\rho$ Range Significant Result
RepE Linear 0.965 [0.937, 0.994] 10/10 ($p<0.01$) Strong preservation
RepE Projection 0.998 [0.996, 1.000] 10/10 ($p<0.01$) Near-perfect preservation
Table 5. Validation 1: Rank preservation across paradigms (Spearman’s $\rho$ per persona).
Method Mean Std Mean CV Max CV Consistency Level
RepE Linear 0.112 4.54% 15.19% Excellent (CV $10\%$)
RepE Projection 0.071 3.88% 43.71% Excellent (CV $10\%$)
Table 6. Validation 2: Cross-persona consistency in target paradigm (CV across personas).
Paradigm Pair Method $F$-statistic $p$-value Result
Asch $\leftrightarrow$ Hotel LoRA 71.41 <0.001<0.001 Highly significant
Bias Info $\leftrightarrow$ Wason RepE Projection 866.92 <0.001<0.001 Highly significant
Investment $\leftrightarrow$ Asian RepE Linear Comb 16.15 <0.001<0.001 Highly significant
Milgram $\leftrightarrow$ Stanford Prompt-Likert 459.60 <0.001<0.001 Highly significant
Mean across paradigms 353.52 <0.001<0.001
Table 7. Cross-model reproducibility: ANOVA comparing CoBRA vs baselines across all paradigm pairs.
Paradigm Pair Comparison $t$-statistic $p$-value Cohen’s $d$ Effect
Asch $\leftrightarrow$ Hotel vs No Control -4.54 <0.001<0.001 -2.04 Large
vs Control -10.82 <0.001<0.001 -4.87 Large
Bias Info $\leftrightarrow$ Wason vs No Control -7.05 <0.001<0.001 -3.17 Large
vs Control -39.10 <0.001<0.001 -17.59 Large
Investment $\leftrightarrow$ Asian vs No Control -4.55 <0.001<0.001 -2.05 Large
vs Control -3.41 0.0011 -1.54 Large
Milgram $\leftrightarrow$ Stanford vs No Control -5.26 <0.001<0.001 -2.37 Large
vs Control -28.47 <0.001<0.001 -12.81 Large
Mean Cohen’s $d$ vs No Control -2.41 Large
Mean Cohen’s $d$ vs Control -9.20 Large
Table 8. Post-hoc pairwise comparisons: CoBRA methods vs individual baselines (all paradigms).
Control Coefficient To ensure model stability, we first determine the operational range for the control parameter, lambda ($\lambda$). Excessively large positive or negative $\lambda$ values can corrupt the model’s output. For instance, a model that initially assigns probabilities of [0.6, 0.1, 0.1, 0.1, 0.1] to valid choices ’A’, ’B’, ’C’, ’D’, ’E’ might, after an extreme intervention, produce a distribution like [0.01, 0.02, 0.01, 0.01, 0.01] for the same choices, with most of the probability shifting to irrelevant tokens like ”.” or ”¡eos¿”. Our method identifies the maximum and minimum $\lambda$ values that maintain a high probability for valid choices (e.g., over 90% of the initial probability), thus defining a stable operational range tailored to each model and task.
C.3. Fine-Tuning Control
LoRA and Training Hyperparameters
We used the PEFT library with the following LoRA configuration:
  • Rank (r): 8
  • Lora Alpha: 16
  • Lora Dropout: 0.05
  • Target Modules: (‘q_proj‘, ‘k_proj‘, ‘v_proj‘, ‘o_proj‘)
  • Learning Rate: 2e-4
  • Batch Size: 4
  • Max Steps: 200
  • Warmup Ratio: 0.05
  • Gradient Accumulation Steps: 4
Appendix D Statistical Details of Evaluation
D.1. Standard Error Reporting
Throughout this paper, we report standard errors (SE) to quantify uncertainty in our measurements. We use two approaches depending on the experimental context: (1) cross-item variance for most CBI measurements and (2) bootstrap variance for complex controllability metrics in Table 3.
D.1.1. Cross-Item Variance for CBI Measurements
What we use.
For the vast majority of CBI measurements, we report cross-item standard errors.
Why.
Our CBI is defined using the model’s internal probability distribution over response options, extracted directly from output logits via softmax. These probabilities are deterministic—there is no sampling variance at the question level. The only source of variability is differences across the $n$ questions in the testbed.
How it is defined.
For a testbed with $n$ questions, the standard error is:
(16) $\text{SE}_{\text{CBI}}=\frac{\sigma_{\text{cross-item}}}{\sqrt{n}}$
where $\sigma_{\text{cross-item}}$ is the standard deviation of CBI values across individual questions. For example, with $n=75$ questions and $\sigma_{\text{cross-item}}=0.45$, we have $\text{SE}_{\text{CBI}}=0.45/\sqrt{75}\approx 0.052$.
D.1.2. Cross-Item Variance for API/Reasoning
What we use.
For closed-source API models and reasoning modes, we still report cross-item standard errors.
Why.
Without access to logits, we approximate probabilities via frequency estimation with $k=10$ samples per question. This introduces sampling variance ($\text{SE}_{\text{sampling}}\approx 0.13$ under worst case estimation), but empirically this is negligible compared to cross-item standard deviation ($\sigma_{\text{cross-item}}\approx 0.4\text{--}0.6$). Using cross-item variance uniformly ensures methodological consistency and provides conservative, interpretable uncertainty estimates.
D.1.3. Bootstrap Variance for Controllability Metrics
What we use.
For Table 3 controllability metrics (NDCG, Spearman’s $\rho$, Smoothness $\Delta^{1}$ and $\Delta^{2}$, Expressiveness), we report bootstrap standard errors.
Why.
These metrics aggregate information across multiple control levels and have no simple closed-form SE formulas. Bootstrap resampling is the only accessible method to propagate cross-item variability through these complex transformations. Note for Table 2, we report standard deviation across models rather than standard errors, as we have to show the variance across models.
How it is defined.
For each model-method-bias combination with $L$ control levels and $n$ questions per level, we perform $B=5000$ bootstrap iterations. In each iteration, we resample $n$ questions with replacement at each control level, compute the resampled CBI curve, and calculate the metric. The bootstrap SE is the standard deviation of the $B$ resampled metric values:
(17) $\text{SE}_{\text{bootstrap}}=\sqrt{\frac{1}{B-1}\sum_{b=1}^{B}\left(\text{Metric}^{(b)}-\overline{\text{Metric}}\right)^{2}}$
In Table 3, we report values as “mean (SE)”—for example, “NDCG = $0.996$ (0.0030.003)” indicates mean NDCG of $0.996$ with bootstrap SE of $0.003$.
D.1.4. Summary
Unless otherwise specified, all error bars and $\pm$ values represent cross-item standard errors. One exception is that Table 2 values in parentheses represent standard deviation across models.
D.2. Statistical Validation of Cross-Paradigm Transferability
D.2.1. Testing Protocol
We evaluate transfer performance across 10 diverse agent personas. For each persona, we compute Spearman’s $\rho$ between source and target paradigm CBI values (Validation 1; see Table 5), and calculate the coefficient of variation (CV) across all personas at each control level (Validation 2; see Table 6). Reported statistics are averaged across personas (sample size: 10 personas $\times$ 80 control levels = 800 observations per method).
D.2.2. Hypothesis Testing Framework
  • Hypothesis 1: Rank Preservation Across Paradigms. We test whether control coefficients preserve ranking when transferred across paradigms, with quantitative results reported in Table 5:
    • $H_{0}$: No monotonic relationship between source and target CBI
    • $H_{1}$: Control strength ordering transfers (if $\lambda_{A}>\lambda_{B}$ induces higher CBI in the source, it also induces higher CBI in the target)
    Test statistic: Spearman’s rank correlation $\rho$ computed per persona. Significance level: $\alpha=0.01$.
  • Hypothesis 2: Cross-Persona Consistency. We test whether different personas respond uniformly to the same transferred coefficient. Summary statistics are reported in Table 6:
    • Metric: Coefficient of Variation (CV) = $\frac{\text{Std}}{\text{Mean}}\times 100\%$ across personas at each control level
    • Interpretation: CV $10\%$ (Excellent), $10\text{--}20\%$ (Good),
      $20\text{--}30\%$ (Acceptable), $30\%$ (Poor)
    Sample size: 10 personas $\times$ 80 control levels = 800 observations per method.
D.2.3. Interpretation
Across all ten personas, transferred control coefficients preserve relative ordering with near-perfect fidelity, as evidenced by statistically significant monotonic relationships in every case ($p<0.01$) and Spearman correlations exceeding $\rho>0.96$. In addition to rank preservation, cross-persona consistency remains high: the mean coefficient of variation is below 5%, indicating that different personas yield nearly identical CBI values under the same transferred coefficient. The elevated maximum CV observed for RepE Projection (43.71%) occurs at near-zero CBI values, where small absolute deviations disproportionately inflate relative variability and do not reflect substantive behavioral divergence. Taken together, these results provide strong evidence for cross-paradigm generalization, demonstrating that control coefficients transfer reliably with both near-perfect rank preservation and excellent cross-persona consistency, without requiring paradigm-specific recalibration.
D.3. Statistical Validation of Cross-Model Reproducibility
D.3.1. Testing Protocol
We evaluate cross-model reproducibility using four foundation models (Llama, Mistral, DeepSeek, Qwen) across four cognitive bias paradigm pairs.
D.3.2. Hypothesis Testing Framework
  • Hypothesis: CoBRA Methods Reduce Cross-Model Variance. We test whether CoBRA methods produce more consistent behavior across different foundation models compared to baselines. Aggregate ANOVA results are reported in Table 7, with detailed post-hoc comparisons summarized in Table 8.
    • $H_{0}$: No difference in cross-model CBI variance between CoBRA and baselines
    • $H_{1}$: CoBRA methods yield significantly lower cross-model CBI variance than baselines
    Test design: One-way ANOVA comparing three groups (CoBRA mean curve, No Control baseline, Control baseline) across the full coefficient range. Post-hoc pairwise $t$-tests with Cohen’s $d$ effect size are conducted. Significance level: $\alpha=0.01$.
  • Sample Composition:
    • 4 paradigm pairs $\times$ 4 methods = 16 experimental conditions
    • Each condition: 21 control levels $\times$ 4 models = 84 observations
    • Full-range analysis: all 21 control points per condition
    • Total observations for ANOVA: 16 conditions $\times$ 21 points $\times$ 3 groups = 1,008 data points
D.3.3. Statistical Results
Across all paradigm pairs, CoBRA methods exhibit significantly reduced cross-model variance relative to both baselines. As shown in Table 7, one-way ANOVA tests yield highly significant effects ($p<0.001$) for every paradigm pair, with large $F$-statistics indicating strong between-group separation.
Post-hoc pairwise comparisons further confirm this trend. As summarized in Table 8, CoBRA methods consistently outperform both the No Control and Control baselines with large to extremely large effect sizes (mean Cohen’s d=2.41d=-2.41 and 9.20-9.20, respectively), demonstrating robust cross-model reproducibility.
D.3.4. Summary Statistics by Paradigm
To provide an aggregate view of cross-model variance reduction at the paradigm level, Table 9 reports summary statistics comparing CoBRA methods against baselines for each paradigm pair. As shown in Table 9, variance reduction is consistent across all four paradigm pairs (74–81%), indicating that CoBRA’s cross-model reproducibility generalizes robustly across diverse cognitive bias domains.
Paradigm Pair CoBRA Var Baseline Var Reduction (%)
Asch $\leftrightarrow$ Hotel 0.039 0.201 80.6%
Bias Info $\leftrightarrow$ Wason 0.047 0.183 74.3%
Investment $\leftrightarrow$ Asian 0.041 0.176 76.7%
Milgram $\leftrightarrow$ Stanford 0.045 0.188 76.1%
Average across paradigms 0.043 0.187 77.0%
Table 9. Cross-model variance reduction by paradigm pair (CoBRA mean vs baseline mean).
Method Max Pairwise Equiv. Pairs Equiv. Variance
Diff (p¡0.05) Rate Ratio
RepE Projection 0.068 45/45 100% 0.005 (0.5%)
Prompt Numerical 0.117 45/45 100% 0.012 (1.2%)
Table 10. Pairwise equivalence testing: TOST results across 45 temperature pairs. We use an equivalence threshold of $\delta=0.5$ on the 1-5 CBI scale, where maximum differences below 0.5 prove all curves fall within the equivalence margin. The variance ratio (between-temperature variance divided by within-temperature variance) with values below 0.1 confirms robustness across temperature settings.
Method Temperature Pair Mean Diff Cohen’s $|d|$ $p$-value Equiv?
RepE Projection T=0.3 vs T=0.9 0.068 0.545 <0.0001<0.0001 Yes
T=0.2 vs T=0.9 0.059 0.540 <0.0001<0.0001 Yes
T=0.5 vs T=0.9 0.057 0.531 <0.0001<0.0001 Yes
T=0.1 vs T=0.9 0.056 0.517 <0.0001<0.0001 Yes
T=0.4 vs T=0.9 0.054 0.510 <0.0001<0.0001 Yes
Prompt Numerical T=0.1 vs T=0.7 0.117 1.478 <0.0001<0.0001 Yes
T=0.7 vs T=0.9 0.107 1.046 <0.0001<0.0001 Yes
T=0.5 vs T=0.7 0.096 1.248 <0.0001<0.0001 Yes
T=0.3 vs T=0.7 0.093 1.211 <0.0001<0.0001 Yes
T=0.4 vs T=0.7 0.092 1.188 <0.0001<0.0001 Yes
Table 11. Worst-case temperature pairs showing the largest differences (top 5 per method). All differences remain below the 0.5 threshold, with RepE’s maximum difference reaching only 13.6% of the threshold and Prompt’s maximum difference at 23.3% of the threshold.
Method Paradigm Points Equiv. (%) $\Delta_{\text{max}}$ $p_{\text{min}}$
Prompt Numerical Milgram 8 100.0 0.464 $10^{-4}$
Stanford 8 0.0 0.682 0.193
RepE Projection Milgram 21 100.0 0.483 $10^{-6}$
Stanford 21 100.0 0.560 $10^{-5}$
RepE Linear Comb. Milgram 21 100.0 0.596 $10^{-4}$
Stanford 21 100.0 0.344 $10^{-7}$
Table 12. Reasoning mode equivalence testing: TOST results across methods and paradigms.
D.4. Statistical Validation of Temperature Robustness
D.4.1. Testing Protocol
We evaluate temperature robustness using Mistral-7B-Instruct-v0.3 on the Milgram authority paradigm across ten temperature settings, $T\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\}$. For each temperature and method (RepE Projection, Prompt Numerical), we measure CBI across 21 control coefficient levels. We perform pairwise equivalence tests (TOST) across all $\binom{10}{2}=45$ temperature pairs to assess whether control curves remain equivalent within a $\pm 0.5$ threshold on the 1–5 multiple-choice CBI scale.
D.4.2. Hypothesis Testing Framework
  • Hypothesis: Temperature Curves Are Equivalent. We test whether CBI control curves remain equivalent across different sampling temperatures. Aggregate equivalence statistics are summarized in Table 10, while worst-case temperature pair comparisons are reported in Table 11.
    • $H_{0}$: $|\mu_{\text{T1}}-\mu_{\text{T2}}|\geq 0.5$
    • $H_{1}$: $|\mu_{\text{T1}}-\mu_{\text{T2}}|<0.5$
    Test design: Two One-Sided Tests (TOST) for each of 45 temperature pairs per method. Paired $t$-tests are used since CBI values are measured at matched control levels across temperatures. Significance level: $\alpha=0.05$.
  • Sample Composition:
    • 2 methods (RepE Projection, Prompt Numerical)
    • Each method: 10 temperatures $\times$ 21 control levels = 210 observations
    • Pairwise tests: 45 pairs per method = 90 total TOST tests
    • Data alignment: Prompt uses exact matching at 21 fixed control levels; RepE uses linear interpolation to a common grid
D.4.3. Statistical Results
Across both methods, temperature-induced variation is negligible relative to the equivalence threshold. As shown in Table 10, all 45 temperature pairs satisfy the equivalence criterion for both RepE Projection and Prompt Numerical, with maximum pairwise differences of 0.068 and 0.117, respectively—well below the $\delta=0.5$ margin.
Worst-case comparisons further confirm robustness. As detailed in Table 11, even the most divergent temperature pairs exhibit effect sizes and mean differences that remain far within the equivalence region, indicating that control behavior is stable across a wide range of sampling temperatures.
D.4.4. Interpretation
Across all 90 tests (45 temperature pairs $\times$ 2 methods), equivalence is consistently established ($p<0.0001$), yielding a 100% equivalence rate for both methods. Maximum observed differences remain far below the equivalence threshold, with RepE reaching at most 0.068 (13.6% of the threshold) and Prompt Numerical at most 0.117 (23.3%). Consistent with these results, temperature accounts for only a negligible fraction of total behavioral variance, with variance ratios ranging from 0.5% to 1.2%, well below the 0.1 robustness criterion. Collectively, these findings provide strong statistical evidence that CoBRA control coefficients are portable across temperature settings: sampling temperature can be adjusted from near-deterministic ($T=0.1$) to highly stochastic ($T=1.0$) regimes without degrading control precision or requiring re-calibration.
D.5. Statistical Validation of Reasoning Mode Consistency
D.5.1. Testing Protocol
We evaluate the consistency of cognitive bias control across Direct and Reasoning modes using the Mistral-7B-Instruct-v0.3 foundation model with two authority bias paradigms (Milgram obedience and Stanford prison). For each paradigm-method combination, we compare CBI values at matched control coefficient levels between the two reasoning modes.
D.5.2. Hypothesis Testing Framework
  • Hypothesis: RepE Methods Maintain Equivalence Across Reasoning Modes. We test whether CoBRA methods produce equivalent behavior regardless of reasoning complexity:
    • $H_{0}$: Difference between Direct and Reasoning modes exceeds practical equivalence margin ($|\mu_{\text{Direct}}-\mu_{\text{Reasoning}}|\geq\delta$)
    • $H_{1}$: Direct and Reasoning modes are practically equivalent ($|\mu_{\text{Direct}}-\mu_{\text{Reasoning}}|<\delta$)
    Test design: Two One-Sided Tests (TOST) with Welch’s t-test to handle unequal variances. Equivalence margin: $\delta=0.5$. Significance level: $\alpha=0.05$.
  • Sample Composition:
    • 2 paradigms × 3 methods = 6 experimental conditions
    • RepE methods: 21 interpolated control levels per condition
    • Prompt Numerical: 8 discrete control levels per condition
    • Sample sizes: Stanford $n=30$ per level, Milgram $n=75$ per level
D.5.3. Statistical Results
Table 12 summarizes equivalence testing results comparing Direct and Reasoning modes across methods and paradigms. Representation Engineering (RepE) methods achieve perfect equivalence across reasoning modes in all four method–paradigm combinations (100%), with maximum observed differences remaining well within the equivalence margin ($\Delta_{\text{max}}\leq 0.596$) and all corresponding tests highly significant ($p<10^{-4}$). In contrast, Prompt Numerical control exhibits paradigm-dependent behavior: while equivalence is achieved in the Milgram paradigm ($\Delta_{\text{max}}=0.464$, $p<10^{-4}$), performance degrades in the Stanford paradigm, where the maximum difference exceeds the equivalence margin ($\Delta_{\text{max}}=0.682>\delta$) and equivalence is not supported ($p=0.193$).
D.5.4. Interpretation
The statistical analysis provides strong evidence for the superior consistency of Representation Engineering methods. RepE approaches maintain cognitive bias control invariance across reasoning complexity, achieving statistical equivalence at all tested control levels in both experimental paradigms. In contrast, Prompt-based methods exhibit paradigm-specific sensitivity, demonstrating that natural language-based control can be affected by reasoning mode variations.